### v3

# Marginal treatment effects in the absence of instrumental variables

Abstract
We propose a method for defining, identifying, and estimating the marginal treatment effect (MTE) without imposing the instrumental variable (IV) assumptions of independence, exclusion, and separability (or monotonicity). Under a new definition of the MTE based on reduced-form treatment error that is statistically independent of the covariates, we find that the relationship between the MTE and standard treatment parameters holds in the absence of IVs. We provide a set of sufficient conditions ensuring the identification of the defined MTE in an environment of essential heterogeneity. The key conditions include a linear restriction on potential outcome regression functions, a nonlinear restriction on the propensity score, and a conditional mean independence restriction which will lead to additive separability. We prove this identification using the notion of semiparametric identification based on functional forms. And we provide an empirical application for the Head Start program to illustrate the usefulness of the proposed method in analyzing heterogenous causal effects when IVs are elusive.
Keywords: causal inference; treatment effect heterogeneity; point identification; Head Start
JEL Codes: C14, C31, C51
1 Introduction
The marginal treatment effect (MTE), which was developed in a series of seminal papers by Heckman and Vytlacil (1999, 2001, 2005), has become one of the most popular tools in social sciences for describing, interpreting, and analyzing the heterogeneity of the effect of a nonrandom treatment. The MTE is defined as the expected treatment effect conditional on observed covariates and a normalized error term that consists of unobservable determinants of the treatment status. Although the definition of the MTE doesn’t necessitate an instrumental variable (IV), nearly all the existing identification strategies for the MTE rely heavily on valid instrumental variation which can induce otherwise similar individuals into different treatment choices. This is mainly because the MTE has been regarded as an extension of and supplement to the standard IV method in the causal inference literature. However, in practical applications, the validity of IVs is usually difficult to justify. In the most common case of just-identification in which only one IV is available, the exogeneity or randomness of the IV is fundamentally untestable. The exclusion restriction, which is another condition underlying the validity of IVs, has also been challenged more often than not (e.g., van den Berg, 2007; Jones, 2015).
Motivated by the potential invalidity of the available instruments, and inspired by the observation that instruments are unnecessary in the definition of the MTE, we attempt to model, identify, and estimate the MTE without imposing the standard IV assumptions of conditional independence, exclusion, and separability in this paper. Specifically, we allow all the observed covariates to be statistically correlated to the error term and have a direct impact on the outcome in addition to an indirect impact through the treatment variable. The cornerstone of our model is a normalized treatment equation, which determines treatment participation by the propensity score crossing a reduced-form error term that is statistically independent of (although functionally dependent on) all the covariates. This is comparable to the conventional IV-based model, in which normalization is performed with respect to only non-IV covariates. The independence of the reduced-form treatment error from all the covariates can partly justify a conditional mean independence assumption on the potential outcome residuals, which can ensure the additive separability of the MTE into observables and unobservables. This separability then facilitates the semiparametric identification of MTE, given a linear restriction on the potential outcome regression functions and a nonlinear restriction on the propensity score function. We prove the identification by representing the MTE as a known function of conditional moments of observed variables. Intuitively, the identifying power comes from the excluded nonlinear variation in the propensity score, which plays the role of the IV in exogenously perturbing the treatment status.
Compared with imposing the IV assumptions, the main cost of the proposed IV-free method is the linear restriction imposed on the outcome equations and the nonlinearity imposed on the treatment equation, which makes our identification strategy for the MTE closely resemble the semiparametric counterpart of identification based on functional forms (Dong, 2010; Escanciano et al., 2016; Lewbel, 2019; Pan et al., 2022). We build our result on the notion of identification based on functional forms mainly because (i) it can lead to point identification and estimation, (ii) the required assumptions are regular and partly testable, and (iii) the resulting estimation procedures are the most compatible with the standard procedures in IV-based MTE models. Although the primary contribution of this work is on the identification of MTE in the absence of IVs, we also propose a semiparametric MTE estimator tailored to the identification result by implementing the kernel-weighted pairwise difference regression (Ahn and Powell, 1993) for each treatment status, and establish its consistency and asymptotic normality that take into account the nonparametrically estimated propensity score in the first step.
The main value of our IV-free framework of the MTE is threefold. First, it provides an approach for consistently estimating heterogeneous causal effects when a valid IV is hard to find. Second, it suggests an easy-to-implement means for testing the validity of a candidate IV, because it nests the models that assume exclusion restrictions. For instance, in studies on returns to education, the validity of most instruments for educational attainment is suspect, such as parental education, distance to the school, and local labor market conditions (Kédagni and Mourifié, 2020; Mourifié et al., 2020). Our framework enables a reliable evaluation of returns to education without imposing IV assumptions on the candidate instruments and implies a simple test for exclusion restrictions by $t$-testing the instruments’ coefficients in potential outcome equations. Third, even though the validity of the IV is verified, identification based on functional forms can be invoked as a way to increase the efficiency of estimation or to check the robustness of the results to alternative identifying assumptions.
Early explorations of the identification of endogenous regression models when IVs are unavailable focused on linear systems of simultaneous equations, in which the lack of IVs is associated with hardly justifiable exclusion restrictions and insufficient moment conditions. The typical strategy for addressing this underidentification problem is to impose second- or higher-order moment restrictions to construct instruments by using the available exogenous covariates (e.g., Fisher, 1966; Lewbel, 1997; Klein and Vella, 2010). In particular, imposing restrictions on the correlation between the covariates and the variance matrix of the vector of model errors leads to identification based on heteroscedasticity (Lewbel, 2012). D’Haultfœuille et al. (2024) considered a more general nonseparable, nonparametric triangular model without the exclusion restriction, and established identification for the control function approach based on a local irrelevance condition. However, these results are restricted to the case of continuous endogenous variable or treatment. Lewbel (2018) showed that, in linear regression models, the moment conditions of identification based on heteroscedasticity can be satisfied when the endogenous treatment is binary, at the expense of strong restrictions on the model errors. Alternatively, Conley et al. (2012) and Van Kippersluis and Rietveld (2018) presented a sensitivity analysis approach for performing inference for linear IV models while relaxing the exclusion restriction to a certain extent. Kang et al. (2016) proposed an identification and estimation method for linear treatment effect models when the exclusion restriction of some IVs may not hold, and Windmeijer et al. (2019) further proposed a robust method to detect invalid IVs that enter the outcome equation as explanatory variables. More recently, Masten and Poirier (2021) suggested a set identification method for linear IV models under violations of the exclusion restriction. A common limitation of this body of literature is that linear models implicitly impose an undesirable homogeneity assumption on treatment effects. Carl et al. (2025) considered invalid IVs in a treatment effect model with nonlinearity in both the treatment and outcome equations, and established the identification based on different functional forms in the two equations, which is similar to our identification strategy in terms of ideas. Moreover, they adopted machine learning methods in estimating the treatment equation to cope with the potentially high nonlinearity. However, their model implicitly assumes homogenous treatment effects as well. As a complement, our model relaxes the restrictive homogeneity assumption and allows for general heterogeneities in treatment effects, which is evidently more realistic.
The literature on sample selection models without the exclusion restriction is also closely related. A general solution to the problem of lack of excluded variables is partial identification and set estimation (Honoré and Hu, 2020, 2024; Westphal et al., 2022). The limitation of this approach is that the identified or estimated set may be too wide to be informative. Another solution is the approach of identification at infinity that uses the data only for large values of a special regressor (Chamberlain, 1986; Lewbel, 2007) or of the outcome (D’Haultfœuille and Maurel, 2013). Although identification at infinity can lead to point identification, it is typically featured as irregular identification (Khan and Tamer, 2010), and the derived estimate will converge at a rate slower than $n^{-1/2}$, where $n$ is the sample size. Heckman (1979) exploited nonlinearity in the selectivity correction function to achieve point identification and root-$n$ consistent estimation for the linear coefficients of a parametric sample selection model, which is the original version of identification based on functional forms. However, Heckman’s model imposes a restrictive bivariate normality assumption on the error terms and thus poses the risk of model misspecification. Escanciano et al. (2016) extended Heckman’s approach to a general semiparametric model and established identification of the linear coefficients by exploiting nonlinearity elsewhere in the model. At the expense of the generality of their model, Escanciano et al. (2016)’s identification result is only up to scale and thus requires a scale normalization assumption. This normalization would be innocuous if we know the sign of the normalized coefficient and are interested in only the sign, but not the magnitude, of the other coefficients. However, the magnitude of the linear coefficients is indispensable to the evaluation of a program or a policy. In addition, the identifying assumption about nonlinearity developed by Escanciano et al. (2016) implicitly rules out the case of the existence of two continuous covariates. We adapt the result of Escanciano et al. (2016) to the MTE model by amending the two defects. First, we take advantage of the specific model structure to relax the scale normalization and identify the magnitude of the linear coefficients. Moreover, the identifying assumption about nonlinearity will be simplified owing to the specific model structure. We give an intuitive interpretation of the new nonlinearity assumption. Second, we combine the nonlinearity assumption with a local irrelevance condition to allow for an arbitrary number of continuous covariates.
The rest of this paper is organized as follows. In Section 2, we introduce our model and definition of the MTE without IVs. In Section 3, we propose a possible set of sufficient conditions for the identification of MTE, in place of the standard IV assumptions. The key conditions include semiparametric functional form restrictions and the conditional mean independence assumption, of which the latter implies the additive separability of the MTE into observed and unobserved components. Section 4 suggests consistent estimation procedures, and Section 5 provides an empirical application to Head Start. Section 6 concludes. Online appendices comprise of (A) additional results of the empirical application, (B) a proof of the main identification result, (C) a discussion on variants of the nonlinearity assumption, (D) an identification result for limited valued outcomes which entails a slight modification of the assumptions, (E) a proof of the asymptotic normality of the semiparametric IV-free MTE estimator, (F) an implementation guidance, and (G) a simulation study.
2 Model
[…]
$Y=DY_{1}+\left(1-D\right)Y_{0},$
and the quantity of interest is the counterfactual treatment effect $Y_{1}-Y_{0}$. We suppose that the treatment status is determined by the following threshold crossing rule:
$D=1\left\{\mu\left(X\right)\geq U\right\},$ (1)
where $X$ is a vector of pretreatment covariates, $1\left\{A\right\}$ is the indicator function of event $A$, $\mu\left(\cdot\right)$ is an unknown function, and $U$ is a structural error term containing unobserved characteristics that may affect participation in the treatment, such as the opportunity costs or intangible benefits of the treatment.
Compared with the IV-based MTE model, we relax the independence and separability assumptions in the treatment equation (1) to account for the absence of IVs. First, we don’t assume that $X$ is stochastically independent of $U$; that is, no exogenous covariate is needed. Second, we allow the treatment decision rule to be intrinsically nonseparable in observed and unobserved characteristics, which is equivalent to relaxing the monotonicity assumption in the IV model (Vytlacil, 2002, 2006). Specifically, let $U_{x}$ be the counterfactual error term denoting what $U$ would have been if $X$ had been externally set to $x$, then the nonseparability of (1) means that at least two vectors $x$ and $\tilde{x}$ exist in the support of $X$ such that $U_{x}\neq U_{\tilde{x}}$ with positive probability. In particular, $U$ is allowed to depend functionally on $X$, as in the following example.
Example 1.
We consider a latent index rule for treatment participation:
$D=1\left\{m\left(X,\varepsilon\right)\geq 0\right\},$ (2)
[…]
Example 1 illustrates that no generality is lost by modeling the treatment variable as equation (1) without imposing the independence and separability assumptions. We define the propensity score function as the conditional probability of receiving the treatment given the covariates,
[…]
$D=1\left\{P\geq V\right\},$ (3)
where
[…]
Example 2.
We suppose that $X$ is a scalar, and
$\left(\begin{array}[]{l}U\\ X\end{array}\right)\sim N\left(\left(\begin{array}[]{l}0\\ 0\end{array}\right),\left(\begin{array}[]{cc}1&\sigma_{UX}\\ \sigma_{UX}&\sigma_{X}^{2}\end{array}\right)\right).$
[…]
$V=F_{U\left|X\right.}\left(U\left|X\right.\right)=\Phi\left(\frac{U-\mu_{U\left|X\right.}\left(X\right)}{\sigma_{U\left|X\right.}}\right)\text{, and }V_{x}=\Phi\left(\frac{U-\mu_{U\left|X\right.}\left(x\right)}{\sigma_{U\left|X\right.}}\right).$
We observe that $V\perp\!\!\!\!\perp X$ because $F_{V\left|X\right.}\left(v\left|x\right.\right)=v$, but $V_{x}$ is not independent of $X$ because
$F_{V_{x}\left|X\right.}\left(v\left|\tilde{x}\right.\right)=\Phi\left(\frac{\left(\sigma_{UX}\left/\sigma_{X}^{2}\right.\right)\left(x-\tilde{x}\right)}{\sigma_{U\left|X\right.}}+\Phi^{-1}\left(v\right)\right),$
and $V_{x}$ is not uniformly distributed because $F_{V_{x}}\left(v\right)=\Phi\left(\mu_{U\left|X\right.}\left(x\right)+\sigma_{U\left|X\right.}\Phi^{-1}\left(v\right)\right)$.
Given this independence, we may consider the reduced-form treatment error $V$ as the orthogonalized unobservables with respect to the observables, or the unobservables that are projected onto the subspace orthogonal to the one spanned by the observables. Another interpretation of $V$ is the ranking of the structural error $U$ conditional on $X$. For instance, $V=0.2$ represents a typical individual whose $U$ value ranks above 20% individuals with identical covariates. $V$ enters the normalized crossing rule on the right, making an individual less likely to receive treatment; thus, it refers to resistance or distaste regarding the treatment in the MTE literature. If $V$ is large, then the propensity score $P$ should be large to induce the individual to participate in the treatment. However, an individual with a $V$ value close to zero will participate even though $P$ is small.
In the above instrument-free model, we define the MTE as the expected treatment effect conditional on the observed and unobserved characteristics:
$\Delta^{\text{MTE}}\left(x,v\right)\equiv E\left[\left.Y_{1}-Y_{0}\right|X=x,V=v\right].$
$\Delta^{\text{MTE}}\left(x,v\right)$ captures all the treatment effect heterogeneity that is consequential for selection bias by conditioning on the orthogonal coordinates of the observable and unobservable dimensions. Given $X$ and $V$, the treatment status $D$ is fixed and thus independent of the treatment effect $Y_{1}-Y_{0}$. Note that if the assumptions of independence and separability hold for some covariates, the normalized error term $V$ defined in (4) will be exactly the same as that in the IV-based MTE model, and if further the independent covariates have no direct effect on the potential outcomes, our defined MTE will reduce to the MTE that is defined in the IV model. Similar to the MTE in the IV model, the IV-free $\Delta^{\text{MTE}}\left(x,v\right)$ can be used as a building block for constructing the commonly used causal parameters, such as the average treatment effect (ATE), the treatment effect on the treated (TT), the treatment effect on the untreated (TUT), and the local average treatment effect (LATE), which can be expressed as weighted averages of $\Delta^{\text{MTE}}\left(x,v\right)$ as follows:
$\Delta^{\text{ATE}}\left(x\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x\right]=\int_{0}^{1}\Delta^{\text{MTE}}\left(x,v\right)dv,$
$\Delta^{\text{TT}}\left(x\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x,D=1\right]=\frac{1}{\pi\left(x\right)}\int_{0}^{\pi\left(x\right)}\Delta^{\text{MTE}}\left(x,v\right)dv,$
$\Delta^{\text{TUT}}\left(x\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x,D=0\right]=\frac{1}{1-\pi\left(x\right)}\int_{\pi\left(x\right)}^{1}\Delta^{\text{MTE}}\left(x,v\right)dv,$
$\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x,v_{1}\leq V\leq v_{2}\right]=\frac{1}{v_{2}-v_{1}}\int_{v_{1}}^{v_{2}}\Delta^{\text{MTE}}\left(x,v\right)dv.$
Somewhat surprisingly, compared with the weights on the IV-based MTE (e.g., Heckman and Vytlacil, 2005, Table IB), which are generally difficult to estimate in practice, the weights on $\Delta^{\text{MTE}}\left(x,v\right)$ are simpler, more intuitive, and easier to compute.
Heckman and Vytlacil (2001, 2005) considered defining the MTE in a similar nonseparable setting by conditioning on the structural error, and showed how to integrate to generate other causal parameters. However, such an MTE cannot be identified even in the presence of IVs. Our definition, based instead on the reduced-form error, can effectively exploit the statistical independence between the observed and unobserved variables to facilitate identification of the MTE. Zhou and Xie (2019) proposed to redefine the MTE by conditioning on $P$ rather than covariates, which is a more parsimonious specification of all the relevant treatment effect heterogeneity for selection bias. The extension of our identification and estimation procedures to this alternative definition is straightforward.
3 Identification
Our identification strategy relies on a linearity restriction on the potential outcome equations and a nonlinearity restriction on the propensity score function. The intuition is that the propensity score minus the linear outcome index will provide the excluded variation that perturbs treatment status, which plays the role of a continuous IV. Furthermore, to ensure the exogeneity of the excluded variation, it is necessary to impose a certain form of independence between the potential outcome residuals and covariates. We denote $h_{d}\left(x\right)=E\left[\left.Y_{d}\right|X=x\right]$ and $U_{d}=Y_{d}-h_{d}\left(X\right)$, $d=0,1$, as the regression functions and residuals of potential outcomes.
[…]
Assumption CMI is standard in the MTE literature and commonly referred to as separability or additive separability assumption (e.g., Brinch et al., 2017; Mogstad et al., 2018; Zhou and Xie, 2019), because it’s a necessary and sufficient condition for the MTE to be additively separable in observables and unobservables (Zhou and Xie, 2019, p.3074):
[…]
Namely, under Assumption CMI, the shape of the MTE curve with respect to $v$ will not vary with covariates. Assumption CMI is partly justified by $E\left[U_{d}\left|X\right.\right]=0$ and $V\perp\!\!\!\!\perp X$, which come directly from the definition. On this basis, it is sufficient for Assumption CMI to hold if the conditional covariance of $U_{d}$ and $V$ is independent of $X$, which is a key assumption in the heteroscedasticity-based identification method as well (Lewbel, 2012). Assumption CMI is implied by and much weaker than the full independence $\left(U_{d},U\right)\perp\!\!\!\!\perp X$ frequently imposed (often implicitly) in applied work, where $U$ is the structural treatment error in (1). In particular, Assumption CMI doesn’t rule out the marginal dependence of $U_{d}$ or $U$ on $X$, as illustrated by Example 3.
Example 3.
Suppose that $X$ is a scalar and
$\left(\begin{array}[]{c}U_{d}\\ U\\ X\end{array}\right)\sim N\left(\left(\begin{array}[]{c}0\\ 0\\ 0\end{array}\right),\left(\begin{array}[]{ccc}\sigma_{d}^{2}&\sigma_{dU}&0\\ \sigma_{dU}&1&\sigma_{UX}\\ 0&\sigma_{UX}&\sigma_{X}^{2}\end{array}\right)\right).$
In this setting, $U$ is correlated to $X$ with an unconstrained correlation coefficient. Through a property of the multivariate normal distribution, we obtain
$\left.\left(\begin{array}[]{c}U_{d}\\ U\end{array}\right)\right|\left(X=x\right)\sim N\left(\left(\begin{array}[]{c}0\\ \mu_{\left.U\right|X}\left(x\right)\end{array}\right),\left(\begin{array}[]{cc}\sigma_{d}^{2}&\sigma_{dU}\\ \sigma_{dU}&\sigma_{U\left|X\right.}^{2}\end{array}\right)\right),$ (5)
[…]
$E\left[\left.U_{d}\right|U=u,X=x\right]=\frac{\sigma_{dU}}{\sigma_{U\left|X\right.}^{2}}\left(u-\mu_{\left.U\right|X}\left(x\right)\right).$
By Example 2, we have $V=\Phi\left(\left.\left[U-\mu_{U\left|X\right.}\left(X\right)\right]\right/\sigma_{U\left|X\right.}\right)$, so that $U=\sigma_{U\left|X\right.}\Phi^{-1}\left(V\right)+\mu_{U\left|X\right.}\left(X\right)$. Consequently,
$E\left[\left.U_{d}\right|V=v,X=x\right]=E\left[\left.U_{d}\right|U=\sigma_{U\left|X\right.}\Phi^{-1}\left(v\right)+\mu_{U\left|X\right.}\left(x\right),X=x\right]=\frac{\sigma_{dU}}{\sigma_{U\left|X\right.}}\Phi^{-1}\left(v\right),$
and Assumption CMI holds. More generally, to allow the dependence of $U_{d}$ on $X$ as well, we can set
$\left.\left(\begin{array}[]{c}U_{d}\\ U\end{array}\right)\right|\left(X=x\right)\sim N\left(\left(\begin{array}[]{c}0\\ \mu_{\left.U\right|X}\left(x\right)\end{array}\right),\left(\begin{array}[]{cc}\sigma_{d}^{2}\left(x\right)&\sigma_{dU}\\ \sigma_{dU}&\sigma_{U\left|X\right.}^{2}\end{array}\right)\right)$
in place of (5), where $\sigma_{d}^{2}\left(x\right)$ is the conditional variance of $U_{d}$ given $X=x$. Since $E\left[\left.U_{d}\right|V=v,X=x\right]$ is irrelevant to the variance of $U_{d}$ according to the above analysis, Assumption CMI still holds in the presence of such heteroscedastic $U_{d}$.
For the purpose of identifying the MTE, we first establish the relationship between $\Delta^{\text{MTE}}\left(x,v\right)$ and observed regression functions. Under Assumption CMI, the observed regression functions for treated and untreated individuals are additively separable in a similar way:
[…]
$=$ $h_{d}\left(X\right)+g_{d}\left(P\right),\text{ }d=0,1,$
where
$g_{1}\left(p\right)$ $=$ $E\left[\left.U_{1}\right|V\leq p\right]=\frac{1}{p}\int_{0}^{p}E\left[\left.U_{1}\right|V=v\right]dv,$
$g_{0}\left(p\right)$ $=$ $E\left[\left.U_{0}\right|V>p\right]=\frac{1}{1-p}\int_{p}^{1}E\left[\left.U_{0}\right|V=v\right]dv.$
The function $g_{d}\left(\cdot\right)$ is usually referred to as selectivity bias correction term and plays a key part in the literature on sample selection models. We introduce $g_{d}\left(\cdot\right)$ into the identification of MTE because $g_{d}\left(\cdot\right)$ is closely related to the MTE and it is generally easier to identify $g_{d}\left(\cdot\right)$ than directly identify the MTE. From its definition, $g_{d}\left(\cdot\right)$ can be viewed as an aggregating function of the unobservable heterogeneity of the MTE. By multiplying $g_{d}\left(p\right)$ with $p$ or 1p1-p and then differentiating with respect to $p$, we can retrieve the unobservable part of the MTE as
$E\left[\left.U_{1}\right|V=p\right]$ $=$ $g_{1}\left(p\right)+pg_{1}^{\left(1\right)}\left(p\right),$
$E\left[\left.U_{0}\right|V=p\right]$ $=$ $g_{0}\left(p\right)-\left(1-p\right)g_{0}^{\left(1\right)}\left(p\right),$
where $g_{d}^{\left(1\right)}$ denotes the derivative function of $g_{d}$. Therefore, the MTE can be represented as
$\Delta^{\text{MTE}}\left(x,v\right)=h_{1}\left(x\right)-h_{0}\left(x\right)+\left[g_{1}\left(v\right)-g_{0}\left(v\right)\right]+vg_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)g_{0}^{\left(1\right)}\left(v\right),$ (7)
and the identification of it can be achieved by identifying $h_{d}$ and $g_{d}$ from the regression equation (6).
Recall that $P=\pi\left(X\right)$, where $\pi$ is a deterministic function, albeit unknown. Without having an IV, especially without having the exclusion restriction that assumes some element of $X$ to drop out of $h_{d}\left(X\right)$, the argument of $g_{d}$ in (6) exhibits no extra variation than $h_{d}$. Therefore, it is impossible to distinguish between $g_{d}$ and $h_{d}$. For example, one may choose to let $h_{d}\left(\cdot\right)$ absorb $g_{d}\left(\pi\left(\cdot\right)\right)$ and set $g_{d}=0$. Inspired by the identification result of Escanciano et al. (2016) for a double index model, functional form restrictions that differentiate $h_{d}\left(\cdot\right)$ and $g_{d}\left(\pi\left(\cdot\right)\right)$ from one another can address this issue and facilitate the identification of both $h_{d}$ and $g_{d}$. To this end, we impose linearity on $h_{d}$ and nonlinearity on $\pi$.
Assumption L (Linearity). Assume that $E\left[\left.Y_{d}\right|X\right]=\alpha_{d}+X^{\prime}\beta_{d}$ with probability one for some fixed $\alpha_{d}$ and $\beta_{d}$, $d=0,1$.
Assumption L imposes a linear restriction on the potential outcome regression functions, which is a common practice in empirical studies. The linear restriction may seem too strong compared with the existing identification strategies for the MTE, which generally allow highly flexible (especially nonparametric) specifications on the potential outcomes. However, when the identification is accomplished and the estimation comes into consideration, the linear specification will turn to be nearly universally adopted due to its tractability and interpretability (e.g., Kirkeboen et al., 2016; Kline and Walters, 2016; Brinch et al., 2017; Heckman et al., 2018; Mogstad et al., 2021; Aryal et al., 2022; Mountjoy, 2022). In addition to the empirical convenience, the linear regression relationship may also be derived from a multivariate normal distribution of $\left(Y_{d},X\right)$, or justified by economic models such as the Cobb-Douglas production function which implies that the logarithm of an output is linearly related to the logarithm of inputs, the capital asset pricing model (CAPM) which implies that the expected return of a portfolio is linearly related to the portfolio risk, and the concavely quadratic utility function which implies a linear demand system for differentiated products (Amir et al., 2017).
With a little abuse of notation, we redenote $U_{d}$ to absorb the intercept $\alpha_{d}$ so that
$Y_{d}=X^{\prime}\beta_{d}+U_{d},$
and thus that the regression equation (6) and the MTE become
[…]
$\Delta^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\left(\beta_{1}-\beta_{0}\right)+\left[g_{1}\left(v\right)-g_{0}\left(v\right)\right]+vg_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)g_{0}^{\left(1\right)}\left(v\right).$ (9)
Noting that Assumption L implicitly requires the potential outcomes to be continuously distributed and supported on the entire real line, we would like to point out that our identification strategy can be adapted to the case of limited valued outcomes as well by specifying a linear latent index. A detailed discussion is left to Appendix D.
To present the nonlinearity assumption, we let $X$ be partitioned as $\left(X^{C},X^{D}\right)$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. We denote $X_{k}$, $X_{k}^{C}$, and $X_{k}^{D}$ as the $k$-th coordinates of $X$, $X^{C}$, and $X^{D}$, respectively. And we denote $x^{C}$ as a generic element in the support of $X^{C}$; likewise for $x^{D}$, $x_{k}^{C}$, and $x_{k}^{D}$. The nonlinearity assumption requires the propensity score function $\pi$ to be nonlinear in $X^{C}$ given a benchmark value of $X^{D}$. Without loss of generality, we suppose that the vector of zeros is in the support of $X^{D}$ and is the benchmark value. We denote $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ for notational convenience, where the discrete covariates are equal to zero.
[…]
– NL2 ($\dim\left(X^{C}\right)\geq 2$): there exist two vectors $x^{C}$, $\tilde{x}^{C}$ in the support of $X^{C}$ and two elements $k$, $j$ in set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq 0$, (ii) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$, (iv) $\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$, and (v) $\left.\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)\neq\left.\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\right/\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)$, where $\partial_{k}\pi_{0}\left(x^{C}\right)=\left.\partial\pi_{0}\left(x^{C}\right)\right/\partial x_{k}^{C}$ is the partial derivative of $\pi_{0}$ with respect to the $k$-th argument.
– NL1 ($\dim\left(X^{C}\right)=1$): there exists a constant $\tilde{x}^{C}$ in the support of $X^{C}$ such that $\pi_{0}^{\left(1\right)}\left(\tilde{x}^{C}\right)=0$, where $\pi_{0}^{\left(1\right)}\left(x^{C}\right)=\left.d\pi_{0}\left(x^{C}\right)\right/dx^{C}$ is the univariate derivative of $\pi_{0}$.
Assumption NL requires the propensity score function to be nonlinear in continuous covariates in a generalized sense. Note that Assumption NL is sufficient but not necessary for nonlinearity of $\pi_{0}$. Appendix C provides other possible forms of sufficient nonlinearity conditions. The combination of Assumptions L and NL will enable identification based on functional forms in a semiparametric version, which can realize the identification of linear coefficients by exploiting nonlinearity elsewhere in the model. Actually, the nonlinear or nonparametric setting of the treatment equation is typically adopted in the applied literature when semiparametric approaches to estimating the MTE are taken, as illustrated in Carneiro et al. (2011) and summarized in Cornelissen et al. (2016). [^1] We thank an anonymous referee for pointing out this observation. The following example provides an economic intuition that underlies the nonlinearity assumption.
Example 4.
Consider the generalized Roy model where individuals choose treatment if the perceived benefit from treatment is greater than the subjective cost of treatment (Eisenhauer et al., 2015). Specifically,
$D=1\left\{Y_{1}-Y_{0}\geq c\left(X\right)+\epsilon\right\},$ (10)
where $c\left(X\right)$ and $\epsilon$ are the observed and unobserved parts of the cost, respectively. Under Assumption L, the treatment equation (10) becomes $D=1\left\{X^{\prime}\left(\beta_{1}-\beta_{0}\right)-c\left(X\right)\geq U\right\}$, where $U=\epsilon-U_{1}+U_{0}$, therefore $\pi\left(x\right)=F_{U\left|X\right.}\left(\left.x^{\prime}\left(\beta_{1}-\beta_{0}\right)-c\left(x\right)\right|x\right)$, which will be a nonlinear function of $x$ if $c\left(x\right)$ is nonlinear or the conditional distribution function of $U$ given $X=x$ is nonlinear. In economic applications, the cost function $c\left(x\right)$ is typically nonlinear in individual and family characteristics. For instance, the subjective cost of higher education would be U-shaped with respect to the family income, because of the relatively high tuition fee for students with disadvantaged backgrounds and high opportunity cost for students with superior backgrounds. On the other hand, the conditional distribution function $F_{U\left|X\right.}\left(\left.u\right|x\right)$ will be nonlinear in $x$ in general if $U$ is not independent of $X$, as shown by Example 6. More generally, the treatment equation (10) is a special case of a utility maximization model of treatment defined as $D=1\left\{\mathfrak{u}_{1}\left(Y_{1},X,\epsilon_{1}\right)\geq\mathfrak{u}_{0}\left(Y_{0},X,\epsilon_{0}\right)\right\}$, where the utility functions $\mathfrak{u}_{1}$ and $\mathfrak{u}_{0}$ are quasi-linear with respect to the first argument. An early version of Lee and Park (2023) gave further examples illustrating why the nonlinearity of the utility functions makes sense in economic studies.
The nonlinearity assumption has two different forms, depending on the number of continuous covariates. When two or more continuous covariates are available, Assumption NL2 will require some variation in $\pi_{0}$ to distinguish it from a single-index function. Concretely, NL2 will not hold if $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ for a smooth function $f$, because in this case, both sides of the inequality in (v) are equal to $\left.\gamma_{k}\right/\gamma_{j}$. Otherwise, however, it is difficult to construct examples that violate (v). Therefore, we can informally test NL2 by constructing a test for the single-index specification of the propensity score against a general nonparametric alternative (e.g., Fan and Li, 1996; Stute and Zhu, 2005; Xia, 2009; Maistre and Patilea, 2019). In practice, NL2 can be fulfilled even when $\pi_{0}$ is single-index specified, if interaction or/and quadratic terms are added, as shown by the following Example.
Example 5.
Consider the case of two continuous covariates. Suppose that for a smooth function $f$, $\pi_{0}\left(x^{C}\right)=f\left(\gamma_{1}x_{1}^{C}+\gamma_{2}x_{2}^{C}+\gamma_{3}x_{1}^{C}x_{2}^{C}\right)$ or $\pi_{0}\left(x^{C}\right)=f\left(\gamma_{1}x_{1}^{C}+\gamma_{2}x_{2}^{C}+\gamma_{3}\left(x_{1}^{C}\right)^{2}\right)$. Then, we obtain $\left.\partial_{1}\pi_{0}\left(x^{C}\right)\right/\partial_{2}\pi_{0}\left(x^{C}\right)=\left.\left(\gamma_{1}+\gamma_{3}x_{2}^{C}\right)\right/\left(\gamma_{2}+\gamma_{3}x_{1}^{C}\right)$ for the interaction case, or $\left.\partial_{1}\pi_{0}\left(x^{C}\right)\right/\partial_{2}\pi_{0}\left(x^{C}\right)=\left.\left(\gamma_{1}+2\gamma_{3}x_{1}^{C}\right)\right/\gamma_{2}$ for the quadratic case. In both cases, Assumption NL2.(v) will generally hold for $x^{C}$ and $\tilde{x}^{C}$ satisfying $x_{1}^{C}\neq\tilde{x}_{1}^{C}$.
Assumption NL2 is a form of the nonlinearity assumption of Escanciano et al. (2016) that is specific to our model, and thus inherits a demerit that it requires the existence of at least two continuous covariates. However, in empirical studies based on survey data, most of the demographic characteristics are documented as discrete or categorical variables, such as age, gender, race, marital status, educational attainment, and so on. Therefore, we also impose Assumption NL1, as a supplement to NL2, to take into account the situation in which only one continuous covariate is available. NL1 requires $\pi_{0}$ to have at least one stationary point. Namely, NL1 holds if the probability of receiving treatment is unaffected by some local change in the continuous covariate. It’s a differential version of the local irrelevance assumption imposed in nonseparable models for attaining point identification (e.g., Torgovitsky, 2015; D’Haultfœuille et al., 2024). Note that NL1 will hold if $\pi_{0}$ is flat everywhere, i.e., if the continuous covariate has no influence on the treatment choice conditional on the benchmark value of $X^{D}$. However, this extreme case will be ruled out by the subsequent Assumption S. Given that, NL1 implies that $\pi_{0}$ is necessarily nonlinear. NL1 can be extended to the case of no continuous covariate, as discussed in Appendix C. However, using only discrete covariates provides not much identifying variation, which may lead to poor performance in the subsequent model estimation (Garlick and Hyman, 2022). Hence, we focus on the case of at least one continuous covariate. NL1 can be visually tested by plotting the estimated propensity score as a function of the continuous covariate given a benchmark value of $X^{D}$, as illustrated in Figure 1 in the subsequent empirical application.
We note that Assumption NL doesn’t exclude the widely specified linear-index treatment equation, as long as the structural error is not independent of covariates.
Example 6.
Suppose the treatment status is determined by a linear-index threshold crossing rule:
$D=1\left\{X^{\prime}\gamma\geq U\right\},$
where $U$ is not independent of $X$. Namely, suppose $\mu\left(X\right)=X^{\prime}\gamma$ in the structural treatment equation (1). Consider a multiplicatively heteroscedastic $U$ such that $U=\sigma\left(X\right)\tilde{U}$, where $\sigma\left(X\right)$ is a positive function and $\tilde{U}$ is an idiosyncratic error independent of $X$. The normalization derives
$D=1\left\{\frac{X^{\prime}\gamma}{\sigma\left(X\right)}\geq\tilde{U}\right\}=1\left\{F_{\tilde{U}}\left(\frac{X^{\prime}\gamma}{\sigma\left(X\right)}\right)\geq V\right\},$
[…]
Unlike in the binary response model, the ubiquitous heteroscedasticity and endogeneity benefit our results while inducing no trouble, because the identification of MTE is irrelevant to the structural coefficients in $\gamma$. Our MTE is defined by the reduced-form treatment error; thus, all we need from the treatment equation is the propensity score, which has a reduced-form nature.
Assumption NL ensures identification of coefficients of continuous covariates. In order to identify coefficients of discrete covariates, we impose a mild support condition. For any $x_{k}^{D}\neq 0$, we denote $x^{Dk}$ as the $\dim\left(X^{D}\right)\times 1$ vector with the $k$-th coordinate being equal to $x_{k}^{D}$ and all the other coordinates being equal to zero.
Assumption S (Support). For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, assume for some $x_{k}^{D}\neq 0$ in the support of $X_{k}^{D}$ that there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$.
A sufficient condition for Assumption S to hold is that $\pi_{0}\left(X^{C}\right)$ has a full support on the unit interval, or, more generally, that the support of $\pi_{0}\left(X^{C}\right)$ is overlapped with those of $\pi\left(X^{C},x^{D}\right)$ for all $x^{D}\neq 0$. Otherwise, we have to find an $x_{k}^{D}\neq 0$ for each $k$ such that the support of $\pi\left(X^{C},x^{Dk}\right)$ is overlapped with that of $\pi_{0}\left(X^{C}\right)$. The following theorem establishes our main identification result.
Theorem 1.
If Assumptions CMI, L, NL, and S hold, then $\beta_{d}$ and $g_{d}\left(p\right)$ at all $p$ in the support of the propensity score $P$ are identified for $d=0,1$.
Theorem 1 is an adaption of the identification result of Escanciano et al. (2016). Since Escanciano et al. (2016) considered a general double index model with an unknown link function, their identification result is only up to scale and thus requires a scale normalization assumption. In comparison, Theorem 1 can identify the magnitude of the linear coefficients without any normalization. The proof of Theorem 1 is delegated to Appendix B. Briefly speaking, the proof is grounded on the observed regression functions (8), which summarize the information from the data. As $g_{d}$ is unknown, we need to eliminate it through some subtraction to realize the identification of $\beta_{d}$. When only one continuous covariate exists, the subtraction can be carried out locally around the point satisfying Assumption NL1. Otherwise, our strategy is to perturb two continuous covariates, such as $x_{k}^{C}$ and $x_{j}^{C}$, in such a way that $\pi\left(x\right)$ remains unchanged. Specifically, for each group $d$, we increase $x_{k}^{C}$ by a small $\epsilon$ and simultaneously change $x_{j}^{C}$ by $\epsilon$ multiplied by $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$, resulting in a perturbed value of the regression function. Subtracting the perturbed regression function from the original (8) will cancel out $g_{d}$ due to the equality of $\pi\left(x\right)$, giving rise to an equation for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Note that the multiplier $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$ is the partial derivative of $x_{j}^{C}$ with respect to $x_{k}^{C}$ if we view $x_{j}^{C}$ as an implicit function of the other continuous covariates by equating $\pi_{0}\left(x^{C}\right)$ to a constant. Accordingly, Assumption NL2.(v) implies two linearly independent equations, ensuring an exact solution (i.e., identification) of $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$.
It is worth mentioning that our strategy naturally features over-identification in the sense that if there is a pair of points satisfying Assumption NL2, then there must be infinitely many pairs of points for which NL2 holds, because $X^{C}$ is continuously distributed. Moreover, there will generally be more than one value of $X^{D}$ that satisfies Assumptions S, as illustrated by the application in Section 5. Such values of $X^{D}$ can be taken as alternative benchmarks, under which Assumption NL may hold for another set of (pairs of) points of $X^{C}$. Consequently, the identification can be represented as the average of solutions over all (pairs of) points of $X^{C}$ satisfying Assumption NL and over all values of $X^{D}$ satisfying Assumptions S.
According to (9), Theorem 1 implies the identification of MTE in the absence of IVs. Specifically, This result allows practitioners to include all the relevant observed characteristics into both treatment and outcome equations, without imposing any exclusion or full independence assumptions. Under Theorem 1, the conventional causal parameters can also be identified without instruments, provided that the support of $P$ contains 0 and/or 1 (which implies identifiability of $g_{d}\left(0\right)$ and/or $g_{d}\left(1\right)$):
$\Delta^{\text{ATE}}\left(x\right)$ $=$ $x^{\prime}\left(\beta_{1}-\beta_{0}\right)+\left[g_{1}\left(1\right)-g_{0}\left(0\right)\right],$ (11)
$\Delta^{\text{TT}}\left(x\right)$ $=$ $x^{\prime}\left(\beta_{1}-\beta_{0}\right)+g_{1}\left(\pi\left(x\right)\right)+\frac{\left(1-\pi\left(x\right)\right)g_{0}\left(\pi\left(x\right)\right)-g_{0}\left(0\right)}{\pi\left(x\right)},$
$\Delta^{\text{TUT}}\left(x\right)$ $=$ $x^{\prime}\left(\beta_{1}-\beta_{0}\right)+\frac{g_{1}\left(1\right)-\pi\left(x\right)g_{1}\left(\pi\left(x\right)\right)}{1-\pi\left(x\right)}-g_{0}\left(\pi\left(x\right)\right),$
$\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $x^{\prime}\left(\beta_{1}-\beta_{0}\right)+\frac{v_{2}g_{1}\left(v_{2}\right)-v_{1}g_{1}\left(v_{1}\right)+\left(1-v_{2}\right)g_{0}\left(v_{2}\right)-\left(1-v_{1}\right)g_{0}\left(v_{1}\right)}{v_{2}-v_{1}}.$
4 Estimation
Our identification strategy for the IV-free MTE implies a separate estimation procedure that works with the partially linear regression (8) by each treatment status. In particular, we recommend the kernel-weighted pairwise difference estimation method proposed by Ahn and Powell (1993) because of its computational simplicity, well-established asymptotic properties, and, most importantly, its capacity to effectively leverage the overidentifying information that underlies the data. Suppose that $\left\{\left(Y_{i},D_{i},X_{i}\right):i=1,2,\cdots,n\right\}$ is a random sample of observations on $\left(Y,D,X\right)$. In the first step, we estimate the nonparametrically specified propensity score using the kernel method, that is,
$\hat{\pi}\left(x\right)=\frac{\sum_{i=1}^{n}D_{i}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\left.\left(X_{il}^{C}-x_{l}^{C}\right)\right/h_{1l}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}}{\sum_{i=1}^{n}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\left.\left(X_{il}^{C}-x_{l}^{C}\right)\right/h_{1l}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}},$ (12)
and construct $\hat{P}_{i}=\hat{\pi}\left(X_{i}\right)$ by leaving the $i$-th observation out in the estimation, where $h_{1l}$, $l=1,2,\cdots,\dim\left(X^{C}\right)$, are bandwidths and $k_{1}$ is a univariate kernel function. If the dimension of $X^{D}$ is large, then a smoothed kernel for discrete covariates (Racine and Li, 2004) can be applied as a substitute for the indicator function, to alleviate the potential problem of inadequate observations in each data cell divided by the support of $X^{D}$. If the number of continuous covariates is not small either, then the well-known curse of dimensionality will appear, and a linear-index specification may thus be practically more relevant when modeling the propensity score. The index should include a series of interaction terms and quadratic or even higher-order terms of continuous covariates to supply sufficient nonlinear variation. The linear-index propensity score can be estimated by parametric probit/logit or semiparametric methods (e.g., Powell et al., 1989; Ichimura, 1993; Klein and Spady, 1993; Lewbel, 2000), depending on the distributional assumption on the error term.
In the second step, we estimate $\beta_{d}$ for each $d$ through a weighted pairwise difference least squares regression:
$\hat{\beta}_{d}$ $=$ $\arg\min_{\beta}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\hat{\omega}_{dij}\left[\left(Y_{i}-Y_{j}\right)-\left(X_{i}-X_{j}\right)^{\prime}\beta\right]^{2}$ (13)
$=$ $\left[\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\hat{\omega}_{dij}\left(X_{i}-X_{j}\right)\left(X_{i}-X_{j}\right)^{\prime}\right]^{-1}\left[\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\hat{\omega}_{dij}\left(X_{i}-X_{j}\right)\left(Y_{i}-Y_{j}\right)\right],$
where the weights are given by
$\hat{\omega}_{dij}=1\left\{D_{i}=D_{j}=d\right\}\frac{1}{h_{2}}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right),$
with $h_{2}$ and $k_{2}$ being the bandwidth and kernel function, respectively, which can be different from those in the first step. Given $\hat{\beta}_{d}$, the nonlinear function $g_{d}\left(p\right)$ and its derivative function $g_{d}^{\left(1\right)}\left(p\right)$ at any $p$ in the support of $P$ can be estimated by the local linear method, namely,
$\left(\begin{array}[]{c}\hat{g}_{d}\left(p\right)\\ \hat{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)=\left[\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)^{\prime}\right]^{-1}\left[\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\hat{\beta}_{d}\right)\right],$
where
$\hat{w}_{di}\left(p\right)=1\left\{D_{i}=d\right\}\frac{1}{h_{3}}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right).$
Finally, we plug $\hat{\beta}_{d}$, $\hat{g}_{d}$, $\hat{g}_{d}^{\left(1\right)}$, and $\hat{\pi}$ into identification equations (9) and (11) to estimate the IV-free MTE and other causal parameters, as follows:
$\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\left[\hat{g}_{1}\left(v\right)-\hat{g}_{0}\left(v\right)\right]+v\hat{g}_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)\hat{g}_{0}^{\left(1\right)}\left(v\right),$ (14)
$\hat{\Delta}^{\text{ATE}}\left(x\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\left[\hat{g}_{1}\left(1\right)-\hat{g}_{0}\left(0\right)\right],$
$\hat{\Delta}^{\text{TT}}\left(x\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\hat{g}_{1}\left(\hat{\pi}\left(x\right)\right)+\frac{\left(1-\hat{\pi}\left(x\right)\right)\hat{g}_{0}\left(\hat{\pi}\left(x\right)\right)-\hat{g}_{0}\left(0\right)}{\hat{\pi}\left(x\right)},$
$\hat{\Delta}^{\text{TUT}}\left(x\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\frac{\hat{g}_{1}\left(1\right)-\hat{\pi}\left(x\right)\hat{g}_{1}\left(\hat{\pi}\left(x\right)\right)}{1-\hat{\pi}\left(x\right)}-\hat{g}_{0}\left(\hat{\pi}\left(x\right)\right),$
$\hat{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\frac{v_{2}\hat{g}_{1}\left(v_{2}\right)-v_{1}\hat{g}_{1}\left(v_{1}\right)+\left(1-v_{2}\right)\hat{g}_{0}\left(v_{2}\right)-\left(1-v_{1}\right)\hat{g}_{0}\left(v_{1}\right)}{v_{2}-v_{1}}.$
Theorem 2.
Suppose the assumptions of Theorem 1 hold. Further suppose Assumptions E.1-E.4 given in Appendix E hold. Then for any interior point $v$ of the support of $P$ and any $x$ satisfying $0<\pi\left(x\right)<1$, we have
$\sqrt{nh_{3}^{3}}\left[\hat{\Delta}^{\text{MTE}}\left(x,v\right)-\Delta^{\text{MTE}}\left(x,v\right)\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{MTE}}^{2}\left(v\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{ATE}}\left(x\right)-\Delta^{\text{ATE}}\left(x\right)-b_{\text{ATE}}h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{ATE}}^{2}\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{TT}}\left(x\right)-\Delta^{\text{TT}}\left(x\right)-b_{\text{TT}}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{TT}}^{2}\left(\pi\left(x\right)\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{TUT}}\left(x\right)-\Delta^{\text{TUT}}\left(x\right)-b_{\text{TUT}}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{TUT}}^{2}\left(\pi\left(x\right)\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)-\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)-b_{\text{LATE}}\left(v_{1},v_{2}\right)h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{LATE}}^{2}\left(v_{1},v_{2}\right)\right),$
where the asymptotic variances $\sigma_{\text{MTE}}^{2},\cdots,\sigma_{\text{LATE}}^{2}$ are defined in (E.21)-(E.25), respectively, and the asymptotic bias terms are defined as
$b_{\text{ATE}}$ $=$ $\frac{\kappa_{2}}{2}\left(g_{1}^{\left(2\right)}\left(1\right)-g_{0}^{\left(2\right)}\left(0\right)\right),$
$b_{\text{TT}}\left(p\right)$ $=$ $\frac{\kappa_{2}}{2}\left(g_{1}^{\left(2\right)}\left(p\right)+\frac{\left(1-p\right)g_{0}^{\left(2\right)}\left(p\right)-g_{0}^{\left(2\right)}\left(0\right)}{p}\right),$
$b_{\text{TUT}}\left(p\right)$ $=$ $\frac{\kappa_{2}}{2}\left(\frac{g_{1}^{\left(2\right)}\left(1\right)-pg_{1}^{\left(2\right)}\left(p\right)}{1-p}-g_{0}^{\left(2\right)}\left(p\right)\right),$
$b_{\text{LATE}}\left(v_{1},v_{2}\right)$ $=$ $\frac{\kappa_{2}}{2}\left(\frac{v_{2}g_{1}^{\left(2\right)}\left(v_{2}\right)-v_{1}g_{1}^{\left(2\right)}\left(v_{1}\right)+\left(1-v_{2}\right)g_{0}^{\left(2\right)}\left(v_{2}\right)-\left(1-v_{1}\right)g_{0}^{\left(2\right)}\left(v_{1}\right)}{v_{2}-v_{1}}\right),$
with $\kappa_{2}=\int k_{3}\left(u\right)u^{2}du$.
The semiparametric MTE estimator $\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ converges at the same rate $O_{p}\left(\left(nh_{3}^{3}\right)^{-1/2}\right)$ as the local linear estimator for the derivative of a regression function (e.g., Li and Racine, 2007, Theorem 2.7), while the estimators of the aggregated causal parameters converge at the same rate $O_{p}\left(\left(nh_{3}\right)^{-1/2}\right)$ as the kernel estimator for a regression function itself. This is because the estimation of MTE involves plugging in $\hat{g}_{d}^{\left(1\right)}$, the nonparametric derivative estimate, which converges slower than $\hat{g}_{d}$, the nonparametric regression function estimate. Taking the usually applied rule-of-thumb bandwidth $h_{3}\sim n^{-1/5}$ for instance, the MTE estimator will converge at the rate $O_{p}\left(n^{-1/5}\right)$ and the aggregated causal parameters will converge at the rate $O_{p}\left(n^{-2/5}\right)$. Since the slope estimate $\hat{\beta}_{d}$ converges at the parametric rate $O_{p}\left(n^{-1/2}\right)$ that is always faster than $\hat{g}_{d}^{\left(1\right)}$ and $\hat{g}_{d}$, the observed part $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)$ is asymptotically negligible and has no impact on the asymptotic distributions of the estimators. Therefore, for the causal parameters that $x$ does not enter the unobserved part such as the MTE, ATE, and LATE, the asymptotic bias and variance of their estimators will not depend on the observed characteristics.
The kernel-weighted pairwise difference estimator has the advantage of having a closed-form expression, so we need not solve any formidable optimization problems. However, it faces the challenging problem of bandwidth selection as most semiparametric estimation methods. Alternatively, we can consider imposing a parametric specification on the unobservable heterogeneity of the MTE such that $E\left[\left.U_{d}\right|V=v\right]=E\left[\left.U_{d}\right|V=v;\theta_{d}\right]$ for finite dimensional $\theta_{d}$ (Heckman and Vytlacil, 2005), e.g., the polynomial specification such that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}v^{j}$ or the normal polynomial specification such that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}\Phi^{-j}\left(v\right)$, for a fixed $J$. In the latter, setting $J=1$ will match Heckman’s normal sample selection model. Under the parametric restriction, the second step becomes a global regression $E\left[Y\left|X,D=d\right.\right]=X^{\prime}\beta_{d}+g_{d}\left(P;\theta_{d}\right)$ with parameterized selectivity bias correction term $g_{d}\left(p;\theta_{d}\right)$ for each $d$; therefore, the tuning of $h_{2}$ and $h_{3}$ is circumvented. Another advantage of a parametrically specified second step is the lower computational burden relative to the pairwise difference estimator that is defined by double summation which requires a squared amount of calculation (Pan and Xie, 2023). The simulation in Appendix G shows that MTE estimators with parametric second-step perform well when the specification is correct or nearly correct, while the semiparametric MTE estimator performs robustly across different designs.
Notably, the local IV (LIV) estimation procedure can be adapted to our model as well, though it needs no IV. Unlike the separate estimation procedure, the adapted LIV approach is based on a whole-sample regression:
[…]
$=$ $\alpha_{0}+X^{\prime}\beta_{0}+PX^{\prime}\left(\beta_{1}-\beta_{0}\right)+q\left(P\right),$
where
$q\left(p\right)=pE\left[U_{1}-U_{0}\left|V\leq p\right.\right]=\int_{0}^{p}E\left[U_{1}-U_{0}\left|V=v\right.\right]dv.$
Note that since
[…]
the MTE is equal to the derivative of the regression function (15) with respect to $P$. As a consequence, it would be sufficient to estimate $\beta_{d}$ for $d=0,1$ and functions $q$ and $q^{\left(1\right)}$. Given the estimated propensity score, we can likewise use the pairwise difference principle to obtain
$\left(\begin{array}[]{c}\check{\beta}_{0}\\ \check{\delta}\end{array}\right)$ $=$ $\left[\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)^{\prime}\right]^{-1}$
$\cdot\left[\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)\left(Y_{i}-Y_{j}\right)\right],$
where $\check{\delta}$ is an estimator for $\beta_{1}-\beta_{0}$ according to (15). Given $\check{\beta}_{0}$ and $\check{\delta}$, we apply the local linear method as well, yielding
$\left(\begin{array}[]{c}\check{r}\left(p\right)\\ \check{q}^{\left(1\right)}\left(p\right)\end{array}\right)$ $=$ $\left[\sum_{i=1}^{n}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)^{\prime}\right]^{-1}$
$\cdot\left[\sum_{i=1}^{n}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\check{\beta}_{0}-\hat{P}_{i}X_{i}^{\prime}\check{\delta}\right)\right],$
where $\check{r}\left(p\right)$ is an estimator for $\alpha_{0}+q\left(p\right)$. Then, we construct the instrument-free LIV estimators for the MTE and other treatment parameters as
$\check{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\check{\delta}+\check{q}^{\left(1\right)}\left(v\right),$
$\check{\Delta}^{\text{ATE}}\left(x\right)$ $=$ $x^{\prime}\check{\delta}+\check{r}\left(1\right)-\check{r}\left(0\right),$
$\check{\Delta}^{\text{TT}}\left(x\right)$ $=$ $x^{\prime}\check{\delta}+\frac{\check{r}\left(\hat{\pi}\left(x\right)\right)-\check{r}\left(0\right)}{\hat{\pi}\left(x\right)},$
$\check{\Delta}^{\text{TUT}}\left(x\right)$ $=$ $x^{\prime}\check{\delta}+\frac{\check{r}\left(1\right)-\check{r}\left(\hat{\pi}\left(x\right)\right)}{1-\hat{\pi}\left(x\right)},$
$\check{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $x^{\prime}\check{\delta}+\frac{\check{r}\left(v_{2}\right)-\check{r}\left(v_{1}\right)}{v_{2}-v_{1}}.$
[…]
In summary, the identification based on functional forms can accommodate most of the frequently-used estimation procedures in IV-based MTE models. However, since it is infeasible to manipulate the nonlinear variation, our identification result does not allow for the counterfactual analysis via estimating the policy-relevant treatment effect (PRTE) which is also an important application of IV-based MTE models (Heckman and Vytlacil, 2005).
5 Empirical application
In this section, we revisit the heterogeneous long-term effects of the Head Start program on educational attainment and labor market outcomes by using the proposed IV-free MTE method. Head Start, which began in 1965, is one of the largest early child care programs in the United States. The program is targeted at children from low-income families and can provide such children with preschool, health, and nutritional services. Currently, Head Start serves more than a million children, at an annual cost of 10 billion dollars. As a federally funded large-scale program, Head Start has encountered concerns about its effectiveness and thus spawned numerous studies to evaluate its educational and economic effects on the participants. Early studies focused mainly on short-term benefits and found that participation in Head Start is associated with improved test scores and reduced grade repetition at the beginning of primary schooling. However, such benefits seem to fade out during the upper primary grades (Currie, 2001). Garces et al. (2002) provided the first empirical evidence for the longer-term effects of Head Start on high school completion, college attendance, earnings, and crime. Since then, the literature has shifted its focus to the medium- and long-term or intergenerational (e.g., Barr and Gibbs, 2022) gains of Head Start enrollment.
However, despite the enormous policy interest, evidence for the long-term effectiveness of Head Start is not unified, as summarized in Figure 1 in De Haan and Leuven (2020). The lack of consistency between these results may be due to differences in the population or problems related to the empirical approach (Elango et al., 2016). For example, the LATE obtained by the family fixed-effects approach (e.g., Garces et al., 2002; Deming, 2009; Bauer and Schanzenbach, 2016) relies on families that differ from other Head Start families in size and in other observable dimensions. Moreover, the sibling comparison design underlying that approach is limited by endogeneity concerns. To reconcile the divergent evidence, De Haan and Leuven (2020) evaluated the heterogeneous long-term effects of Head Start by using a distributional treatment effect approach that relies on two weak stochastic dominance assumptions, instead of restrictive IV assumptions. The authors found substantial heterogeneity in the returns to Head Start. Specifically, they found that the program has positive and statistically significant effects on education and wage income for the lower end of the distribution of participants.
To produce a complete picture, we assess the causal heterogeneity of Head Start from another perspective; namely, we examine the effects across different levels of unobserved resistance to participation in the program, rather than across the distribution of long-term outcomes. By relating the treatment effects to participation decision, the MTE is informative about the nature of selection into treatment and allows the computation of various causal parameters, such as the ATE, TT, and TUT. Another feature of the MTE is that its description of the effect heterogeneity is irrelevant to the specific outcome variable. Instead, the MTE curve depicts the treatment effects on the unobserved determinants of the treatment. This is another advantage of the MTE in the case of multiple outcomes of interest, as in this application, where the interpretation of the heterogeneity is kept consistent across different outcomes. Finally, in contrast to the distributional treatment effects partially identified by De Haan and Leuven (2020), our MTE method can achieve point identification and estimation.
We use the data provided by De Haan and Leuven (2020), which are from Round 16 (1994 survey year) of the National Longitudinal Study of Youth 1979 (NLSY79). The sample is restricted to the 1960–64 cohorts, because the first cohort eligible for Head Start was born in 1960. In addition, the sample excludes individuals who participated in any preschool programs other than Head Start, implying that we estimate the returns to Head Start relative to informal care. The treatment variable is whether the respondents attended the Head Start program as a child, and the outcome variables are the respondents’ highest years of education and logarithmic yearly wage incomes in their early 30s (they were between 30 and 34 years old in 1994). The covariates are age, gender, race, parental education, and family income in 1978. We refer the reader to the original paper for additional details on the data, sample, and variables. For our analysis, since we apply nonparametric estimation in the first step, we recode parental education into two categories to reduce the number of the data cells or subsamples split by different values of discrete covariates. Specifically, we redefine parental education as a binary variable equal to one if at least one parent went to college or zero if both parents are high school graduates or lower.
We first verify the credibility of Assumptions S and NL for the Head Start data. Table F.2 in Appendix F lists the support of the nonparametrically estimated propensity score for all data cells that are split by different values of discrete covariates. We find that the 34th data cell has a full support on the unit interval. Therefore, Assumption S must hold if we choose the corresponding values (32 years old, female, black race, parental education being college or higher) as the benchmark. We then verify Assumption NL for the 34th data cell. Note that only one continuous covariate exists in the data, that is, family income in 1978. Hence, we invoke Assumption NL1, which requires the propensity score function to have at least one stationary point. The left panel of Figure 1 plots the estimated propensity score as a univariate function of the continuous covariate for the 34th data cell. We find that the propensity score is highly nonlinear and nonmonotonic in family income in 1978, possibly owing to the parents’ various self-selection on participation in the program. So Assumption NL1 is fulfilled.
According to our identification result, it is enough to find one benchmark value of $X^{D}$ for which the support of the propensity score satisfies Assumption S and the propensity score function satisfies Assumption NL, as fulfilled above. In another application, however, the chances are that the propensity score function corresponding to the fully supported data cell violates Assumption NL or even there is no fully supported data cell. If this is the case, we need to try values of $X^{D}$ for which the verification of Assumption S may be not so trivial. As an example, we consider an additional data cell and retest Assumptions S and NL. We choose the largest data cell (i.e., the 29th data cell with 175 observations) and reset the benchmark to be 32 years old, male, white race, and high school or lower parental education. Now closer examination of Assumption S is necessary, as the support for this data cell is limited to $\left[0,0.194\right]$, which is totally separate from the supports for some other data cells. To this end, we need to find a distinct value of each discrete variable such that the propensity score has overlapping support with $\left[0,0.194\right]$. For example, if the value of age is altered from 32 to 30 (or 31, 33, 34) while the other discrete variables remain unaltered, we will go to the 5th (or 17th, 41st, 53rd) data cell with support $\left[0,0.177\right]$ (or $\left[0,0.242\right],\left[0,0.154\right],\left[0,0.204\right]$), which is overlapped with $\left[0,0.194\right]$ as required. Altering the values of gender, race, and parental education leads to the 35th, 25th (or 27th), and 30th data cells with supports $\left[0,0.526\right]$, $\left[0,0.771\right]$ (or $\left[0,0.623\right]$), and $\left[0,0.061\right]$, respectively, which all meet the overlapping condition as well. Therefore, Assumption S holds for the largest data cell. The validation of Assumption NL1 for this data cell can be demonstrated as well by the existence of stationary points of the estimated propensity score function, as illustrated in the right panel of Figure 1.
We next turn to the estimation. In the first step, we nonparametrically regress the treatment variable (Head Start) on all of the covariates to generate propensity scores for the sample, by employing the kernel estimation method in (12) with the Gaussian kernel and the rule-of-thumb bandwidth as given in Table F.1. Figure 2 plots the frequency distribution of the estimated propensity score by treatment status. The figure shows that the propensity score in our sample follows a bimodal distribution, with the main peak being at approximately 0.5 for the participants and approximately 0.1 for the nonparticipants. To reduce the potential impact of outliers, we trim the observations of the 1% smallest and 1% largest propensity scores in the following steps. This trimming leads to a common support ranging from 0 to 0.6, as indicated by the two dashed vertical lines. Given the estimated propensity score, we then estimate in sequence the linear coefficients $\beta_{1}$ and $\beta_{0}$, MTE, and summary treatment effect measures including ATE, TT, and TUT. Since the separate estimation procedure exploits more identifying information behind the data than the LIV procedure in general (Brinch et al., 2017), we focus on the results of the separate estimation.
Figure 3 shows the estimated MTE under Heckman’s normal specification in which $E\left[\left.U_{d}\right|V=v\right]=\rho_{d0}+\rho_{dV}\Phi^{-1}\left(v\right)$ for $d=0,1$, where $\Phi\left(\cdot\right)$ is the CDF of standard normal distribution, which can be derived by assuming that $\left(U_{d},U\right)$ follows a bivariate normal distribution as in Example 3. The MTE curves, evaluated at the mean values of $X$, relate the unobserved component $U_{1}-U_{0}$ of the treatment effect to the unobserved component $V$ of the treatment choice. A high value of $V$ implies a low probability of treatment; thus, we interpret $V$ as resistance to participation in Head Start. The MTE curve for wage income plotted in panel B decreases with resistance, revealing a pattern of selection on gains, as expected. In other words, based on the unobserved characteristics, the children who were most likely to enroll in Head Start benefitted the most from the program in terms of their labor market outcome. However, when the outcome is educational attainment in panel A, the positive slope of the MTE curve points to a pattern of reverse selection on gains. In consequence, the TUT exceeds the ATE, which in turn exceeds the TT. The same pattern was observed by Cornelissen et al. (2018) when estimating school-readiness return to a preschool program in Germany. This phenomenon may be attributed partly to the fact that parents have their own objectives in deciding childcare arrangements. Nevertheless, the disagreement between selection patterns for the two outcomes raises concern about the possible functional form misspecification of the normal MTE, which is strictly restricted to be monotonic in the resistance to treatment. Therefore, we consider a nonparametric specification for $E\left[\left.U_{d}\right|V=v\right]$ and implement a semiparametric separate estimation procedure.
Figure 4 plots the semiparametric MTE curves for education and labor market outcomes in panels A and B, respectively. Under the flexible specification, the MTE curves are no longer monotonic, and the clear pattern of selection disappears. In the case of education outcome, the curve is initially flat, then becomes an inverted U shape, with a statistically significant positive effect appearing in the region of the peak, corresponding to the children with resistance to treatment ranging from 0.32 to 0.42. The complex feature of this curve is hardly captured by any parsimonious parametric function. Similar observations are seen in the case of wage income, in which the MTE curve is nonmonotonic, with a complex shape, and significantly greater than zero for less than 10% of the children who were most likely to attend childcare early. A comparison of the summary treatment effect measures indicates weak selection on gains for both outcomes. Table 1 reports the semiparametric estimates for the effects of the covariates on potential outcomes and their difference based on (13). Columns 3 and 6 show that girls gain significantly more returns to Head Start attendance than boys. However, other than gender, no substantial observable heterogeneity exists in the treatment effects of the program, though parental education and family income in 1978 have a significantly positive effect on the respondents’ potential education and potential labor market outcomes in both the treated and untreated states.
6 Conclusion
We propose a novel method for defining, identifying, and estimating the MTE in the absence of IVs. Our MTE model allows all the covariates to be correlated to the structural treatment error. In this model, we define the MTE based on a reduced-form treatment error that (i) is uniformly distributed on the unit interval, (ii) is statistically independent of covariates, and (iii) has several economic meanings. The independence property facilitates the identification of our defined MTE. We provide sufficient conditions under which the proposed IV-free MTE can be point identified based on functional forms. The conditions are standard in a certain sense. The conditional mean independence assumption is equivalent to the separability assumption commonly imposed in the MTE literature, and is implied by and much weaker than the full independence assumption. The linearity and nonlinearity assumptions are the foundation of identification based on functional forms, and can make sense in most empirical studies. We prove the identification by using a construction method. Our identification strategy allows the adaptation of most of the existing estimation procedures for the IV-based MTE, such as separate estimation, LIV estimation, parametric estimation, and semiparametric estimation. For the empirical application, we evaluate the MTE of the Head Start program on long-term education and labor market outcomes, in which an IV for Head Start participation is difficult to acquire. We find significant positive effects for individuals with medium-level or low resistance to treatment, and substantial heterogeneity exists.
Acknowledgements
We wish to thank the Editor and two anonymous reviewers for their valuable comments and suggestions. Financial supports for this research are provided through the National Social Science Foundation of China (Grant 24FTJB010) and the National Natural Science Foundation of China (Grants 72173142, 72034006, 72342034, 72173083).
Appendix A Additional results of the empirical application
Based on $\hat{\beta}_{1}$ and $\hat{\beta}_{0}$ reported in Table 1, and the separate estimates of $E\left[\left.U_{d}\right|V=v\right]$ for $d=0,1$ under the semiparametric specification, we can estimate the marginal structural functions
$E\left[\left.Y_{d}\right|V=v\right]=E\left[X\right]^{\prime}\beta_{d}+E\left[\left.U_{d}\right|V=v\right]$
for potential outcomes $Y_{1}$ and $Y_{0}$, which we plot in Figure A.1. Panel A sheds light on the significantly positive effect of Head Start on education for the respondents with medium-level resistance to treatment, revealing that their gains from the program are driven mainly by the remarkably low educational attainment when untreated. Panel B leads to similar results for wage income, where significant gain emerges for low values of $V$. Moreover, the relatively flat curve of potential wage income in the treated state implies that early childcare attendance serves as an equalizer that diminishes the intergroup difference in the labor market outcome.
Figure A.1: Semiparametric estimates of marginal structural functions
Notes: The marginal structural function refers to the expected potential outcome conditional on the unobserved resistance to treatment. Panels A and B depict the estimated marginal structural function curves for education outcome and labor market outcome, respectively. The 90% confidence interval is based on bootstrapping with 1,000 replications.
In addition, we can semiparametrically estimate conditional expectations of the observed responses given the unobserved determinants of treatment choice, such as the marginal probability of participation
$E\left[\left.D\right|V=v\right]=\Pr\left(\left.P\geq v\right|V=v\right)=\Pr\left(P\geq v\right),$
which mirrors the distribution of the propensity score, and the marginal observed outcome
$E\left[\left.Y\right|V=v\right]$ $=$ $E\left[\left.Y_{0}\right|V=v\right]+E\left[\left.D\left(Y_{1}-Y_{0}\right)\right|V=v\right]$
$=$ $E\left[\left.Y_{0}\right|V=v\right]+E\left[\left.DX\right|V=v\right]^{\prime}\left(\beta_{1}-\beta_{0}\right)+E\left[\left.D\left(U_{1}-U_{0}\right)\right|V=v\right]$
$=$ $E\left[\left.Y_{0}\right|V=v\right]+E\left[1\left\{P\geq v\right\}X\right]^{\prime}\left(\beta_{1}-\beta_{0}\right)+\Pr\left(P\geq v\right)\cdot E\left[\left.U_{1}-U_{0}\right|V=v\right],$
where the last equality follows from the independence of $X$ and $V$ and from the conditional mean independence of $X$ and $U_{d}$ given $V$. Plugging in proper estimates of each component in the above equations, we obtain the estimated marginal response functions and plot them in Figure A.2. In particular, the marginal observed outcome curves relate the individuals with significant positive returns (with medium $V$ in Panel A and small $V$ in Panel B) to those with low education and low wage income, thereby bridging our results on the MTE and those of De Haan and Leuven (2020) on the distributional treatment effect.
Figure A.2: Semiparametric estimates of marginal response functions
Notes: This figure plots the (conditional) expectation of the observed outcome (solid line on the left y-axis), and the conditional probability of participation (dashed line on the right y-axis). Panels A and B illustrate the estimated marginal response function curves for education outcome and labor market outcome, respectively. The 90% confidence interval is based on bootstrapping with 1,000 replications.
Appendix B Proof of Theorem 1
[…]
$m_{d0}\left(x^{C}\right)=x^{C\prime}\beta_{d}^{C}+g_{d}\left(\pi_{0}\left(x^{C}\right)\right).$ (B.1)
When $\dim\left(X^{C}\right)=1$, by differentiating both sides of (B.1) at $\tilde{x}^{C}$ and invoking Assumption NL1, we obtain
$m_{d0}^{\left(1\right)}\left(\tilde{x}^{C}\right)=\beta_{d}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\tilde{x}^{C}\right)\right)\pi_{0}^{\left(1\right)}\left(\tilde{x}^{C}\right)=\beta_{d}^{C},$
which identifies $\beta_{d}^{C}$ for $d=0,1$. When $\dim\left(X^{C}\right)\geq 2$, for $k,j\in\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ satisfying Assumption NL2, taking the partial derivatives of $m_{d0}\left(x^{C}\right)$ with respect to $x_{k}^{C}$ and $x_{j}^{C}$ yields that
$\partial_{k}m_{d0}\left(x^{C}\right)$ $=$ $\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)\partial_{k}\pi_{0}\left(x^{C}\right),$
$\partial_{j}m_{d0}\left(x^{C}\right)$ $=$ $\beta_{d,j}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)\partial_{j}\pi_{0}\left(x^{C}\right).$
It follows from Assumption NL2.(i)-(ii) that
$\frac{\partial_{k}m_{d0}\left(x^{C}\right)-\beta_{d,k}^{C}}{\partial_{k}\pi_{0}\left(x^{C}\right)}=g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)=\frac{\partial_{j}m_{d0}\left(x^{C}\right)-\beta_{d,j}^{C}}{\partial_{j}\pi_{0}\left(x^{C}\right)},$
so that
$\partial_{k}m_{d0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)-\partial_{j}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)=\partial_{j}\pi_{0}\left(x^{C}\right)\beta_{d,k}^{C}-\partial_{k}\pi_{0}\left(x^{C}\right)\beta_{d,j}^{C},$ (B.2)
which is linear in $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. The same equation is obtained if we evaluate the expression at another point $\tilde{x}^{C}$ that satisfies Assumption NL2.(iii)-(iv), which gives
$\left(\begin{array}[]{c}\partial_{k}m_{d0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)-\partial_{j}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)\\ \partial_{k}m_{d0}\left(\tilde{x}^{C}\right)\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)-\partial_{j}m_{d0}\left(\tilde{x}^{C}\right)\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\end{array}\right)=\Psi\left(\begin{array}[]{c}\beta_{d,k}^{C}\\ \beta_{d,j}^{C}\end{array}\right),$ (B.3)
where
$\Psi=\left(\begin{array}[]{cc}\partial_{j}\pi_{0}\left(x^{C}\right),&-\partial_{k}\pi_{0}\left(x^{C}\right)\\ \partial_{j}\pi_{0}\left(\tilde{x}^{C}\right),&-\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\end{array}\right).$
Assumption NL2.(v) ensures that the determinant of $\Psi$ is nonzero, which shows that $\Psi$ is nonsingular. Therefore, equation (B.3) can be solved for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$ by inverting $\Psi$, thereby identifying $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Given identification of $\beta_{d,k}^{C}$, we can then identify all other coefficient $\beta_{d,l}^{C}$ in $\beta_{d}^{C}$ by solving (B.2) with the subscript $j$ replaced by $l$, which gives
$\beta_{d,l}^{C}=\frac{\partial_{l}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)-\partial_{k}m_{d0}\left(x^{C}\right)\partial_{l}\pi_{0}\left(x^{C}\right)+\partial_{l}\pi_{0}\left(x^{C}\right)\beta_{d,k}^{C}}{\partial_{k}\pi_{0}\left(x^{C}\right)}.$
Given the identification of $\beta_{d}^{C}$, the function $g_{d}$ is identified on the support of $\pi_{0}\left(X^{C}\right)$ by
$g_{d}\left(p\right)=E\left[\left.m_{d0}\left(X^{C}\right)-X^{C\prime}\beta_{d}^{C}\right|\pi_{0}\left(X^{C}\right)=p\right].$
Next, we consider identifying $\beta_{d}^{D}$, the coefficients of discrete covariates. For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, we have
$m_{d}\left(x^{C},x^{Dk}\right)=x^{C\prime}\beta_{d}^{C}+x_{k}^{D}\beta_{d,k}^{D}+g_{d}\left(\pi\left(x^{C},x^{Dk}\right)\right)$
for any $x^{C}$ in the support of $X^{C}$. By Assumption S, there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$. It follows from the above identification result that $g_{d}\left(\pi\left(x^{C}\left(k\right),x^{Dk}\right)\right)$ is identified. Consequently, $\beta_{d,k}^{D}$ is identified by
$\beta_{d,k}^{D}=\frac{m_{d}\left(x^{C}\left(k\right),x^{Dk}\right)-x^{C}\left(k\right)^{\prime}\beta_{d}^{C}-g_{d}\left(\pi\left(x^{C}\left(k\right),x^{Dk}\right)\right)}{x_{k}^{D}}.$
This argument holds for each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, thereby identifying $\beta_{d}^{D}$. Finally, given the identification of $\beta_{d}=\left(\beta_{d}^{C},\beta_{d}^{D}\right)$, it follows from (8) that the function $g_{d}$ is identified on the support of $P=\pi\left(X\right)$ by
$g_{d}\left(p\right)=E\left[\left.m_{d}\left(X\right)-X^{\prime}\beta_{d}\right|\pi\left(X\right)=p\right],$
for $d=0,1$.
Appendix C Further discussion on the nonlinearity assumption
In light of the proof of Theorem 1, our identification strategy may work even when no continuous covariate is available in the data. In that case, we put a discretely distributed covariate in $X^{C}$.
Theorem C.1.
Theorem 1 holds if Assumption NL1 is replaced by that there exist two different constants $x^{C}$, $\tilde{x}^{C}$ in the support of $X^{C}$ such that $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$.
Proof.
When $\dim\left(X^{C}\right)=1$, it follows from (B.1) and the assumption of Theorem C.1 that
$m_{d0}\left(x^{C}\right)-x^{C}\beta_{d}^{C}=g_{d}\left(\pi_{0}\left(x^{C}\right)\right)=g_{d}\left(\pi_{0}\left(\tilde{x}^{C}\right)\right)=m_{d0}\left(\tilde{x}^{C}\right)-\tilde{x}^{C}\beta_{d}^{C}.$
Hence, $\beta_{d}^{C}$ is identified as
$\beta_{d}^{C}=\frac{m_{d0}\left(x^{C}\right)-m_{d0}\left(\tilde{x}^{C}\right)}{x^{C}-\tilde{x}^{C}}.$
The remaining part of the proof is the same as the proof of Theorem 1. ∎
The assumption of Theorem C.1 in place of NL1 requires the univariate function $\pi_{0}$ to be not one-to-one, but imposes no other smoothness or continuity restriction on $\pi_{0}$. Therefore, it doesn’t require $X^{C}$ to be continuously valued. In the extreme case of $X^{C}$ containing only one binary covariate, this assumption will be equivalent to the full irrelevance of $X^{C}$ to the treatment probability, which is a condition suggested by Chamberlain (1986) for identification of the sample selection model. The following two theorems show that Assumption NL1 and the assumption of Theorem C.1 can be extended to the case of multiple continuous covariates.
Theorem C.2.
Theorem 1 holds if Assumption NL is replaced by that (i) there are two points $x^{C}$ and $\tilde{x}^{C}$ in the support of $X^{C}$ and an index $k$ such that $x_{k}^{C}\neq\tilde{x}_{k}^{C}$, $x_{-k}^{C}=\tilde{x}_{-k}^{C}$, and $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$, and (ii) there exists $\breve{x}^{C}$ in the support of $X^{C}$ such that functions $\pi_{0}$ and $m_{d0}$ are differentiable at $\breve{x}^{C}$, with $\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)\neq 0$.
Proof.
It follows from equation (B.1) and $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$ that
$m_{d0}\left(x^{C}\right)-x_{k}^{C}\beta_{d,k}^{C}-x_{-k}^{C\prime}\beta_{d,-k}^{C}=m_{d0}\left(\tilde{x}^{C}\right)-\tilde{x}_{k}^{C}\beta_{d,k}^{C}-\tilde{x}_{-k}^{C^{\prime}}\beta_{d,-k}^{C}.$
Hence, $\beta_{d,k}^{C}$ is identified by condition (i) as
$\beta_{d,k}^{C}=\frac{m_{d0}\left(x^{C}\right)-m_{d0}\left(\tilde{x}^{C}\right)}{x_{k}^{C}-\tilde{x}_{k}^{C}}.$
For any index $j\neq k$, taking the partial derivatives of $m_{d0}\left(x^{C}\right)$ with respect to $x_{k}^{C}$ and $x_{j}^{C}$ at $\breve{x}^{C}$ yields that
$\partial_{k}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{k}\pi_{0}\left(\breve{x}^{C}\right),$ (C.1)
$\partial_{j}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,j}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{j}\pi_{0}\left(\breve{x}^{C}\right).$ (C.2)
Since $\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)\neq 0$ by condition (ii), we can find $g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)$ from (C.1) to be
$g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)=\frac{\partial_{k}m_{d0}\left(\breve{x}^{C}\right)-\beta_{d,k}^{C}}{\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)},$
and substitute it into (C.2) to identify $\beta_{d,j}^{C}$ as
$\beta_{d,j}^{C}=\partial_{j}m_{d0}\left(\breve{x}^{C}\right)-\frac{\partial_{j}\pi_{0}\left(\breve{x}^{C}\right)}{\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)}\left[\partial_{k}m_{d0}\left(\breve{x}^{C}\right)-\beta_{d,k}^{C}\right].$
Therefore, $\beta_{d}^{C}$ is identified. The remaining part of the proof is the same as the proof of Theorem 1. ∎
Theorem C.3.
Theorem 1 holds if Assumption NL is replaced by that there exist two points $x^{C}$ and $\tilde{x}^{C}$ in the support of $X^{C}$ and an index $k$ such that functions $\pi_{0}$ and $m_{d0}$ are differentiable at $x^{C}$ and $\tilde{x}^{C}$, with $\partial_{k}\pi_{0}\left(x^{C}\right)=0$ and $\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$.
Proof.
The identification of $\beta_{d,k}^{C}$ is analogous to Theorem 1. Then, the identification of all the other coefficients $\beta_{d,j}^{C}$ in $\beta_{d}^{C}$ and the remaining components of the model is the same as in Theorem C.2 and Theorem 1, respectively. ∎
Alternatively, the identification of coefficients of continuous covariates can be attained by exploiting the local irrelevance of the function $g_{d}\left(p\right)$.
Theorem C.4.
Theorem 1 holds if Assumption NL is replaced by the existence of a vector $x^{C}$ in the support of $X^{C}$ such that $g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)=0$ for $d=0,1$, and that functions $\pi_{0}$ and $m_{d0}$ are differentiable at $x^{C}$.
Proof.
For each $k\in\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$, taking the partial derivative of both sides of (B.1) with respect to $x_{k}^{C}$ gives
$\partial_{k}m_{d0}\left(x^{C}\right)=\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)\partial_{k}\pi_{0}\left(x^{C}\right)=\beta_{d,k}^{C},$
which immediately identifies all $\beta_{d,k}^{C}$ in $\beta_{d}^{C}$. The remaining part of the proof is the same as that of Theorem 1. ∎
The above discussion further illustrates the over-identification feature of our strategy. In fact, the strength of identification of our method is determined by the number of nonlinearity points in the data satisfying Assumption NL or Theorems C.1-C.4, given different benchmark values of $X^{D}$.
Appendix D Limited valued outcomes
To accommodate limited valued outcomes, we first generalize the linear additive model to a linear latent index model.
Assumption D.L (Linear index). Assume that $Y_{d}=l\left(X^{\prime}\gamma_{d},V_{d}\right)$, $d=0,1$, holds for some known link function $l$ and unknown coefficients $\gamma_{d}$, where $V_{d}$ is unobserved error term of $Y_{d}$.
The generalized linear model applies to a variety of frequently encountered limited dependent variables, such as $l\left(t,v\right)=1\left\{t\geq v\right\}$ for binary $Y\in\left\{0,1\right\}$, $l\left(t,v\right)=\max\left(t-v,0\right)$ for censored (at zero) $Y\geq 0$, $l\left(t,v\right)=\exp\left(t+v\right)\left/\left[1+\exp\left(t+v\right)\right]\right$. for proportion-valued $Y\in\left(0,1\right)$, $l\left(t,v\right)=\sum_{j=1}^{s}1\left\{t+v\geq c_{j}\right\}$ for ordered $Y\in\left\{0,1,\cdots,s\right\}$ with known thresholds $c_{1}<c_{2}<\cdots<c_{s}$, and so forth. Although setting $l\left(t,v\right)=t+v$ reduces to the linear additive model, the following discussion doesn’t include Theorem 1 as a special case because $\gamma_{d}$ here can be only identified up to scale for a general link function $l$. Any changes in the scaling of $\gamma_{d}$ can be freely absorbed into $l$, and a scale normalization is needed for identification of $\gamma_{d}$ and $l$.
Assumption D.N (Normalization). Decompose $X^{\prime}\gamma_{d}=X^{C\prime}\gamma_{d}^{C}+X^{D\prime}\gamma_{d}^{D}$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. Assume that $\gamma_{d,1}^{C}$, the first element of $\gamma_{d}^{C}$, equal to 1.
Assumption D.N imposes the convenient normalization that the first continuous covariate has a unit coefficient. This scaling of $\gamma_{d}$ is arbitrary and is innocuous because our focus is on the identification of MTE, which is a difference in the conditional expectations of $l\left(X^{\prime}\gamma_{d},V_{d}\right)$, rather than on separate identification of $\gamma_{d}$ and $l$. Since in the generalized linear model the potential outcomes are not necessarily additive in the error term, the mean independence assumption needs to be strengthened to a stricter distributional independence assumption.
Assumption D.CI (Conditional Independence). Assume that $\left.V_{d}\perp\!\!\!\!\perp X\right|V$ for $d=0,1$, namely, $V_{d}$ is independent of $X$ conditional on $V$, where $V$ is the reduced-form treatment error in equation (3).
Recalling that $V\perp\!\!\!\!\perp X$ by definition, Assumption D.CI is equivalent to the full independence $\left(V_{d},V\right)\perp\!\!\!\!\perp X$ for $d=0,1$, which implies that both the marginal distribution of $V_{d}$ and the copula of $V_{d}$ and $V$ are independent of $X$. Under Assumptions D.L and D.CI, we have
$\Delta^{\text{MTE}}\left(x,v\right)=E\left[\left.l\left(x^{\prime}\gamma_{1},V_{1}\right)-l\left(x^{\prime}\gamma_{0},V_{0}\right)\right|V=v\right].$
The additive nonseparability of the observables and unobservables constitutes the primary difficulty in identifying MTE in the limited outcome case. Meanwhile, Assumptions D.L and D.CI lead to a double index form of the observable outcome regression functions for each treatment status, that is,
$E\left[\left.Y\cdot 1\left\{D=d\right\}\right|X=x\right]=r_{d}\left(x^{\prime}\gamma_{d},\pi\left(x\right)\right)$ (D.1)
for $d=0,1$, where
$r_{0}\left(t,p\right)$ $=$ $\int_{p}^{1}E\left[\left.l\left(t,V_{0}\right)\right|V=v\right]dv,$ (D.2)
$r_{1}\left(t,p\right)$ $=$ $\int_{0}^{p}E\left[\left.l\left(t,V_{1}\right)\right|V=v\right]dv.$ (D.3)
Provided that $\pi\left(x\right)$ is a nonlinear index, $\gamma_{d}$ and thus $r_{d}$ can be identified based on functional forms, which ensures identification of MTE since
$\Delta^{\text{MTE}}\left(x,v\right)=\partial_{2}r_{1}\left(x^{\prime}\gamma_{1},v\right)+\partial_{2}r_{0}\left(x^{\prime}\gamma_{0},v\right),$ (D.4)
where $\partial_{2}r_{d}$ is the partial derivative of $r_{d}$ with respect to its second argument. Like in the case of unlimited outcomes, the key powers of this identification strategy are supplied by the nonlinearity of $\pi\left(x\right)$, as specified by the following assumption.
Assumption D.NL (Non-Linearity). Assume that the functions $\pi_{0}$ and $r_{d}$ are differentiable and denote their partial derivatives with respect to the $k$-th argument as $\partial_{k}\pi_{0}$ and $\partial_{k}r_{d}$, where $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ and $r_{d}$, $d=0,1$, are defined in (D.1) and (D.1). Assume that there exist two vectors $x^{C}$ and $\tilde{x}^{C}$ on the support of $X^{C}$ and two elements $k$ and $j$ of the set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{1}r_{d}\left(x^{C\prime}\gamma_{d}^{C},\pi_{0}\left(x^{C}\right)\right)\neq 0$, (ii) $\partial_{1}r_{d}\left(\tilde{x}^{C\prime}\gamma_{d}^{C},\pi_{0}\left(\tilde{x}^{C}\right)\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,k}^{C}$, (iv) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,j}^{C}$, (v) $\partial_{l}\pi_{0}\left(\tilde{x}^{C}\right)\neq\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\gamma_{d,l}^{C}$ for $l=2,\cdots,\dim\left(X^{C}\right)$, and
(vi) $\partial_{k}\pi_{0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)-\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)$
$\neq\left[\partial_{1}\pi_{0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)-\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)\right]\gamma_{d,k}^{C}$
$-\left[\partial_{1}\pi_{0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)-\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)\right]\gamma_{d,j}^{C}.$
Assumption D.NL.(i)-(ii) require $r_{d}$ to depend on the linear index, and (iii)-(vi) essentially require some nonlinear variation in $\pi_{0}$ under the scale normalization. As in Assumption NL, it is difficult to construct examples other than $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ that violates Assumption D.NL. Finally, a support assumption and an invertibility assumption are imposed for technical reasons.
Assumption D.S (Support). For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, assume for some $x_{k}^{D}\neq 0$ in the support of $X_{k}^{D}$ that there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$ and that $x^{C}\left(k\right)^{\prime}\gamma_{d}^{C}+x_{k}^{D}\gamma_{d,k}^{D}$ is in the support of $X^{C\prime}\gamma_{d}^{C}$ for $d=0,1$.
Assumption D.I (Invertibility). Assume that the function $r_{d}$ is invertible on its first argument for $d=0,1$.
For the case of a binary outcome, we have $r_{0}\left(t,p\right)=\int_{p}^{1}F_{\left.V_{0}\right|V}\left(\left.t\right|v\right)dv$ and $r_{1}\left(t,p\right)=\int_{0}^{p}F_{\left.V_{1}\right|V}\left(\left.t\right|v\right)dv$, hence a sufficient condition for Assumption D.I to hold is that $V_{d}$ is continuously distributed with support $\mathbb{R}$, conditional on $V$, for $d=0,1$. Interestingly, the same continuous distribution condition also suffices for Assumption D.I in the censored case with $l\left(t,v\right)=\max\left(t-v,0\right)$, where $\partial_{1}r_{0}\left(t,p\right)=\int_{p}^{1}F_{\left.V_{0}\right|V}\left(\left.t\right|v\right)dv$ and $\partial_{1}r_{1}\left(t,p\right)=\int_{0}^{p}F_{\left.V_{1}\right|V}\left(\left.t\right|v\right)dv$. Under the imposed assumptions, the following identification theorem for the generalized linear model follows immediately from Escanciano et al. (2016, Theorems 3.1-3.2).
Theorem D.1.
If Assumptions D.L, D.NL, D.CI, D.S, D.N, and D.I hold, then $\gamma_{d}$ and $r_{d}\left(t,p\right)$ at all points in the support of $X^{\prime}\gamma_{d}$ and $P$ are identified for $d=0,1$.
Having established identification of $\gamma_{d}$ and $r_{d}$, the MTE for limited valued outcomes can be identified by (D.4) and estimated by the semiparametric least squares method of Escanciano et al. (2016). Alternatively, the LIV estimation procedure may be adapted according to
$E\left[\left.Y\right|X=x\right]=s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},\pi\left(x\right)\right)$
and
$\Delta^{\text{MTE}}\left(x,v\right)$ $=$ $\frac{\partial s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v\right)}{\partial v},$
$\Delta^{\text{ATE}}\left(x\right)$ $=$ $s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},1\right)-s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},0\right),$
$\Delta^{\text{TT}}\left(x\right)$ $=$ $\frac{s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},\pi\left(x\right)\right)-s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},0\right)}{\pi\left(x\right)},$
$\Delta^{\text{TUT}}\left(x\right)$ $=$ $\frac{s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},1\right)-s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},\pi\left(x\right)\right)}{1-\pi\left(x\right)},$
$\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $\frac{s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v_{2}\right)-s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v_{1}\right)}{v_{2}-v_{1}},$
where
$s\left(t_{0},t_{1},p\right)=r_{0}\left(t_{0},p\right)+r_{1}\left(t_{1},p\right).$
Appendix E Sufficient conditions and proofs for Theorem 2
E.1 Sufficient conditions
We first give the assumptions that have been mentioned when presenting Theorem 2. We denote $f_{X}\left(\cdot\right)$ and $f_{P}\left(\cdot\right)$ as the probability density function (PDF) of $X=\left(X^{C},X^{D}\right)$ and $P=\pi\left(X\right)$, respectively, and $f_{P}^{\left(1\right)}\left(p\right)=\left.\partial f_{P}\left(p\right)\right/\partial p$ as the derivative function of $f_{P}\left(\cdot\right)$ if it is differentiable. For $d=0,1$, we define $V_{d}=Y-X^{\prime}\beta_{d}-g_{d}\left(P\right)$, which satisfies $E\left[\left.V_{d}\right|X,D=d\right]=0$ because by definition $g_{d}\left(p\right)=E\left[\left.Y-X^{\prime}\beta_{d}\right|\pi\left(X\right)=p,D=d\right]$. And we denote $\sigma_{V_{d}}^{2}\left(p\right)=Var\left(\left.V_{d}\right|P=p,D=d\right)=E\left[\left.V_{d}^{2}\right|P=p,D=d\right]$. Note that
$\sigma_{V_{d}}^{2}\left(p\right)$ $=$ $E\left[\left.\left(U_{d}-g_{d}\left(P\right)\right)^{2}\right|P=p,D=d\right]$
$=$ $E\left[\left.U_{d}^{2}\right|P=p,D=d\right]-g_{d}\left(p\right)$
$=$ $\sigma_{U_{d}}^{2}\left(p\right)-g_{d}\left(p\right).$
In addition, we denote
$\delta_{d}\left(p\right)=E\left[\left.1\left\{D=d\right\}\right|P=p\right]=\left\{\begin{array}[]{cc}p,&\text{if }d=1,\\ 1-p,&\text{if }d=0,\end{array}\right$.
and
$\delta_{d}^{\left(1\right)}\left(p\right)=\frac{\partial\delta_{d}\left(p\right)}{\partial p}=\left\{\begin{array}[]{cc}1,&\text{if }d=1,\\ -1,&\text{if }d=0.\end{array}\right$.
Assumption E.1 (First step). For an integer $s\geq 2$, we assume that:
(i) $\pi\left(x^{C},x^{D}\right)$ and $f_{X}\left(x^{C},x^{D}\right)$ are $s$-times differentiable with respect to $x^{C}$ for all $x^{D}$, and the derivative functions are all Lipschitz continuous and bounded;
(ii) $\inf_{x^{C}}f_{X}\left(x^{C},x^{D}\right)\geq c$ for a positive constant $c$ for all $x^{D}$;
(iii) $k_{1}\left(\cdot\right)$ is a bounded $s$-th order kernel function satisfying that $\int k_{1}\left(u\right)du=1$, $\int u^{t}k_{1}\left(u\right)du=0$ for $t=1,\cdots,s-1$, and $\int u^{s}k_{1}\left(u\right)du\neq 0$, and the functions $K_{1l}\left(u\right)=\left|u\right|^{l}k_{1}\left(u\right)$ are Lipschitz continuous for all $0\leq l\leq s+1$.
(iv) $h_{1l}\rightarrow 0$ for $l=1,2,\cdots,\dim\left(X^{C}\right)$, and $\left.nh_{11}h_{12}\cdots h_{1,\dim\left(X^{C}\right)}\right/\ln n\rightarrow\infty$, as $n\rightarrow\infty$.
Assumption E.2 (Second step). For $d=0,1$, we assume that:
(i) $\left\|\hat{\beta}_{d}-\beta_{d}\right\|=O_{p}\left(n^{-1/2}\right)$;
(ii) $g_{d}\left(\cdot\right)$, $f_{P}\left(\cdot\right)$, and $\sigma_{V_{d}}^{2}\left(\cdot\right)$ are twice continuously differentiable over the support of $P$, and $\inf_{p\in\left[0,1\right]}f_{P}\left(p\right)\geq c$ for a positive constant $c$.
(iii) $k_{3}\left(\cdot\right)$ is a bounded, symmetric, compactly supported, and three times continuously differentiable kernel function;
(iv) $h_{3}\rightarrow 0$, $nh_{3}^{3}\rightarrow\infty$, and $nh_{3}^{7}\rightarrow 0$, as $n\rightarrow\infty$.
Assumption E.3 (Rate). We assume that as $n\rightarrow\infty$, (i) $nh_{3}\eta_{1}^{2}\rightarrow 0$, (ii) $\left.n\eta_{1}^{6}\right/h_{3}^{7}\rightarrow 0$, (iii) $nh_{3}^{3}\left|h_{1}\right|\rightarrow\infty$, and (iv) $n^{3/2}h_{3}^{5}\left|h_{1}\right|^{2}\rightarrow\infty$, where
$\eta_{1}$ $=$ $tr\left(h_{1}^{s}\right)+\sqrt{\frac{\ln n}{n\left|h_{1}\right|}},$
$tr\left(h_{1}^{s}\right)$ $=$ $\sum_{l=1}^{\dim\left(X^{C}\right)}h_{1l}^{s},$
$\left|h_{1}\right|$ $=$ $h_{11}h_{12}\cdots h_{1,\dim\left(X^{C}\right)}.$
Assumption E.4 (Moment). We assume that $E\left[\left\|X\right\|^{2}\right]<\infty$ and $E\left[V_{d}^{2}\right]<\infty$ for $d=0,1$.
Assumption E.1.(iii) employs the higher-order kernel device to reduce the bias and improve the convergence rate of the first-step estimation (12). Since the first step is a multivariate kernel regression, the usual symmetric (second-order) kernel will necessarily lead to a rate of convergence that dominates the subsequent local linear and MTE estimation, which is apparently undesirable. By using the higher-order kernel, Lemma E.1 below implies that
$\sup_{x}\left|\hat{\pi}\left(x\right)-\pi\left(x\right)\right|=O_{p}\left(\eta_{1}\right).$
Therefore, Assumption E.3.(i) ensures the asymptotic negligibility of the first-step estimation error relative to the subsequent local linear estimation. If $h_{1l}$ follows a rate $O\left(n^{-1\left/\left(2s+\dim\left(X^{C}\right)\right)\right.}\right)$ that optimally balances the squared bias and the variance, the first-step rate of convergence will become $O_{p}\left(\left(\ln n\right)^{1/2}n^{-s\left/\left(2s+\dim\left(X^{C}\right)\right)\right.}\right)$. Further suppose that $h_{3}$ follows a rule-of-thumb rate $O\left(n^{-1/5}\right)$, then Assumption E.3.(i) will reduce to $s>2\dim\left(X^{C}\right)$. Note that under the supposed rates of $h_{1l}$ and $h_{3}$, Assumption E.3.(ii), (iii), and (iv) will reduce to $s>2\dim\left(X^{C}\right)$, $s>\left(3/4\right)\dim\left(X^{C}\right)$, and $s>\left(3/2\right)\dim\left(X^{C}\right)$, respectively, which are all implied by Assumption E.3.(i).
Assumption E.2.(i) imposes a weak rate condition on the slope estimation in the second step. The $\sqrt{n}$-consistency of the weighted pairwise difference estimation (13) has been established by Ahn and Powell (1993, Corollary 3.1). Actually, most semiparametric estimators for sample selection models satisfy Assumption E.2.(i) as well, such as the Robinson-type partialling-out estimator (Pagan and Ullah, 1999, p.307), the symmetry-based pairwise difference estimator (Chen and Zhou, 2010), and the integrated least squares estimator (Pan et al., 2022), among others.
E.2 Proof of Theorem 2
We first develop the asymptotic normality of the local linear estimator $\left(\hat{g}_{d}\left(p\right),\hat{g}_{d}^{\left(1\right)}\left(p\right)\right)$ for $d=0,1$. To this end, we define the infeasible estimator as
$\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)=\left[\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)^{\prime}\right]^{-1}\left[\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\beta_{d}\right)\right],$
where
$w_{di}\left(p\right)=1\left\{D_{i}=d\right\}\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right).$
We will analyze $\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$ and $\left(\hat{g}_{d}\left(p\right),\hat{g}_{d}^{\left(1\right)}\left(p\right)\right)-\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$ in order.
The asymptotic normality of $\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$ can be established by standard arguments for kernel regression (e.g., Li and Racine, 2007, Subsection 2.7.2), except for the presence of the treatment status indicator $1\left\{D_{i}=d\right\}$ in our case. For completeness, we provide a proof for the asymptotic normality of $\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$ in Lemma E.2. In the proof, we obtain an asymptotically linear representation of $\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$ in (E.52):
$J_{n}\left[\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)\right]=\Gamma_{d}J_{n}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)V_{di}+o_{p}\left(1\right),$ (E.1)
where
[…]
$J_{n}$ $=$ $\left(\begin{array}[]{cc}\sqrt{nh_{3}}&0\\ 0&\sqrt{nh_{3}^{3}}\end{array}\right).$
We next consider $J_{n}\left[\left(\hat{g}_{d}\left(p\right),\hat{g}_{d}^{\left(1\right)}\left(p\right)\right)-\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)\right]$.
For $r=0,1,2$, we denote
$\hat{B}_{r}$ $=$ $\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)^{r},$
$B_{r}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{r},$
and
[…]
then
$\left(\begin{array}[]{c}\hat{g}_{d}\left(p\right)\\ \hat{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)$ $=$ $\hat{Q}^{-1}\frac{1}{n}\sum_{i}\hat{w}_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\hat{\beta}_{d}\right)$
$=$ $\frac{1}{\left|\hat{Q}\right|}\frac{1}{n}\sum_{i}\hat{w}_{di}\left(p\right)\left(\begin{array}[]{c}\hat{B}_{2}-\left(\hat{P}_{i}-p\right)\hat{B}_{1}\\ -\hat{B}_{1}+\left(\hat{P}_{i}-p\right)\hat{B}_{0}\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\hat{\beta}_{d}\right),$
$\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)$ $=$ $\frac{1}{\left|Q\right|}\frac{1}{n}\sum_{i}w_{di}\left(p\right)\left(\begin{array}[]{c}B_{2}-\left(P_{i}-p\right)B_{1}\\ -B_{1}+\left(P_{i}-p\right)B_{0}\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\beta_{d}\right),$
where
[…]
For $r=0,1$, Further denote
$\hat{L}_{r}$ $=$ $\frac{1}{n}\sum_{i}\hat{w}_{di}\left(p\right)\left[\hat{B}_{r+1}-\left(\hat{P}_{i}-p\right)\hat{B}_{r}\right]\left(Y_{i}-X_{i}^{\prime}\hat{\beta}_{d}\right),$
$L_{r}$ $=$ $\frac{1}{n}\sum_{i}w_{di}\left(p\right)\left[B_{r+1}-\left(P_{i}-p\right)B_{r}\right]\left(Y_{i}-X_{i}^{\prime}\beta_{d}\right),$
then we have
[…]
It follows from Lemma E.2 that
$\frac{L_{1}}{\left|Q\right|}$ $=$ $g_{d}\left(p\right)+O_{p}\left(h_{3}^{2}+\frac{1}{\sqrt{nh_{3}}}\right)=g_{d}\left(p\right)+o_{p}\left(h_{3}\right),$
$\frac{L_{0}}{\left|Q\right|}$ $=$ $-g_{d}^{\left(1\right)}\left(p\right)+O_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)=-g_{d}^{\left(1\right)}\left(p\right)+o_{p}\left(1\right).$
Therefore,
$\hat{g}_{d}\left(p\right)-\tilde{g}_{d}\left(p\right)$ $=$ $\frac{\hat{L}_{1}}{\left|\hat{Q}\right|}-\frac{L_{1}}{\left|Q\right|}=\left[\left(\hat{L}_{1}-L_{1}\right)-\left(\left|\hat{Q}\right|-\left|Q\right|\right)\frac{L_{1}}{\left|Q\right|}\right]\frac{1}{\left|\hat{Q}\right|}$ (E.7)
$=$ $\frac{\left(\hat{L}_{1}-L_{1}\right)-\left(\left|\hat{Q}\right|-\left|Q\right|\right)g_{d}\left(p\right)+o_{p}\left(h_{3}\left(\left|\hat{Q}\right|-\left|Q\right|\right)\right)}{\left|\hat{Q}\right|}$,
$\hat{g}_{d}^{\left(1\right)}\left(p\right)-\tilde{g}_{d}^{\left(1\right)}\left(p\right)$ $=$ $-\frac{\left(\hat{L}_{0}-L_{0}\right)+\left(\left|\hat{Q}\right|-\left|Q\right|\right)g_{d}^{\left(1\right)}\left(p\right)+o_{p}\left(\left|\hat{Q}\right|-\left|Q\right|\right)}{\left|\hat{Q}\right|}$. (E.8)
Denote
$V_{di}\left(p\right)=Y_{i}-X_{i}^{\prime}\beta_{d}-g_{d}\left(p\right)=V_{di}+g_{d}\left(P_{i}\right)-g_{d}\left(p\right),$ (E.9)
which satisfies $E\left[\left.V_{di}\left(p\right)\right|X_{i},D_{i}=d\right]=g_{d}\left(P_{i}\right)-g_{d}\left(p\right)$ by definition, and denote
$\hat{C}_{r}$ $=$ $\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)^{r}V_{di}\left(p\right),$
$C_{r}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{r}V_{di}\left(p\right),$
$\hat{G}_{r}$ $=$ $\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)^{r}X_{i},$
for $r=0,1$. By several calculations,
$\left(\hat{L}_{1}-L_{1}\right)-\left(\left|\hat{Q}\right|-\left|Q\right|\right)g_{d}\left(p\right)$ $=$ $\left(\hat{C}_{0}\hat{B}_{2}-C_{0}B_{2}\right)-\left(\hat{C}_{1}\hat{B}_{1}-C_{1}B_{1}\right)$
$-\left(\hat{\beta}_{d}-\beta_{d}\right)^{\prime}\left(\hat{B}_{2}\hat{G}_{0}-\hat{B}_{1}\hat{G}_{1}\right),$
$\hat{L}_{0}-L_{0}$ $=$ $\left(\hat{C}_{0}\hat{B}_{1}-C_{0}B_{1}\right)-\left(\hat{C}_{1}\hat{B}_{0}-C_{1}B_{0}\right)$
$-\left(\hat{\beta}_{d}-\beta_{d}\right)^{\prime}\left(\hat{B}_{1}\hat{G}_{0}-\hat{B}_{0}\hat{G}_{1}\right),$
$\left|\hat{Q}\right|-\left|Q\right|$ $=$ $\left(\hat{B}_{0}\hat{B}_{2}-B_{0}B_{2}\right)-\left(\hat{B}_{1}^{2}-B_{1}^{2}\right).$
It follows from Lemmas E.3-E.7 that
$\hat{C}_{0}\hat{B}_{2}-C_{0}B_{2}$ $=$ $\left(\hat{C}_{0}-C_{0}\right)\left(\hat{B}_{2}-B_{2}\right)+C_{0}\left(\hat{B}_{2}-B_{2}\right)+\left(\hat{C}_{0}-C_{0}\right)B_{2}$
$=$ $O_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)+o_{p}\left(h_{3}\right)o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)$
$+\left(\hat{C}_{0}-C_{0}\right)\left[\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)h_{3}^{2}+o_{p}\left(h_{3}^{3}\right)\right]$
$=$ $\frac{\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(D_{i}-P_{i}\right)$
$+o_{p}\left(\sqrt{\frac{h_{3}^{3}}{n}}\right),$
$\hat{C}_{1}\hat{B}_{1}-C_{1}B_{1}$ $=$ $\left(\hat{C}_{1}-C_{1}\right)\left(\hat{B}_{1}-B_{1}\right)+\left(\hat{C}_{1}-C_{1}\right)B_{1}+C_{1}\left(\hat{B}_{1}-B_{1}\right)$
$=$ $O_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)+O_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)O_{p}\left(h_{3}^{2}\right)+O_{p}\left(h_{3}^{2}\right)o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)$
$=$ $o_{p}\left(\sqrt{\frac{h_{3}^{3}}{n}}\right),$
$\left(\hat{\beta}_{d}-\beta_{d}\right)^{\prime}\left(\hat{B}_{2}\hat{G}_{0}-\hat{B}_{1}\hat{G}_{1}\right)$ $\leq$ $\left\|\hat{\beta}_{d}-\beta_{d}\right\|\left[\begin{array}[]{c}\left(\left|\hat{B}_{2}-B_{2}\right|+\left|B_{2}\right|\right)\left\|\hat{G}_{0}\right\|\\ +\left(\left|\hat{B}_{1}-B_{1}\right|+\left|B_{1}\right|\right)\left\|\hat{G}_{1}\right\|\end{array}\right]$
$=$ $O_{p}\left(\frac{1}{\sqrt{n}}\right)\left[\begin{array}[]{c}\left(o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)+O_{p}\left(h_{3}^{2}\right)\right)O_{p}\left(1\right)\\ +\left(o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)+O_{p}\left(h_{3}^{2}\right)\right)O_{p}\left(h_{3}\right)\end{array}\right]$
$=$ $o_{p}\left(\sqrt{\frac{h_{3}^{3}}{n}}\right),$
$\hat{C}_{0}\hat{B}_{1}-C_{0}B_{1}$ $=$ $\left(\hat{C}_{0}-C_{0}\right)\left(\hat{B}_{1}-B_{1}\right)+C_{0}\left(\hat{B}_{1}-B_{1}\right)+\left(\hat{C}_{0}-C_{0}\right)B_{1}$
$=$ $O_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)+o_{p}\left(h_{3}\right)o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)+O_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)O_{p}\left(h_{3}^{2}\right)$
$=$ $o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
$\hat{C}_{1}\hat{B}_{0}-C_{1}B_{0}$ $=$ $\left(\hat{C}_{1}-C_{1}\right)\left(\hat{B}_{0}-B_{0}\right)+C_{1}\left(\hat{B}_{0}-B_{0}\right)+\left(\hat{C}_{1}-C_{1}\right)B_{0}$
$=$ $O_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)+O_{p}\left(h_{3}^{2}\right)o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)$
$+\left(\hat{C}_{1}-C_{1}\right)\left[\delta_{d}\left(p\right)f_{P}\left(p\right)+o_{p}\left(h_{3}\right)\right]$
$=$ $\frac{\delta_{d}\left(p\right)f_{P}\left(p\right)g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\left[k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)+k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\frac{P_{i}-p}{h_{3}}\right]$
$\cdot\frac{P_{i}-p}{h_{3}}\left(D_{i}-P_{i}\right)+o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
$\left(\hat{\beta}_{d}-\beta_{d}\right)^{\prime}\left(\hat{B}_{1}\hat{G}_{0}-\hat{B}_{0}\hat{G}_{1}\right)$ $\leq$ $\left\|\hat{\beta}_{d}-\beta_{d}\right\|\left[\begin{array}[]{c}\left(\left|\hat{B}_{1}-B_{1}\right|+\left|B_{1}\right|\right)\left\|\hat{G}_{0}\right\|\\ +\left(\left|\hat{B}_{0}-B_{0}\right|+\left|B_{0}\right|\right)\left\|\hat{G}_{1}\right\|\end{array}\right]$
$=$ $O_{p}\left(\frac{1}{\sqrt{n}}\right)\left[\begin{array}[]{c}\left(o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)+O_{p}\left(h_{3}^{2}\right)\right)O_{p}\left(1\right)\\ +\left(o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)+O_{p}\left(1\right)\right)O_{p}\left(h_{3}\right)\end{array}\right]$
$=$ $o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
$\hat{B}_{0}\hat{B}_{2}-B_{0}B_{2}$ $=$ $\left(\hat{B}_{0}-B_{0}\right)\left(\hat{B}_{2}-B_{2}\right)+B_{0}\left(\hat{B}_{2}-B_{2}\right)+\left(\hat{B}_{0}-B_{0}\right)B_{2}$
$=$ $o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)+O_{p}\left(1\right)o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)O_{p}\left(h_{3}^{2}\right)$
$=$ $o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
and
$\hat{B}_{1}^{2}-B_{1}^{2}$ $=$ $\left(\hat{B}_{1}-B_{1}\right)\left[\left(\hat{B}_{1}-B_{1}\right)+2B_{1}\right]$
$=$ $o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)\left[o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)+O_{p}\left(h_{3}^{2}\right)\right]$
$=$ $o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
Therefore,
$\left(\hat{L}_{1}-L_{1}\right)-\left(\left|\hat{Q}\right|-\left|Q\right|\right)g_{d}\left(p\right)$ $=$ $\frac{\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)$ (E.14)
$\cdot\left(P_{i}-p\right)\left(D_{i}-P_{i}\right)+o_{p}\left(\sqrt{\frac{h_{3}^{3}}{n}}\right),$
$\hat{L}_{0}-L_{0}$ $=$ $-\frac{\delta_{d}\left(p\right)f_{P}\left(p\right)g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\left[k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)+k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\frac{P_{i}-p}{h_{3}}\right]$ (E.15)
$\cdot\frac{P_{i}-p}{h_{3}}\left(D_{i}-P_{i}\right)+o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
and
$\left|\hat{Q}\right|-\left|Q\right|=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$ (E.16)
In addition, it follows from Lemma E.3 that
$\left|Q\right|=\kappa_{2}\delta_{d}^{2}\left(p\right)f_{P}^{2}\left(p\right)h_{3}^{2}+o_{p}\left(h_{3}^{2}\right).$
Hence,
$\frac{\left|\hat{Q}\right|}{\left|Q\right|}-1=\frac{\left|\hat{Q}\right|-\left|Q\right|}{\left|Q\right|}=\frac{o_{p}\left(1\left/\sqrt{nh_{3}^{3}}\right.\right)}{\kappa_{2}\delta_{d}^{2}\left(p\right)f_{P}^{2}\left(p\right)+o_{p}\left(1\right)}=o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)=o_{p}\left(1\right),$
and
$\frac{1}{\left|\hat{Q}\right|}=\frac{1}{\left|Q\right|}\left(1+o_{p}\left(1\right)\right)=\frac{1}{\kappa_{2}\delta_{d}^{2}\left(p\right)f_{P}^{2}\left(p\right)h_{3}^{2}}\left(1+o_{p}\left(1\right)\right).$ (E.17)
Substituting (E.14)-(E.17) into (E.7)-(E.8) yields that
$\hat{g}_{d}\left(p\right)-\tilde{g}_{d}\left(p\right)$ $=$ $\left(\frac{g_{d}^{\left(1\right)}\left(p\right)}{\delta_{d}\left(p\right)f_{P}\left(p\right)}\right)\frac{1}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\frac{P_{i}-p}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(D_{i}-P_{i}\right)$
$+o_{p}\left(\sqrt{\frac{1}{nh_{3}}}\right),$
$\hat{g}_{d}^{\left(1\right)}\left(p\right)-\tilde{g}_{d}^{\left(1\right)}\left(p\right)$ $=$ $\left(\frac{g_{d}^{\left(1\right)}\left(p\right)}{\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)}\right)\frac{1}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\left[\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\right$.
$\left.+\frac{P_{i}-p}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right]\frac{P_{i}-p}{h_{3}^{2}}\left(D_{i}-P_{i}\right)+o_{p}\left(\sqrt{\frac{1}{nh_{3}^{3}}}\right).$
Further combining with (E.1), we obtain
$J_{n}\left[\left(\begin{array}[]{c}\hat{g}_{d}\left(p\right)\\ \hat{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)\right]=\Gamma_{d}A_{dn}\left(p\right)+o_{p}\left(1\right),$ (E.18)
where
$A_{dn}\left(p\right)$ $=$ $J_{n}\frac{1}{n}\sum_{i=1}^{n}\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)\left[w_{di}\left(p\right)V_{di}+\left(\begin{array}[]{c}\Lambda_{di}\left(p\right)\\ \Lambda_{di}\left(p\right)+\Upsilon_{di}\left(p\right)\end{array}\right)\left(D_{i}-P_{i}\right)\right],$
$\Lambda_{di}\left(p\right)$ $=$ $g_{d}^{\left(1\right)}\left(p\right)\delta_{d}\left(P_{i}\right)\frac{P_{i}-p}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right),$
$\Upsilon_{di}\left(p\right)$ $=$ $g_{d}^{\left(1\right)}\left(p\right)\delta_{d}\left(P_{i}\right)\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right).$
Since
$E\left[\left.w_{di}\left(p\right)V_{di}\right|P_{i}\right]$ $=$ $\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\delta_{d}\left(P_{i}\right)E\left[\left.V_{di}\right|P_{i},D_{i}=d\right]=0,$
$E\left[\left.D_{i}-P_{i}\right|P_{i}\right]$ $=$ $0,$
$E\left[\left.w_{di}\left(p\right)V_{di}\left(D_{i}-P_{i}\right)\right|P_{i}\right]$ $=$ $\left(d-P_{i}\right)E\left[\left.w_{di}\left(p\right)V_{di}\right|P_{i}\right]=0,$
we have
$E\left[A_{dn}\left(p\right)\right]$ $=$ $0,$
$Var\left(A_{dn\left(1\right)}\left(p\right)\right)$ $=$ $h_{3}\left(E\left[w_{di}^{2}\left(p\right)V_{di}^{2}\right]+E\left[\Lambda_{di}^{2}\left(p\right)\left(D_{i}-P_{i}\right)^{2}\right]\right)$
$=$ $\kappa\delta_{d}\left(p\right)f_{P}\left(p\right)\sigma_{V_{d}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p\left(1-p\right)\delta_{d}^{2}\left(p\right)f_{P}\left(p\right)\left(g_{d}^{\left(1\right)}\left(p\right)\right)^{2}+o\left(1\right),$
$Var\left(A_{dn\left(2\right)}\left(p\right)\right)$ $=$ $h_{3}^{3}\left(E\left[\frac{\left(P_{i}-p\right)^{2}}{h_{3}^{4}}w_{di}^{2}\left(p\right)V_{di}^{2}\right]+E\left[\frac{\left(P_{i}-p\right)^{2}}{h_{3}^{4}}\left(\Lambda_{di}\left(p\right)+\Upsilon_{di}\left(p\right)\right)^{2}\left(D_{i}-P_{i}\right)^{2}\right]\right)$
$=$ $\kappa_{22}\delta_{d}\left(p\right)f_{P}\left(p\right)\sigma_{V_{d}}^{2}\left(p\right)+\left(\kappa_{24}^{\left(1\right)}-2\kappa_{22}\right)p\left(1-p\right)\delta_{d}^{2}\left(p\right)f_{P}\left(p\right)\left(g_{d}^{\left(1\right)}\left(p\right)\right)^{2}+o\left(1\right),$
$E\left[A_{dn\left(1\right)}\left(p\right)A_{dn\left(2\right)}\left(p\right)\right]$ $=$ $h_{3}^{2}\left(E\left[\frac{P_{i}-p}{h_{3}^{2}}w_{di}^{2}\left(p\right)V_{di}^{2}\right]+E\left[\frac{P_{i}-p}{h_{3}^{2}}\left(\Lambda_{di}^{2}\left(p\right)+\Lambda_{di}\left(p\right)\Upsilon_{di}\left(p\right)\right)\left(D_{i}-P_{i}\right)^{2}\right]\right)$
$=$ $O\left(h_{3}\right)=o\left(1\right),$
where
$\kappa_{22}^{\left(1\right)}$ $=$ $\int\left(k_{3}^{\left(1\right)}\left(u\right)\right)^{2}u^{2}du,$
$\kappa_{24}^{\left(1\right)}$ $=$ $\int\left(k_{3}^{\left(1\right)}\left(u\right)\right)^{2}u^{4}du.$
It then follows from Liapunov’s central limit theorem that
$A_{dn}\left(p\right)\rightarrow N\left(0,\text{diag}\left(\begin{array}[]{c}\kappa\delta_{d}\left(p\right)f_{P}\left(p\right)\sigma_{V_{d}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p\left(1-p\right)\delta_{d}^{2}\left(p\right)f_{P}\left(p\right)\left(g_{d}^{\left(1\right)}\left(p\right)\right)^{2}\\ \kappa_{22}\delta_{d}\left(p\right)f_{P}\left(p\right)\sigma_{V_{d}}^{2}\left(p\right)+\left(\kappa_{24}^{\left(1\right)}-2\kappa_{22}\right)p\left(1-p\right)\delta_{d}^{2}\left(p\right)f_{P}\left(p\right)\left(g_{d}^{\left(1\right)}\left(p\right)\right)^{2}\end{array}\right)\right),$
and thus from (E.18) that
$J_{n}\left[\left(\begin{array}[]{c}\hat{g}_{d}\left(p\right)\\ \hat{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)\right]\rightarrow N\left(0,\Sigma_{d}\left(p\right)\right),$
where
$\Sigma_{d}\left(p\right)=\left(\begin{array}[]{cc}\frac{\kappa\sigma_{V_{d}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p\left(1-p\right)\delta_{d}\left(p\right)\left(g_{d}^{\left(1\right)}\left(p\right)\right)^{2}}{\delta_{d}\left(p\right)f_{P}\left(p\right)}&0\\ 0&\frac{\kappa_{22}\sigma_{V_{d}}^{2}\left(p\right)+\left(\kappa_{24}^{\left(1\right)}-2\kappa_{22}\right)p\left(1-p\right)\delta_{d}\left(p\right)\left(g_{d}^{\left(1\right)}\left(p\right)\right)^{2}}{\kappa_{2}^{2}\delta_{d}\left(p\right)f_{P}\left(p\right)}\end{array}\right).$
Moreover, we can deduce the joint asymptotic normality of $\left(\hat{g}_{1}\left(p\right),\hat{g}_{1}^{\left(1\right)}\left(p\right)\right)$ and $\left(\hat{g}_{0}\left(p\right),\hat{g}_{0}^{\left(1\right)}\left(p\right)\right)$ from the asymptotically linear representation (E.18). Due to
$E\left[A_{1n\left(1\right)}\left(p\right)A_{0n\left(1\right)}\left(p\right)\right]$ $=$ $h_{3}E\left[\Lambda_{1i}\left(p\right)\Lambda_{0i}\left(p\right)\left(D_{i}-P_{i}\right)^{2}\right]=\kappa_{22}^{\left(1\right)}g_{1}^{\left(1\right)}\left(p\right)g_{0}^{\left(1\right)}\left(p\right)p^{2}\left(1-p\right)^{2}f_{P}\left(p\right)+o\left(1\right),$
$E\left[A_{1n\left(1\right)}\left(p\right)A_{0n\left(2\right)}\left(p\right)\right]$ $=$ $h_{3}^{2}E\left[\frac{P_{i}-p}{h_{3}^{2}}\Lambda_{1i}\left(p\right)\left(\Lambda_{0i}\left(p\right)+\Upsilon_{0i}\left(p\right)\right)\left(D_{i}-P_{i}\right)^{2}\right]=O\left(h_{3}\right)=o\left(1\right),$
$E\left[A_{1n\left(2\right)}\left(p\right)A_{0n\left(1\right)}\left(p\right)\right]$ $=$ $h_{3}^{2}E\left[\frac{P_{i}-p}{h_{3}^{2}}\left(\Lambda_{1i}\left(p\right)+\Upsilon_{1i}\left(p\right)\right)\Lambda_{0i}\left(p\right)\left(D_{i}-P_{i}\right)^{2}\right]=O\left(h_{3}\right)=o\left(1\right),$
$E\left[A_{1n\left(2\right)}\left(p\right)A_{0n\left(2\right)}\left(p\right)\right]$ $=$ $h_{3}^{3}E\left[\frac{\left(P_{i}-p\right)^{2}}{h_{3}^{4}}\left(\Lambda_{1i}\left(p\right)+\Upsilon_{1i}\left(p\right)\right)\left(\Lambda_{0i}\left(p\right)+\Upsilon_{0i}\left(p\right)\right)\left(D_{i}-P_{i}\right)^{2}\right]$
$=$ $\left(\kappa_{24}^{\left(1\right)}-2\kappa_{22}\right)g_{1}^{\left(1\right)}\left(p\right)g_{0}^{\left(1\right)}\left(p\right)p^{2}\left(1-p\right)^{2}f_{P}\left(p\right)+o\left(1\right),$
it follows from Liapunov’s central limit theorem that
$\left(\begin{array}[]{cc}J_{n}&0\\ 0&J_{n}\end{array}\right)\left[\left(\begin{array}[]{c}\hat{g}_{1}\left(p\right)\\ \hat{g}_{1}^{\left(1\right)}\left(p\right)\\ \hat{g}_{0}\left(p\right)\\ \hat{g}_{0}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{1}\left(p\right)\\ g_{1}^{\left(1\right)}\left(p\right)\\ g_{0}\left(p\right)\\ g_{0}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{1}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\\ \frac{\kappa_{2}}{2}g_{0}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)\right]\rightarrow N\left(0,\Sigma\left(p\right)\right),$ (E.20)
where
$\Sigma\left(p\right)=\left(\begin{array}[]{cc}\Sigma_{1}\left(p\right)&\Sigma_{10}\left(p\right)\\ \Sigma_{10}\left(p\right)&\Sigma_{0}\left(p\right)\end{array}\right),$
and
$\Sigma_{10}\left(p\right)=\left(\begin{array}[]{cc}\displaystyle\frac{\kappa_{22}^{\left(1\right)}p\left(1-p\right)g_{1}^{\left(1\right)}\left(p\right)g_{0}^{\left(1\right)}\left(p\right)}{f_{P}\left(p\right)}&0\\ 0&\displaystyle\frac{\left(\kappa_{24}^{\left(1\right)}-2\kappa_{22}\right)p\left(1-p\right)g_{1}^{\left(1\right)}\left(p\right)g_{0}^{\left(1\right)}\left(p\right)}{\kappa_{2}^{2}f_{P}\left(p\right)}\end{array}\right).$
Now the asymptotic normality of the estimators for MTE and relevant causal parameters can be readily deduced from (E.20). For the MTE estimator, we have
$\sqrt{nh_{3}^{3}}\left[\hat{\Delta}^{\text{MTE}}\left(x,v\right)-\Delta^{\text{MTE}}\left(x,v\right)\right]$
$=$ $x^{\prime}\left[\sqrt{nh_{3}^{3}}\left(\hat{\beta}_{1}-\beta_{1}\right)-\sqrt{nh_{3}^{3}}\left(\hat{\beta}_{0}-\beta_{0}\right)\right]$
$+\left[\sqrt{nh_{3}^{3}}\left(\hat{g}_{1}\left(v\right)-g_{1}\left(v\right)\right)-\sqrt{nh_{3}^{3}}\left(\hat{g}_{0}\left(v\right)-g_{0}\left(v\right)\right)\right]$
$+v\sqrt{nh_{3}^{3}}\left(\hat{g}_{1}^{\left(1\right)}\left(v\right)-g_{1}^{\left(1\right)}\left(v\right)\right)+\left(1-v\right)\sqrt{nh_{3}^{3}}\left(\hat{g}_{0}^{\left(1\right)}\left(v\right)-g_{0}^{\left(1\right)}\left(v\right)\right)$
$=$ $O_{p}\left(\sqrt{h_{3}^{3}}\right)+O\left(\sqrt{nh_{3}^{7}}\right)+O_{p}\left(h_{3}\right)$
$+v\sqrt{nh_{3}^{3}}\left(\hat{g}_{1}^{\left(1\right)}\left(v\right)-g_{1}^{\left(1\right)}\left(v\right)\right)+\left(1-v\right)\sqrt{nh_{3}^{3}}\left(\hat{g}_{0}^{\left(1\right)}\left(v\right)-g_{0}^{\left(1\right)}\left(v\right)\right)$
$\rightarrow$ $N\left(0,\sigma_{\text{MTE}}^{2}\left(v\right)\right),$
where
$\sigma_{\text{MTE}}^{2}\left(v\right)=\frac{\kappa_{22}\left[v\sigma_{V_{1}}^{2}\left(v\right)+\left(1-v\right)\sigma_{V_{0}}^{2}\left(v\right)\right]+\left(\kappa_{24}^{\left(1\right)}-2\kappa_{22}\right)v\left(1-v\right)\left[vg_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)g_{0}^{\left(1\right)}\left(v\right)\right]^{2}}{\kappa_{2}^{2}f_{P}\left(v\right)}.$ (E.21)
For the relevant causal parameter estimators, we have
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{ATE}}\left(x\right)-\Delta^{\text{ATE}}\left(x\right)-\frac{\kappa_{2}}{2}\left(g_{1}^{\left(2\right)}\left(1\right)-g_{0}^{\left(2\right)}\left(0\right)\right)h_{3}^{2}\right]$
$=$ $x^{\prime}\left[\sqrt{nh_{3}}\left(\hat{\beta}_{1}-\beta_{1}\right)-\sqrt{nh_{3}}\left(\hat{\beta}_{0}-\beta_{0}\right)\right]$
$+\sqrt{nh_{3}}\left[\hat{g}_{1}\left(1\right)-g_{1}\left(1\right)-\frac{\kappa_{2}}{2}g_{1}^{\left(2\right)}\left(1\right)h_{3}^{2}\right]-\sqrt{nh_{3}}\left[\hat{g}_{0}\left(0\right)-g_{0}\left(0\right)-\frac{\kappa_{2}}{2}g_{0}^{\left(2\right)}\left(0\right)h_{3}^{2}\right]$
$=$ $O_{p}\left(\sqrt{h_{3}}\right)+\sqrt{nh_{3}}\left[\hat{g}_{1}\left(1\right)-g_{1}\left(1\right)-\frac{\kappa_{2}}{2}g_{1}^{\left(2\right)}\left(1\right)h_{3}^{2}\right]-\sqrt{nh_{3}}\left[\hat{g}_{0}\left(0\right)-g_{0}\left(0\right)-\frac{\kappa_{2}}{2}g_{0}^{\left(2\right)}\left(0\right)h_{3}^{2}\right]$
$\rightarrow$ $N\left(0,\sigma_{\text{ATE}}^{2}\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{TT}}\left(x\right)-\Delta^{\text{TT}}\left(x\right)-\frac{\kappa_{2}}{2}\left(g_{1}^{\left(2\right)}\left(\pi\left(x\right)\right)+\frac{\left(1-\pi\left(x\right)\right)g_{0}^{\left(2\right)}\left(\pi\left(x\right)\right)-g_{0}^{\left(2\right)}\left(0\right)}{\pi\left(x\right)}\right)h_{3}^{2}\right]$
$=$ $x^{\prime}\left[\sqrt{nh_{3}}\left(\hat{\beta}_{1}-\beta_{1}\right)-\sqrt{nh_{3}}\left(\hat{\beta}_{0}-\beta_{0}\right)\right]+\sqrt{nh_{3}}\left[\hat{g}_{1}\left(\hat{\pi}\left(x\right)\right)-g_{1}\left(\pi\left(x\right)\right)\right]$
$+\sqrt{nh_{3}}\left[\frac{\left(1-\hat{\pi}\left(x\right)\right)\hat{g}_{0}\left(\hat{\pi}\left(x\right)\right)-\hat{g}_{0}\left(0\right)}{\hat{\pi}\left(x\right)}-\frac{\left(1-\pi\left(x\right)\right)g_{0}\left(\pi\left(x\right)\right)-g_{0}\left(0\right)}{\pi\left(x\right)}\right]$
$-\sqrt{nh_{3}}\frac{\kappa_{2}}{2}\left[g_{1}^{\left(2\right)}\left(\pi\left(x\right)\right)+\frac{1-\pi\left(x\right)}{\pi\left(x\right)}g_{0}^{\left(2\right)}\left(\pi\left(x\right)\right)-\frac{1}{\pi\left(x\right)}g_{0}^{\left(2\right)}\left(0\right)\right]h_{3}^{2}$
$=$ $O_{p}\left(\sqrt{h_{3}}\right)+O_{p}\left(\sqrt{nh_{3}}\eta_{1}\right)+\sqrt{nh_{3}}\left[\hat{g}_{1}\left(\pi\left(x\right)\right)-g_{1}\left(\pi\left(x\right)\right)-\frac{\kappa_{2}}{2}g_{1}^{\left(2\right)}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$
$+O_{p}\left(\sqrt{nh_{3}}\eta_{1}\right)+\frac{1-\pi\left(x\right)}{\pi\left(x\right)}\sqrt{nh_{3}}\left[\hat{g}_{0}\left(\pi\left(x\right)\right)-g_{0}\left(\pi\left(x\right)\right)-\frac{\kappa_{2}}{2}g_{0}^{\left(2\right)}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$
$+O_{p}\left(\sqrt{nh_{3}}\eta_{1}\right)-\frac{1}{\pi\left(x\right)}\sqrt{nh_{3}}\left[\hat{g}_{0}\left(0\right)-g_{0}\left(0\right)-\frac{\kappa_{2}}{2}g_{0}^{\left(2\right)}\left(0\right)h_{3}^{2}\right]$
$\rightarrow$ $N\left(0,\sigma_{\text{TT}}^{2}\left(\pi\left(x\right)\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{TUT}}\left(x\right)-\Delta^{\text{TUT}}\left(x\right)-\frac{\kappa_{2}}{2}\left(\frac{g_{1}^{\left(2\right)}\left(1\right)-\pi\left(x\right)g_{1}^{\left(2\right)}\left(\pi\left(x\right)\right)}{1-\pi\left(x\right)}-g_{0}^{\left(2\right)}\left(\pi\left(x\right)\right)\right)h_{3}^{2}\right]$
$=$ $-\sqrt{nh_{3}}\left[\hat{g}_{0}\left(\pi\left(x\right)\right)-g_{0}\left(\pi\left(x\right)\right)-\frac{\kappa_{2}}{2}g_{0}^{\left(2\right)}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$
$-\frac{\pi\left(x\right)}{1-\pi\left(x\right)}\sqrt{nh_{3}}\left[\hat{g}_{1}\left(\pi\left(x\right)\right)-g_{1}\left(\pi\left(x\right)\right)-\frac{\kappa_{2}}{2}g_{1}^{\left(2\right)}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$
$+\frac{1}{1-\pi\left(x\right)}\sqrt{nh_{3}}\left[\hat{g}_{1}\left(1\right)-g_{1}\left(1\right)-\frac{\kappa_{2}}{2}g_{1}^{\left(2\right)}\left(1\right)h_{3}^{2}\right]+o_{p}\left(1\right)$
$\rightarrow$ $N\left(0,\sigma_{\text{TUT}}^{2}\left(\pi\left(x\right)\right)\right),$
$\sqrt{nh_{3}}\left[\begin{array}[]{c}\hat{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)-\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)\\ -\frac{\kappa_{2}}{2}\left(\frac{v_{2}g_{1}^{\left(2\right)}\left(v_{2}\right)-v_{1}g_{1}^{\left(2\right)}\left(v_{1}\right)+\left(1-v_{2}\right)g_{0}^{\left(2\right)}\left(v_{2}\right)-\left(1-v_{1}\right)g_{0}^{\left(2\right)}\left(v_{1}\right)}{v_{2}-v_{1}}\right)h_{3}^{2}\end{array}\right]\rightarrow N\left(0,\sigma_{\text{LATE}}^{2}\left(v_{1},v_{2}\right)\right),$
where
$\sigma_{\text{ATE}}^{2}=\frac{\kappa\sigma_{V_{1}}^{2}\left(1\right)}{f_{P}\left(1\right)}+\frac{\kappa\sigma_{V_{0}}^{2}\left(0\right)}{f_{P}\left(0\right)},$ (E.22)
$\sigma_{\text{TT}}^{2}\left(p\right)$ $=$ $\frac{\kappa\sigma_{V_{1}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p^{2}\left(1-p\right)\left(g_{1}^{\left(1\right)}\left(p\right)\right)^{2}}{pf_{P}\left(p\right)}$ (E.23)
$+\left(\frac{1-p}{p}\right)^{2}\frac{\kappa\sigma_{V_{0}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p\left(1-p\right)^{2}\left(g_{0}^{\left(1\right)}\left(p\right)\right)^{2}}{\left(1-p\right)f_{P}\left(p\right)}$
$+\frac{2\kappa_{22}^{\left(1\right)}\left(1-p\right)^{2}g_{1}^{\left(1\right)}\left(p\right)g_{0}^{\left(1\right)}\left(p\right)}{f_{P}\left(p\right)}+\frac{\kappa\sigma_{V_{0}}^{2}\left(0\right)}{p^{2}f_{P}\left(0\right)},$
$\sigma_{\text{TUT}}^{2}\left(p\right)$ $=$ $\frac{\kappa\sigma_{V_{0}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p\left(1-p\right)^{2}\left(g_{0}^{\left(1\right)}\left(p\right)\right)^{2}}{\left(1-p\right)f_{P}\left(p\right)}$ (E.24)
$+\left(\frac{p}{1-p}\right)^{2}\frac{\kappa\sigma_{V_{1}}^{2}\left(p\right)+\kappa_{22}^{\left(1\right)}p^{2}\left(1-p\right)\left(g_{1}^{\left(1\right)}\left(p\right)\right)^{2}}{pf_{P}\left(p\right)}$
$+\frac{2\kappa_{22}^{\left(1\right)}p^{2}g_{1}^{\left(1\right)}\left(p\right)g_{0}^{\left(1\right)}\left(p\right)}{f_{P}\left(p\right)}+\frac{\kappa\sigma_{V_{1}}^{2}\left(1\right)}{\left(1-p\right)^{2}f_{P}\left(1\right)},$
$\sigma_{\text{LATE}}^{2}\left(v_{1},v_{2}\right)$ $=$ $\frac{\kappa}{\left(v_{2}-v_{1}\right)^{2}f_{P}\left(v_{1}\right)}\left[v_{1}\sigma_{V_{1}}^{2}\left(v_{1}\right)+\left(1-v_{1}\right)\sigma_{V_{0}}^{2}\left(v_{1}\right)\right]$ (E.25)
$+\frac{\kappa}{\left(v_{2}-v_{1}\right)^{2}f_{P}\left(v_{2}\right)}\left[v_{2}\sigma_{V_{1}}^{2}\left(v_{2}\right)+\left(1-v_{2}\right)\sigma_{V_{0}}^{2}\left(v_{2}\right)\right]$
$+\frac{\kappa_{22}^{\left(1\right)}v_{1}\left(1-v_{1}\right)}{\left(v_{2}-v_{1}\right)^{2}f_{P}\left(v_{1}\right)}\left[v_{1}g_{1}^{\left(1\right)}\left(v_{1}\right)+\left(1-v_{1}\right)g_{0}^{\left(1\right)}\left(v_{1}\right)\right]^{2}$
$+\frac{\kappa_{22}^{\left(1\right)}v_{2}\left(1-v_{2}\right)}{\left(v_{2}-v_{1}\right)^{2}f_{P}\left(v_{2}\right)}\left[v_{2}g_{1}^{\left(1\right)}\left(v_{2}\right)+\left(1-v_{2}\right)g_{0}^{\left(1\right)}\left(v_{2}\right)\right]^{2},$
which completes the proof.
E.3 Auxiliary lemmas
Lemma E.1.
Under Assumption E.1, we have
$\sup_{x}\left|\hat{\pi}\left(x\right)-\pi\left(x\right)\right|$ $=$ $O\left(\sum_{l=1}^{\dim\left(X^{C}\right)}h_{1l}^{s}+\sqrt{\frac{\ln n}{nh_{11}h_{12}\cdots h_{1,\dim\left(X^{C}\right)}}}\right)\text{ almost surely,}$
$\sup_{x}\left|\hat{f}_{X}\left(x\right)-f_{X}\left(x\right)\right|$ $=$ $O\left(\sum_{l=1}^{\dim\left(X^{C}\right)}h_{1l}^{s}+\sqrt{\frac{\ln n}{nh_{11}h_{12}\cdots h_{1,\dim\left(X^{C}\right)}}}\right)\text{ almost surely,}$
where $\hat{\pi}\left(x\right)$ is defined in (12) and $\hat{f}_{X}\left(x\right)$ is the normalized denominator of $\hat{\pi}\left(x\right)$, namely,
$\hat{f}_{X}\left(x\right)=\frac{1}{nh_{11}h_{12}\cdots h_{1,\dim\left(X^{C}\right)}}\sum_{i=1}^{n}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\frac{X_{il}^{C}-x_{l}^{C}}{h_{1l}}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}.$ (E.26)
Proof.
This proof follows straightforwardly from combining Subsection 1.11 and Theorems 1.4 and 2.6 of Li and Racine (2007), thus is omitted here. ∎
Lemma E.2.
Under Assumption E.2.(ii)-(iv), the infeasible local linear estimators $\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$, $d=0,1$, are asymptotically normally distributed for any interior point $p$ of the support of $P$:
$\left(\begin{array}[]{cc}\sqrt{nh_{3}}&0\\ 0&\sqrt{nh_{3}^{3}}\end{array}\right)\left[\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)\right]\rightarrow N\left(0,\Xi_{d}\left(p\right)\right),$
where
$\Xi_{d}\left(p\right)=\left(\begin{array}[]{cc}\left.\kappa\sigma_{V_{d}}^{2}\left(p\right)\right/\left[\delta_{d}\left(p\right)f_{P}\left(p\right)\right]&0\\ 0&\left.\kappa_{22}\sigma_{V_{d}}^{2}\left(p\right)\right/\left[\kappa_{2}^{2}\delta_{d}\left(p\right)f_{P}\left(p\right)\right]\end{array}\right),$
and
$\kappa$ $=$ $\int k_{3}^{2}\left(u\right)du,$
$\kappa_{2}$ $=$ $\int k_{3}\left(u\right)u^{2}du,$
$\kappa_{22}$ $=$ $\int k_{3}^{2}\left(u\right)u^{2}du.$
Proof.
Denote
$Q_{d}=\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)^{\prime}$
as the normalized denominator of $\left(\tilde{g}_{d}\left(p\right),\tilde{g}_{d}^{\left(1\right)}\left(p\right)\right)$. By $Y-X^{\prime}\beta_{d}=g_{d}\left(P\right)+V_{d}$, we have
$\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)$ $=$ $Q_{d}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)\left[g_{d}\left(P_{i}\right)+V_{di}\right]$
$=$ $\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)+Q_{d}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)\left[\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}+R_{di}+V_{di}\right],$
where the second equality follows from a Taylor expansion of $g_{d}\left(P_{i}\right)$ around $P_{i}=p$, and $R_{di}=\frac{1}{6}g_{d}^{\left(3\right)}\left(P_{i}^{\ast}\right)\left(P_{i}-p\right)^{3}$ with $P_{i}^{\ast}$ an intermediate value. Note that the three times differentiability of $g_{d}\left(\cdot\right)$ required in the expansion is stronger than Assumption E.2.(ii) and is imposed only for notational simplicity. Actually, the twice continuous differentiability of $g_{d}\left(\cdot\right)$ imposed by Assumption E.2.(ii) suffices for the same conclusion to hold.
By Lemma E.3,
$Q_{d}=\left(\begin{array}[]{cc}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right),&\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)\\ \frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right),&\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{2}\end{array}\right)=\left(\begin{array}[]{cc}O_{p}\left(1\right)&O_{p}\left(h_{3}^{2}\right)\\ O_{p}\left(h_{3}^{2}\right)&O_{p}\left(h_{3}^{2}\right)\end{array}\right),$
which is an asymptotically singular matrix. Hence, we further normalize $Q_{d}$ by a matrix,
$G=\left(\begin{array}[]{cc}1,&0\\ 0,&h_{3}^{-2}\end{array}\right),$
such that
$GQ_{d}$ $=$ $\left(\begin{array}[]{cc}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right),&\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)\\ \frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right),&\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{2}\end{array}\right)$
$\overset{p}{\longrightarrow}$ $T_{d}\equiv\left(\begin{array}[]{cc}\delta_{d}\left(p\right)f_{P}\left(p\right),&0\\ \kappa_{2}\left[\delta_{d}^{\left(1\right)}\left(p\right)f_{P}\left(p\right)+\delta_{d}\left(p\right)f_{P}^{\left(1\right)}\left(p\right)\right],&\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)\end{array}\right),$
Since $T_{d}$ is invertible, we have
$\left(GQ_{d}\right)^{-1}\overset{p}{\longrightarrow}T_{d}^{-1}=\left(\begin{array}[]{cc}1,&0\\ -\left[\left.\delta_{d}^{\left(1\right)}\left(p\right)\right/\delta_{d}\left(p\right)+\left.f_{P}^{\left(1\right)}\left(p\right)\right/f_{P}\left(p\right)\right],&1\left/\kappa_{2}\right.\end{array}\right)\frac{1}{\delta_{d}\left(p\right)f_{P}\left(p\right)}.$
Therefore,
$J_{n}\left[\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)\right]$
$=$ $J_{n}\left(GQ_{d}\right)^{-1}G\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ P_{i}-p\end{array}\right)\left[\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}+R_{di}+V_{di}\right]$
$=$ $J_{n}T_{d}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)\left[\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}+R_{di}+V_{di}\right]+o_{p}\left(1\right),$
where
$J_{n}=\left(\begin{array}[]{cc}\sqrt{nh_{3}}&0\\ 0&\sqrt{nh_{3}^{3}}\end{array}\right).$
Note that for
[…]
we have
$J_{n}T_{d}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)\left[\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}+R_{di}+V_{di}\right]$
$=$ $J_{n}\Gamma_{d}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)\left[\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}+R_{di}+V_{di}\right]+o_{p}\left(1\right),$
because
$\sqrt{nh_{3}^{3}}T_{d\left(2,1\right)}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}$ $=$ $O\left(\sqrt{nh_{3}^{7}}+h_{3}^{3}\right)=o\left(1\right),$
$\sqrt{nh_{3}^{3}}T_{d\left(2,1\right)}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)R_{di}$ $=$ $O\left(\sqrt{nh_{3}^{9}}+h_{3}^{4}\right)=o\left(1\right),$
$\sqrt{nh_{3}^{3}}T_{d\left(2,1\right)}^{-1}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)V_{di}$ $=$ $O\left(h_{3}\right)=o\left(1\right),$
where
$T_{d\left(2,1\right)}^{-1}=-\frac{\left.\delta_{d}^{\left(1\right)}\left(p\right)\right/\delta_{d}\left(p\right)+\left.f_{P}^{\left(1\right)}\left(p\right)\right/f_{P}\left(p\right)}{\delta_{d}\left(p\right)f_{P}\left(p\right)}.$
Therefore,
$J_{n}\left[\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)\right]$ (E.40)
$=$ $J_{n}\Gamma_{d}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)\left[\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2}+R_{di}+V_{di}\right]+o_{p}\left(1\right)$ (E.43)
$=$ $\Gamma_{d}\left[A_{1d}\left(p\right)+A_{2d}\left(p\right)+A_{3d}\left(p\right)\right]+o_{p}\left(1\right),$ (E.44)
where
$A_{1d}\left(p\right)$ $=$ $\left(\begin{array}[]{c}A_{1d\left(1\right)}\left(p\right)\\ A_{1d\left(2\right)}\left(p\right)\end{array}\right)\equiv J_{n}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)\frac{1}{2}g_{d}^{\left(2\right)}\left(p\right)\left(P_{i}-p\right)^{2},$
$A_{2d}\left(p\right)$ $=$ $\left(\begin{array}[]{c}A_{2d\left(1\right)}\left(p\right)\\ A_{2d\left(2\right)}\left(p\right)\end{array}\right)\equiv J_{n}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)R_{di}$
$=$ $J_{n}\frac{1}{6n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)g_{d}^{\left(3\right)}\left(P_{i}^{\ast}\right)\left(P_{i}-p\right)^{3},$
$A_{3d}\left(p\right)$ $=$ $\left(\begin{array}[]{c}A_{3d\left(1\right)}\left(p\right)\\ A_{3d\left(2\right)}\left(p\right)\end{array}\right)\equiv J_{n}\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\begin{array}[]{c}1\\ \frac{P_{i}-p}{h_{3}^{2}}\end{array}\right)V_{di}.$
We first consider $A_{1d}\left(p\right)$. Since
$E\left[A_{1d\left(1\right)}\left(p\right)\right]$ $=$ $\frac{\sqrt{nh_{3}}}{2}g_{d}^{\left(2\right)}\left(p\right)E\left[w_{di}\left(p\right)\left(P_{i}-p\right)^{2}\right]=\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)\sqrt{nh_{3}^{5}}+o\left(1\right),$
$E\left[A_{1d\left(2\right)}\left(p\right)\right]$ $=$ $\frac{\sqrt{nh_{3}^{3}}}{2}g_{d}^{\left(2\right)}\left(p\right)E\left[w_{di}\left(p\right)\frac{P_{i}-p}{h_{3}^{2}}\left(P_{i}-p\right)^{2}\right]=O\left(\sqrt{nh_{3}^{7}}\right)=o\left(1\right),$
and
$Var\left(A_{1d\left(1\right)}\left(p\right)\right)$ $\leq$ $\frac{h_{3}}{4}\left[g_{d}^{\left(2\right)}\left(p\right)\right]^{2}E\left[w_{di}^{2}\left(p\right)\left(P_{i}-p\right)^{4}\right]=O\left(h_{3}^{2}\right)=o\left(1\right),$
$Var\left(A_{1d\left(2\right)}\left(p\right)\right)$ $\leq$ $\frac{h_{3}^{3}}{4}\left[g_{d}^{\left(2\right)}\left(p\right)\right]^{2}E\left[w_{di}^{2}\left(p\right)\frac{\left(P_{i}-p\right)^{6}}{h_{3}^{4}}\right]=O\left(h_{3}^{4}\right)=o\left(1\right),$
it follows from Markov’s inequality that
$A_{1d\left(1\right)}\left(p\right)$ $=$ $\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)\sqrt{nh_{3}^{5}}+o_{p}\left(1\right),$
$A_{1d\left(2\right)}\left(p\right)$ $=$ $o_{p}\left(1\right),$
namely,
$A_{1d}\left(p\right)=J_{n}\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)+o_{p}\left(1\right).$ (E.49)
Similarly, we can show that
$A_{2d}\left(p\right)=o_{p}\left(1\right).$ (E.50)
We next consider $A_{3d}\left(p\right)$. By the definition of $V_{di}$, we have $E\left[A_{3d}\left(p\right)\right]=0$. For the variance-covariance term, we have
$Var\left(A_{3d\left(1\right)}\left(p\right)\right)=h_{3}E\left[w_{di}^{2}\left(p\right)V_{di}^{2}\right]=\kappa\sigma_{V_{d}}^{2}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)+O\left(h_{3}\right),$
$Var\left(A_{3d\left(2\right)}\left(p\right)\right)=h_{3}^{3}E\left[w_{di}^{2}\left(p\right)\frac{\left(P_{i}-p\right)^{2}}{h_{3}^{4}}V_{di}^{2}\right]=\kappa_{22}\sigma_{V_{d}}^{2}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)+O\left(h_{3}\right),$
$Cov\left(A_{3d\left(1\right)}\left(p\right),A_{3d\left(2\right)}\left(p\right)\right)=h_{3}^{2}E\left[w_{di}^{2}\left(p\right)\frac{P_{i}-p}{h_{3}^{2}}V_{di}^{2}\right]=O\left(h_{3}\right).$
It then follows from Liapunov’s central limit theorem that
$A_{3d}\left(p\right)\rightarrow N\left(0,\Omega_{d}\right),$ (E.51)
where
$\Omega_{d}=\left(\begin{array}[]{cc}\kappa\sigma_{V_{d}}^{2}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)&0\\ 0&\kappa_{22}\sigma_{V_{d}}^{2}\left(p\right)\delta_{d}\left(p\right)f_{P}\left(p\right)\end{array}\right).$
Substituting (E.49) and (E.50) into (E.44) yields that
$J_{n}\left[\left(\begin{array}[]{c}\tilde{g}_{d}\left(p\right)\\ \tilde{g}_{d}^{\left(1\right)}\left(p\right)\end{array}\right)-\left(\begin{array}[]{c}g_{d}\left(p\right)\\ g_{d}^{\left(1\right)}\left(p\right)\end{array}\right)\right]=J_{n}\left(\begin{array}[]{c}\frac{\kappa_{2}}{2}g_{d}^{\left(2\right)}\left(p\right)h_{3}^{2}\\ 0\end{array}\right)+\Gamma_{d}A_{3d}\left(p\right)+o_{p}\left(1\right),$ (E.52)
and the conclusion follows from (E.51). ∎
Lemma E.3.
Under Assumption E.2, we have
$\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)$ $=$ $\delta_{d}\left(p\right)f_{P}\left(p\right)+o_{p}\left(h_{3}\right),$
$\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)$ $=$ $\kappa_{2}\left[\delta_{d}^{\left(1\right)}\left(p\right)f_{P}\left(p\right)+\delta_{d}\left(p\right)f_{P}^{\left(1\right)}\left(p\right)\right]+O_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right),$
$\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{2}$ $=$ $\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)+o_{p}\left(h_{3}\right).$
Proof.
Since the data are i.i.d., we have
$E\left[\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\right]$ $=$ $E\left[1\left\{D_{i}=d\right\}\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\right]$
$=$ $E\left[\delta_{d}\left(P_{i}\right)\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\right]$
$=$ $\int\delta_{d}\left(v\right)\frac{1}{h_{3}}k_{3}\left(\frac{v-p}{h_{3}}\right)f_{P}\left(v\right)dv$
$=$ $\int\delta_{d}\left(p+uh_{3}\right)k_{3}\left(u\right)f_{P}\left(p+uh_{3}\right)du$
$=$ $\delta_{d}\left(p\right)f_{P}\left(p\right)+O\left(h_{3}^{2}\right),$
where the second equality follows from the law of iterated expectation. On the other hand,
$Var\left(\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\right)$ $=$ $\frac{1}{nh_{3}^{2}}Var\left(1\left\{D_{i}=d\right\}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\right)$
$\leq$ $\frac{1}{nh_{3}^{2}}E\left[1\left\{D_{i}=d\right\}k_{3}^{2}\left(\frac{P_{i}-p}{h_{3}}\right)\right]$
$=$ $O\left(\frac{1}{nh_{3}}\right).$
Therefore,
$\left(E\left[\left(\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)-\delta_{d}\left(p\right)f_{P}\left(p\right)\right)^{2}\right]\right)^{1/2}=O\left(h_{3}^{2}+\frac{1}{\sqrt{nh_{3}}}\right)=o\left(h_{3}\right),$
where the last equality is by Assumption E.2.(iv). The first assertion of the lemma then follows from Markov’s inequality.
For the second assertion of the lemma, we have
$E\left[\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)\right]$ $=$ $E\left[\delta_{d}\left(P_{i}\right)\frac{1}{h_{3}^{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\right]$
$=$ $\int\delta_{d}\left(v\right)\frac{1}{h_{3}^{3}}k_{3}\left(\frac{v-p}{h_{3}}\right)\left(v-p\right)f_{P}\left(v\right)dv$
$=$ $\frac{1}{h_{3}}\int\delta_{d}\left(p+uh_{3}\right)f_{P}\left(p+uh_{3}\right)k_{3}\left(u\right)udu$
$=$ $\kappa_{2}\left[\delta_{d}^{\left(1\right)}\left(p\right)f_{P}\left(p\right)+\delta_{d}\left(p\right)f_{P}^{\left(1\right)}\left(p\right)\right]+O\left(h_{3}^{2}\right),$
and
$Var\left(\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)\right)$ $\leq$ $\frac{1}{nh_{3}^{6}}E\left[1\left\{D_{i}=d\right\}k_{3}^{2}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)^{2}\right]$
$=$ $O\left(\frac{1}{nh_{3}^{3}}\right).$
Therefore,
$\left(E\left[\left(\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)-\kappa_{2}\left[\delta_{d}^{\left(1\right)}\left(p\right)f_{P}\left(p\right)+\delta_{d}\left(p\right)f_{P}^{\left(1\right)}\left(p\right)\right]\right)^{2}\right]\right)^{1/2}$ $=$ $O\left(h_{3}^{2}+\frac{1}{\sqrt{nh_{3}^{3}}}\right)$
$=$ $O\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right),$
and the second assertion of the lemma follows. Analogously, the third assertion of the lemma can be readily shown. ∎
Lemma E.4.
Under Assumption E.2, we have
$\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)V_{di}\left(p\right)$ $=$ $o_{p}\left(h_{3}\right),$
$\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)V_{di}\left(p\right)$ $=$ $\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)g_{d}^{\left(1\right)}\left(p\right)+O_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right),$
where $V_{di}\left(p\right)$ is defined in (E.9).
Proof.
Note that $E\left[\left.V_{di}\left(p\right)\right|P_{i},D_{i}=d\right]=g_{d}\left(P_{i}\right)-g_{d}\left(p\right)$ because $E\left[\left.V_{di}\right|P_{i},D_{i}=d\right]=0$ by definition. Hence,
$E\left[\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)V_{di}\left(p\right)\right]$ $=$ $E\left[\delta_{d}\left(P_{i}\right)\left(g_{d}\left(P_{i}\right)-g_{d}\left(p\right)\right)\frac{1}{h_{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\right]$
$=$ $O\left(h_{3}^{2}\right),$
$E\left[\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)V_{di}\left(p\right)\right]$ $=$ $E\left[\delta_{d}\left(P_{i}\right)\left(g_{d}\left(P_{i}\right)-g_{d}\left(p\right)\right)\frac{1}{h_{3}^{3}}k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\right]$
$=$ $\kappa_{2}\delta_{d}\left(p\right)f_{P}\left(p\right)g_{d}^{\left(1\right)}\left(p\right)+O\left(h_{3}^{2}\right).$
For the variances, the same arguments as in Lemma E.3 show that
$Var\left(\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)V_{di}\left(p\right)\right)$ $=$ $O\left(\frac{1}{nh_{3}}\right),$
$Var\left(\frac{1}{nh_{3}^{2}}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)V_{di}\left(p\right)\right)$ $=$ $O\left(\frac{1}{nh_{3}^{3}}\right).$
The conclusion then follows from Markov’s inequality. ∎
Lemma E.5.
Under Assumptions E.1, E.2, and E.3, we have
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)$ $=$ $o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right),$
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)$ $=$ $o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)^{2}-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{2}$ $=$ $o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
Proof.
By a Taylor expansion,
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)$
$+\frac{1}{2n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{2}$
$+\frac{1}{6n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{3},$
where $\hat{P}_{i}^{\ast}$ is an intermediate value between $\hat{P}_{i}$ and $P_{i}$. For any i.i.d. nonnegative random variables $\left\{Z_{ni}:1\leq i\leq n\right\}$ with $0<EZ_{ni}<\infty$, we have $\left.\left(1/n\right)\sum_{i=1}^{n}Z_{ni}\right/EZ_{ni}=O_{p}\left(1\right)$ by Markov’s inequality. Thus, it follows from Lemma E.1 and Assumption E.3.(i)-(ii) that
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)\right|$ $\leq$ $\sup_{i}\left|\hat{P}_{i}-P_{i}\right|\frac{1}{n}\sum_{i=1}^{n}\frac{1}{h_{3}^{2}}\left|k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|$
$=$ $O_{p}\left(\eta_{1}\right)O_{p}\left(\frac{1}{h_{3}^{2}}E\left|k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|\right)$
$=$ $O_{p}\left(\frac{\eta_{1}}{h_{3}}\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right),$
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{2}\right|$ $\leq$ $O_{p}\left(\eta_{1}^{2}\right)O_{p}\left(\frac{1}{h_{3}^{3}}E\left|k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|\right)$
$=$ $O_{p}\left(\frac{\eta_{1}^{2}}{h_{3}^{2}}\right)=o_{p}\left(\frac{1}{nh_{3}^{3}}\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right),$
and
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{3}\right|\leq O_{p}\left(\frac{\eta_{1}^{3}}{h_{3}^{4}}\right)=o_{p}\left(\frac{1}{h_{3}^{4}}\sqrt{\frac{h_{3}^{7}}{n}}\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right).$
Therefore,
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right).$
For the second claim of the lemma, we decompose the left-hand side into three terms:
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)$ (E.53)
$=$ $\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)+\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)$
$+\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right).$
For the first term, we have
$\left|\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)\right|\leq O_{p}\left(\eta_{1}\right)O_{p}\left(E\left|w_{di}\left(p\right)\right|\right)=O_{p}\left(\eta_{1}\right)O_{p}\left(1\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$ (E.54)
For the second term, a Taylor expansion yields that
$\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)$
$+\frac{1}{2n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}$
$+\frac{1}{6n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{3}.$
It follows that
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)\right|$
$\leq$ $\sup_{i}\left|\hat{P}_{i}-P_{i}\right|\frac{1}{n}\sum_{i=1}^{n}\frac{1}{h_{3}^{2}}\left|k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|\left|P_{i}-p\right|$
$=$ $O_{p}\left(\eta_{1}\right)O_{p}\left(\frac{1}{h_{3}}E\left[\left|k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|\left|\frac{P_{i}-p}{h_{3}}\right|\right]\right)$
$=$ $O_{p}\left(\eta_{1}\right)O_{p}\left(1\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}\right|$
$\leq$ $O_{p}\left(\eta_{1}^{2}\right)O_{p}\left(\frac{1}{h_{3}^{2}}E\left|k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|\left|\frac{P_{i}-p}{h_{3}}\right|\right)$
$=$ $O_{p}\left(\frac{\eta_{1}^{2}}{h_{3}}\right)=o_{p}\left(\frac{1}{nh_{3}^{2}}\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$
and
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{3}\right|\leq O_{p}\left(\frac{\eta_{1}^{3}}{h_{3}^{4}}\right)=o_{p}\left(\frac{1}{h_{3}^{4}}\sqrt{\frac{h_{3}^{7}}{n}}\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$
Therefore,
$\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$ (E.55)
For the third term of (E.53), we have
$\left|\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)\right|\leq\sup_{i}\left|\hat{P}_{i}-P_{i}\right|\frac{1}{n}\sum_{i=1}^{n}\left|\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right|.$
As for the first claim of the lemma, we can show that
$\frac{1}{n}\sum_{i=1}^{n}\left|\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right|=o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)=o_{p}\left(1\right).$ (E.56)
It then follows that
$\left|\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)\right|\leq O_{p}\left(\eta_{1}\right)o_{p}\left(1\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$ (E.57)
Substituting (E.54), (E.55), and (E.57) into (E.53), we obtain
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$
For the third claim of the lemma, we decompose the left-hand side into five terms:
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)^{2}-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)^{2}$ (E.58)
$=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)^{2}+\frac{2}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)\left(P_{i}-p\right)$
$+\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}+\frac{2}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)\left(P_{i}-p\right)$
$+\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)^{2}.$
For the first term, it follows from (E.56) that
$\left|\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)^{2}\right|$ $\leq$ $\sup_{i}\left|\hat{P}_{i}-P_{i}\right|^{2}\frac{1}{n}\sum_{i=1}^{n}\left|\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right|$ (E.59)
$=$ $O_{p}\left(\eta_{1}^{2}\right)o_{p}\left(1\right)=o_{p}\left(\frac{1}{nh_{3}}\right)$
$=$ $o_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\sqrt{\frac{h_{3}}{n}}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
For the second term, a Taylor expansion yields that
$\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)\left(P_{i}-p\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}$
$+\frac{1}{2n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{3}.$
Analogously, it follows that
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}\right|$ $\leq$ $O_{p}\left(\eta_{1}^{2}\right)=o_{p}\left(\frac{1}{nh_{3}}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
$\left|\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{3}\right|$ $\leq$ $O_{p}\left(\frac{\eta_{1}^{3}}{h_{3}^{3}}\right)=o_{p}\left(\frac{\sqrt{h_{3}^{7}}}{h_{3}^{3}\sqrt{n}}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
Therefore,
$\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)\left(P_{i}-p\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$ (E.60)
For the third and fourth terms of (E.58), we have
$\left|\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}\right|$ $\leq$ $O_{p}\left(\eta_{1}^{2}\right)O_{p}\left(E\left|w_{di}\left(p\right)\right|\right)$ (E.61)
$=$ $O_{p}\left(\eta_{1}^{2}\right)O_{p}\left(1\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
$\left|\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)\left(P_{i}-p\right)\right|$ $\leq$ $O_{p}\left(\eta_{1}\right)O_{p}\left(E\left|w_{di}\left(p\right)\left(P_{i}-p\right)\right|\right)$ (E.62)
$=$ $O_{p}\left(\eta_{1}h_{3}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
For the last term of (E.58), we can show as well by a Taylor expansion that
$\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)^{2}=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$ (E.63)
The conclusion follows from combining (E.58)-(E.63). ∎
Lemma E.6.
Under Assumptions E.1, E.2, E.3, and E.4, we have
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)V_{di}\left(p\right)-\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)V_{di}\left(p\right)$
$=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\frac{1}{h_{3}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\frac{P_{i}-p}{h_{3}}\left(D_{i}-P_{i}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)$
$=$ $O_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$
and
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)V_{di}\left(p\right)-\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)V_{di}\left(p\right)$
$=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\left[k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)+k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\frac{P_{i}-p}{h_{3}}\right]\frac{P_{i}-p}{h_{3}}\left(D_{i}-P_{i}\right)+o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right)$
$=$ $O_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
where $V_{di}\left(p\right)$ is defined in (E.9).
Proof.
By a Taylor expansion,
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)V_{di}\left(p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)V_{di}\left(p\right)$ (E.64)
$=$ $\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)V_{di}\left(p\right)$
$+\frac{1}{2n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{2}V_{di}\left(p\right)$
$+\frac{1}{6n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{3}V_{di}\left(p\right)$
$\triangleq$ $\left(\text{I}\right)+\left(\text{II}\right)+\left(\text{III}\right),$
where $\hat{P}_{i}^{\ast}$ is an intermediate value between $\hat{P}_{i}$ and $P_{i}$. Note that $\left(\text{I}\right)$ and $\left(\text{II}\right)$ can be represented as U-statistics of order two and three, respectively, plus higher order terms. Specifically, by denoting $W_{i}=\left(Y_{i},D_{i},X_{i}\right)$ as the $i$-th observation, and
$K_{1h}\left(X_{j},X_{i}\right)$ $=$ $\prod_{l=1}^{\dim\left(X^{C}\right)}\frac{1}{\left|h_{1}\right|}k_{1}\left(\frac{X_{jl}^{C}-X_{il}^{C}}{h_{1l}}\right)1\left\{X_{j}^{D}=X_{i}^{D}\right\},$
$m_{1}\left(W_{i},W_{j}\right)$ $=$ $1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right)\left(D_{j}-P_{i}\right)K_{1h}\left(X_{j},X_{i}\right),$
$m_{2}\left(W_{i},W_{j},W_{l}\right)$ $=$ $1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right)\left(D_{j}-P_{i}\right)K_{1h}\left(X_{j},X_{i}\right)\left(D_{l}-P_{i}\right)K_{1h}\left(X_{l},X_{i}\right),$
$M_{1}$ $=$ $\frac{1}{n\left(n-1\right)}\sum_{i}\sum_{j\neq i}\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)},$
$M_{2}$ $=$ $\frac{1}{n\left(n-1\right)\left(n-2\right)}\sum_{i}\sum_{j\neq i}\sum_{l\neq j,i}\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)},$
we have
$\left(\text{I}\right)$ $=$ $\frac{1}{n\left(n-1\right)}\sum_{i}\sum_{j\neq i}\frac{m_{1}\left(W_{i},W_{j}\right)}{\hat{f}_{X}\left(X_{i}\right)}=M_{1}\cdot\left(1+O_{p}\left(\eta_{1}\right)\right),$ (E.65)
$\left(\text{II}\right)$ $=$ $\frac{1}{n\left(n-1\right)^{2}}\sum_{i}\sum_{j\neq i}\sum_{l\neq i}\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{\hat{f}_{X}^{2}\left(X_{i}\right)}=M_{2}\cdot\left(1+O_{p}\left(\eta_{1}\right)\right),$ (E.66)
where $\hat{f}_{X}\left(x\right)$ is defined in (E.26) and the last equalities follow from Lemma E.1.
To analyze the U-statistics $M_{1}$ and $M_{2}$, we need to calculate the conditional expectations of their kernels:
$E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{i}\right]$ $=$ $O\left(tr\left(h_{1}^{s}\right)\right)\cdot 1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right),$
$E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{j}\right]$ $=$ $\delta_{d}\left(P_{j}\right)\frac{g_{d}\left(P_{j}\right)-g_{d}\left(p\right)}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)\left[1+O\left(tr\left(h_{1}^{s}\right)\right)\right],$
$E\left[\left.\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)}\right|W_{i}\right]$ $=$ $O\left(tr\left(h_{1}^{2s}\right)\right)\cdot 1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right),$
$E\left[\left.\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)}\right|W_{j}\right]$ $=$ $O\left(tr\left(h_{1}^{s}\right)\right)\cdot\delta_{d}\left(P_{j}\right)\frac{g_{d}\left(P_{j}\right)-g_{d}\left(p\right)}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)$
$\cdot\left[1+O\left(tr\left(h_{1}^{s}\right)\right)\right],$
$E\left[\left.\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)}\right|W_{l}\right]$ $=$ $O\left(tr\left(h_{1}^{s}\right)\right)\cdot\delta_{d}\left(P_{l}\right)\frac{g_{d}\left(P_{l}\right)-g_{d}\left(p\right)}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{l}-p}{h_{3}}\right)\left(D_{l}-P_{l}\right)$
$\cdot\left[1+O\left(tr\left(h_{1}^{s}\right)\right)\right],$
where we use $E\left[\left.V_{di}\left(p\right)\right|X_{i},D_{i}=d\right]=g_{d}\left(P_{i}\right)-g_{d}\left(p\right)$, and hence,
$E\left[M_{1}\right]$ $=$ $E\left[E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{j}\right]\right]=0,$
$E\left[M_{2}\right]$ $=$ $E\left[E\left[\left.\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)}\right|W_{j}\right]\right]=0.$
Since by Assumptions E.1.(ii) and E.3.(iii)-(iv),
$E\left[\frac{m_{1}^{2}\left(W_{i},W_{j}\right)}{f_{X}^{2}\left(X_{i}\right)}\right]$ $=$ $O\left(\frac{1}{h_{3}^{3}\left|h_{1}\right|}\right)=o\left(n\right),$
$E\left[\frac{m_{2}^{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{4}\left(X_{i}\right)}\right]$ $=$ $O\left(\frac{1}{h_{3}^{5}\left|h_{1}\right|^{2}}\right)=o\left(n^{3/2}\right),$
it follows from the projection method for U-statistics (e.g., Powell et al., 1989, Lemma 3.1) that
$M_{1}=\frac{1}{n}\sum_{i=1}^{n}E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{i}\right]+\frac{1}{n}\sum_{j=1}^{n}E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{j}\right]+o_{p}\left(\frac{1}{\sqrt{n}}\right),$
and from the variances of U-statistics (e.g., Shao, 2003, Corollary 3.2) that
$Var\left(M_{2}\right)$ $\leq$ $\frac{3}{n}\left[Var\left(E\left[\left.\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)}\right|W_{i}\right]\right)+2Var\left(E\left[\left.\frac{m_{2}\left(W_{i},W_{j},W_{l}\right)}{f_{X}^{2}\left(X_{i}\right)}\right|W_{j}\right]\right)\right]+O\left(\frac{1}{n^{2}}\right)$
$=$ $O\left(\frac{tr\left(h_{1}^{4s}\right)}{nh_{3}^{5}}+\frac{tr\left(h_{1}^{2s}\right)}{nh_{3}}+\frac{1}{n^{2}}\right)=O\left(\frac{1}{n^{3}h_{3}^{7}}+\frac{1}{n^{2}h_{3}^{2}}+\frac{1}{n^{2}}\right)=O\left(\frac{1}{n^{3}h_{3}^{7}}+\frac{1}{n^{2}h_{3}^{2}}\right),$
[…]
$\frac{1}{n}\sum_{i=1}^{n}E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{i}\right]=O\left(tr\left(h_{1}^{s}\right)\right)\cdot\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right),$
and
$E\left[\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right)\right]$ $=$ $E\left[\delta_{d}\left(P_{i}\right)\frac{g_{d}\left(P_{i}\right)-g_{d}\left(p\right)}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right]$
$=$ $O\left(1\right),$
$Var\left(\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)V_{di}\left(p\right)\right)$ $\leq$ $\frac{1}{n}E\left[\frac{1}{h_{3}^{4}}\left|k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\right|^{2}V_{di}^{2}\left(p\right)\right]$
$=$ $O\left(\frac{1}{nh_{3}^{3}}\right)=o\left(1\right),$
it follows from Markov’s inequality that
$\frac{1}{n}\sum_{i=1}^{n}E\left[\left.\frac{m_{1}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{i}\right]=O\left(tr\left(h_{1}^{s}\right)\right)\cdot O_{p}\left(1\right)=O_{p}\left(tr\left(h_{1}^{s}\right)\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$
Hence,
$M_{1}$ $=$ $\frac{1}{n}\sum_{j=1}^{n}\delta_{d}\left(P_{j}\right)\frac{g_{d}\left(P_{j}\right)-g_{d}\left(p\right)}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)$ (E.67)
$=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{j=1}^{n}\delta_{d}\left(P_{j}\right)\frac{P_{j}-p}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$
On the other hand,
$E\left[M_{2}^{2}\right]=Var\left(M_{2}\right)=O\left(\frac{1}{n^{3}h_{3}^{7}}+\frac{1}{n^{2}h_{3}^{2}}\right)=o\left(\frac{1}{nh_{3}}\right),$
and thus by Markov’s inequality,
$M_{2}=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$ (E.68)
Substituting (E.67) and (E.68) into (E.65) and (E.66), respectively, we obtain
$\left(\text{I}\right)+\left(\text{II}\right)=\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\frac{1}{h_{3}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\frac{P_{i}-p}{h_{3}}\left(D_{i}-P_{i}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$ (E.69)
For the term $\left(\text{III}\right)$, it follows from Lemma E.1 that
$\left(\text{III}\right)\leq\frac{1}{6n}\sum_{i=1}^{n}\frac{1}{h_{3}^{4}}\left|k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\right|\left|\hat{P}_{i}-P_{i}\right|^{3}=O_{p}\left(\frac{\eta_{1}^{3}}{h_{3}^{4}}\right)=o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$ (E.70)
where the last equality follows from Assumption E.3.(ii). The first claim of the lemma follows from substituting (E.69) and (E.70) into (E.64).
For the second claim, we decompose the left-hand side into three terms:
$\frac{1}{n}\sum_{i=1}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)V_{di}\left(p\right)-\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(P_{i}-p\right)V_{di}\left(p\right)$ (E.71)
$=$ $\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)V_{di}\left(p\right)+\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)V_{di}\left(p\right)$
$+\frac{1}{n}\sum_{i=1}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)V_{di}\left(p\right).$
Denote
$m_{3}\left(W_{i},W_{j}\right)$ $=$ $\frac{1}{h_{3}}w_{di}\left(p\right)V_{di}\left(p\right)\left(D_{j}-P_{i}\right)K_{1h}\left(X_{j},X_{i}\right),$
$M_{3}$ $=$ $\frac{1}{n\left(n-1\right)}\sum_{i}\sum_{j\neq i}\frac{m_{3}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)},$
then it follows from Lemma E.1 that
$\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)V_{di}\left(p\right)=\frac{h_{3}}{n\left(n-1\right)}\sum_{i}\sum_{j\neq i}\frac{m_{3}\left(W_{i},W_{j}\right)}{\hat{f}_{X}\left(X_{i}\right)}=h_{3}M_{3}\cdot\left(1+O_{p}\left(\eta_{1}\right)\right).$
Since
$E\left[\frac{m_{3}^{2}\left(W_{i},W_{j}\right)}{f_{X}^{2}\left(X_{i}\right)}\right]$ $=$ $O\left(\frac{1}{h_{3}^{3}\left|h_{1}\right|}\right)=o\left(n\right),$
$E\left[\left.\frac{m_{3}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{i}\right]$ $=$ $O\left(tr\left(h_{1}^{s}\right)\right)\cdot\frac{1}{h_{3}}w_{di}\left(p\right)V_{di}\left(p\right),$
$E\left[\left.\frac{m_{3}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{j}\right]$ $=$ $\delta_{d}\left(P_{j}\right)\left[g_{d}\left(P_{j}\right)-g_{d}\left(p\right)\right]\frac{1}{h_{3}^{2}}k_{3}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)\left[1+O\left(tr\left(h_{1}^{s}\right)\right)\right],$
$E\left[\frac{m_{3}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right]$ $=$ $0,$
it follows from the projection of U-statistics that
$M_{3}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}E\left[\left.\frac{m_{3}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{i}\right]+\frac{1}{n}\sum_{j=1}^{n}E\left[\left.\frac{m_{3}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)}\right|W_{j}\right]+o_{p}\left(\frac{1}{\sqrt{n}}\right)$
$=$ $\frac{1}{n}\sum_{j=1}^{n}\delta_{d}\left(P_{j}\right)\frac{g_{d}\left(P_{j}\right)-g_{d}\left(p\right)}{h_{3}^{2}}k_{3}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)$
$=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{j=1}^{n}\delta_{d}\left(P_{j}\right)\frac{P_{j}-p}{h_{3}^{2}}k_{3}\left(\frac{P_{j}-p}{h_{3}}\right)\left(D_{j}-P_{j}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$
Therefore,
$\frac{1}{n}\sum_{i=1}^{n}w_{di}\left(p\right)\left(\hat{P}_{i}-P_{i}\right)V_{di}\left(p\right)$ $=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)k_{3}\left(\frac{P_{i}-p}{h_{3}}\right)\frac{P_{i}-p}{h_{3}}\left(D_{i}-P_{i}\right)$ (E.72)
$+o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
For the second term in the decomposition (E.53), a Taylor expansion gives
$\frac{1}{n}\sum_{i}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)V_{di}\left(p\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}V_{di}\left(p\right)\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)$
$+\frac{1}{2n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}V_{di}\left(p\right)\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{2}$
$+\frac{1}{6n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}V_{di}\left(p\right)\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(\hat{P}_{i}-P_{i}\right)^{3}$
$\triangleq$ $\left(\text{IV}\right)+\left(\text{V}\right)+\left(\text{VI}\right).$
Denote
$m_{4}\left(W_{i},W_{j}\right)$ $=$ $1\left\{D_{i}=d\right\}V_{di}\left(p\right)\frac{1}{h_{3}^{3}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(D_{j}-P_{i}\right)K_{1h}\left(X_{j},X_{i}\right),$
$M_{4}$ $=$ $\frac{1}{n\left(n-1\right)}\sum_{i}\sum_{j\neq i}\frac{m_{4}\left(W_{i},W_{j}\right)}{f_{X}\left(X_{i}\right)},$
then we have
$\left(\text{IV}\right)=\frac{h_{3}}{n\left(n-1\right)}\sum_{i}\sum_{j\neq i}\frac{m_{4}\left(W_{i},W_{j}\right)}{\hat{f}_{X}\left(X_{i}\right)}=h_{3}M_{4}\cdot\left(1+O_{p}\left(\eta_{1}\right)\right).$
Analogously, it follows from the projection of U-statistics that
$M_{4}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\frac{g_{d}\left(P_{i}\right)-g_{d}\left(p\right)}{h_{3}^{3}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\left(D_{i}-P_{i}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right)$
$=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)\frac{1}{h_{3}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\frac{P_{i}-p}{h_{3}}\right)^{2}\left(D_{i}-P_{i}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right),$
and hence
$\left(\text{IV}\right)=\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\frac{P_{i}-p}{h_{3}}\right)^{2}\left(D_{i}-P_{i}\right)+o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
Moreover, it follows from Lemma E.1 that
$\left|\left(\text{V}\right)\right|$ $\leq$ $O_{p}\left(\eta_{1}^{2}\right)\cdot O_{p}\left(E\left|V_{di}\left(p\right)\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(P_{i}-p\right)\right|\right)=O_{p}\left(\eta_{1}^{2}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right),$
$\left|\left(\text{VI}\right)\right|$ $\leq$ $O_{p}\left(\eta_{1}^{3}\right)\cdot O_{p}\left(E\left|V_{di}\left(p\right)\frac{1}{h_{3}^{4}}k_{3}^{\left(3\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(P_{i}-p\right)\right|\right)=O_{p}\left(\frac{\eta_{1}^{3}}{h_{3}^{3}}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
Therefore, we have
$\frac{1}{n}\sum_{i}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(P_{i}-p\right)V_{di}\left(p\right)$ (E.73)
$=$ $\frac{g_{d}^{\left(1\right)}\left(p\right)}{n}\sum_{i=1}^{n}\delta_{d}\left(P_{i}\right)k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\frac{P_{i}-p}{h_{3}}\right)^{2}\left(D_{i}-P_{i}\right)+o_{p}\left(\frac{1}{\sqrt{nh_{3}}}\right).$
For the third term in the decomposition (E.53), a Taylor expansion gives
$\frac{1}{n}\sum_{i}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left(\hat{P}_{i}-P_{i}\right)V_{di}\left(p\right)$ (E.74)
$=$ $\frac{1}{n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}V_{di}\left(p\right)\frac{1}{h_{3}^{2}}k_{3}^{\left(1\right)}\left(\frac{P_{i}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{2}$
$+\frac{1}{2n}\sum_{i=1}^{n}1\left\{D_{i}=d\right\}V_{di}\left(p\right)\frac{1}{h_{3}^{3}}k_{3}^{\left(2\right)}\left(\frac{\hat{P}_{i}^{\ast}-p}{h_{3}}\right)\left(\hat{P}_{i}-P_{i}\right)^{3}$
$=$ $O_{p}\left(\eta_{1}^{2}\right)+O_{p}\left(\frac{\eta_{1}^{3}}{h_{3}^{3}}\right)=o_{p}\left(\sqrt{\frac{h_{3}}{n}}\right).$
The second claim of the lemma then follows from inserting (E.72), (E.73), and (E.74) into the decomposition (E.71). ∎
Lemma E.7.
Under Assumptions E.1, E.2, E.3, and E.4, we have
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)X_{i}$ $=$ $O_{p}\left(1\right),$
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)X_{i}$ $=$ $O_{p}\left(h_{3}\right).$
Proof.
Note that
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)X_{i}=\frac{1}{n}\sum_{i}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]X_{i}+\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)X_{i}.$ (E.75)
Similar to the first claim of Lemma E.5, we can show that
$\frac{1}{n}\sum_{i}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]X_{i}=O_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)=o_{p}\left(1\right).$ (E.76)
For the second term, we have
$E\left[\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)X_{i}\right]=O\left(1\right)$
and
$Var\left(\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)X_{i}\right)=O\left(\frac{1}{nh_{3}}\right)=o\left(1\right),$
hence,
$\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)X_{i}=O_{p}\left(1\right).$ (E.77)
The first claim of the lemma follows from substituting (E.76) and (E.77) into (E.75).
For the second claim, it follows from the compactness of the support of $k_{3}\left(\cdot\right)$ imposed in Assumption E.2.(iii) that
$\left\|\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)X_{i}\right\|\leq\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left|\hat{P}_{i}-p\right|\left\|X_{i}\right\|=O\left(h_{3}\right)\cdot\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left\|X_{i}\right\|.$
As above, we can show that
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left\|X_{i}\right\|$ $=$ $\frac{1}{n}\sum_{i}^{n}\left[\hat{w}_{di}\left(p\right)-w_{di}\left(p\right)\right]\left\|X_{i}\right\|+\frac{1}{n}\sum_{i}^{n}w_{di}\left(p\right)\left\|X_{i}\right\|$
$=$ $O_{p}\left(\frac{1}{\sqrt{nh_{3}^{3}}}\right)+O_{p}\left(1\right)=O_{p}\left(1\right).$
Therefore,
$\frac{1}{n}\sum_{i}^{n}\hat{w}_{di}\left(p\right)\left(\hat{P}_{i}-p\right)X_{i}=O_{p}\left(h_{3}\right).$
Appendix F Implementation guidance
Table F.1 summarizes our estimation procedure and preliminary tests for the assumptions. In general, we recommend that the data cell we choose for testing should have at least 50 observations. For example, there are 40 among all 60 data cells in Table F.2 satisfying this criterion. If we examine the supports of the 40 data cells one by one using the verification method as in our application, we will find that Assumption S holds for all of them. Moreover, we plot the conditional propensity score functions for all the 40 data cells in Figure F.1, which verifies that Assumption NL1 (nonmonotonicity) holds for all of them as well. In other words, we can verify the validity of Assumptions S and NL1 in our application regardless of which data cell we choose. This shows that Assumptions S and NL1 are quite weak and hardly limit the applicability of our identification strategy in practice. On the other hand, if the number of discrete covariates is not small and/or all data cells have less than 50 observations, we should consider to use a semiparametric or parametric approach that imposes a linear index structure on the treatment function. In this case, we should remember to include nonlinear terms of continuous covariates in the linear index to ensure the identification.
Appendix G Simulation
We examine the finite sample property of the MTE estimation based on our non-IV identification strategy by a Monte Carlo simulation. The treatment variable data are generated according to the structural treatment equation (1):
$D_{i}=1\left\{\mu\left(X_{i}\right)\geq U_{i}\right\},\text{ }i=1,2,\cdots,n,$
where the function $\mu\left(\cdot\right)$ and the distribution of error $U_{i}$ will vary across different designs. The vector $X_{i}$ consists of one continuous covariate and five discrete covariates, namely, $X_{i}=\left(X_{i}^{C},X_{1i}^{D},X_{2i}^{D},X_{3i}^{D},X_{4i}^{D},X_{5i}^{D}\right)^{\prime}$, where $X_{i}^{C}$ follows a standard normal distribution and $X_{ji}^{D}$ ($j=1,\cdots,5$) are independent $\left\{0,1\right\}$-valued binary random variables with identical distribution that $\Pr\left(X_{ji}^{D}=1\right)=0.5$. The potential and observed outcomes are generated by
$Y_{1i}$ $=$ $\beta_{1}^{C}X_{i}^{C}+\sum_{j=1}^{5}\left(X_{ji}^{D}-0.5\right)+U_{1i},$
$Y_{0i}$ $=$ $\beta_{0}^{C}X_{i}^{C}+\sum_{j=1}^{5}\frac{1}{6-j}\left(X_{ji}^{D}-0.5\right)+U_{0i},$
$Y_{i}$ $=$ $D_{i}Y_{1i}+\left(1-D_{i}\right)Y_{0i},\text{ }i=1,2,\cdots,n,$
where $U_{1i}$ is generated by $U_{1i}=-0.75U_{i}+\sqrt{1-0.75^{2}}\varrho_{i}$ with a standard normal random variable $\varrho_{i}$ that is independent of $U_{i}$ and $X_{i}$, and $U_{0i}$ follows an independent standard normal distribution as well. The sample size $n$ is set 4000.
We consider 16 different designs. In the benchmark design (Design 1), the structural treatment error $U_{i}$ follows a standard normal distribution that is independent of $X_{i}$, and the treatment function $\mu\left(\cdot\right)$ is quadratic in the continuous covairate and linear in the discrete covariates. Specifically, we set
$\mu\left(X_{i}\right)=X_{i}^{C}+\frac{1}{2}\left[\left(X_{i}^{C}\right)^{2}-1\right]+\sum_{j=1}^{5}\frac{1}{j}\left(X_{ji}^{D}-0.5\right).$ (G.1)
Note that $E\left[\mu\left(X_{i}\right)\right]=0$ and the probability of treatment is about $0.5$. In addition, we impose the exclusion restriction $\beta_{1}^{C}=\beta_{0}^{C}=0$ in Design 1. Other designs include cases of no exclusion restriction, linear or cubic treatment function $\mu\left(\cdot\right)$, nonnormal treatment error $U_{i}$, nonzero correlation between $U_{i}$ and $X_{i}^{C}$ as in Example 3, heteroskedastic $U_{i}$ or $U_{di}$, nonnormal forms of MTE, and nonlinear outcome equations. Table G.1 collects detailed descriptions of the designs. For each design, the simulation replicates 1000 times.
In every replication $b=1,\cdots,1000$, we compute different estimators for the MTE evaluated at $x=\bar{X}$, the sample mean of $X_{i}$, and $v=0.01,0.02,\cdots,0.99$. The first estimator is fully parametric, assuming jointly normally distributed errors and quadratic $\mu\left(\cdot\right)$: $\mu\left(X_{i}\right)=\gamma_{0}+\gamma_{1}X_{i}^{C}+\gamma_{2}\left(X_{i}^{C}\right)^{2}+\sum_{j=1}^{5}\gamma_{2+j}X_{ji}^{D}$. The normal assumption leads to the widely-used Heckman’s two-step estimation. Nevertheless, the parametric estimator has a high risk of misspecification. More flexible estimators comprise of a nonparametric first step (12) and a parametric second step. In the first step, we use the second-order ($s=2$) Gaussian kernel as $k_{1}\left(\cdot\right)$ because higher-order kernels are likely to result in unstable estimates in practice (Pan et al., 2022), and use the rule-of-thumb bandwidth as given in Table F.1. In the second step, we consider three parametric specifications on the unobservable heterogeneity of MTE, namely, a normal specification that $E\left[\left.U_{di}\right|V_{i}=v\right]=\rho_{d0}+\rho_{dV}\Phi^{-1}\left(v\right)$, a linear specification that $E\left[\left.U_{di}\right|V_{i}=v\right]=\theta_{d0}+\theta_{d1}v$, and a quadratic specification that $E\left[\left.U_{di}\right|V_{i}=v\right]=\theta_{d0}+\theta_{d1}v+\theta_{d2}v^{2}$, where $V_{i}=F_{\left.U\right|X}\left(\left.U_{i}\right|X_{i}\right)$ is the reduced-form treatment error. The most flexible estimator under Assumptions L and NL is the semiparametric one that comprises of a nonparametric first step and a partially linear second step, leading to the kernel-weighted pairwise difference estimator (13)-(14). We use the Gaussian kernel and rule-of-thumb bandwidths as well. Theorem 2 has established the asymptotic normality of this semiparametric MTE estimator. Recall that all the above estimators impose no exclusion restriction and base the identification on different functional forms between the treatment and outcome equations. As a comparison, we also consider the conventional IV-based MTE estimator under the semiparametric specification, whose implementation differs from the semiparametric non-IV MTE estimator only in that it assumes the exclusion restriction $\beta_{1}^{C}=\beta_{0}^{C}=0$. For each MTE estimator, we calculate the average over replications, that is, $\left.\sum_{b=1}^{1000}\hat{\Delta}_{b}^{\text{MTE}}\left(\bar{X},v\right)\right/1000$. And we depict the averaged MTE curves, along with the true MTE curve, for each design in Figures G.1-G.4.
In the benchmark design, the parametric and normal MTE estimators are correctly specified. Therefore, we can see from the upper-left panel of Figure G.1 that the parametric MTE performs perfectly and the normal MTE comes in second, and both of them are superior to the semiparametric one which exploits the least information of the model. Since the exclusion restriction holds in this design, the IV-based MTE is also consistent and performs well. However, by comparing the two semiparametric estimators, we find little improvement of the IV-based MTE over the non-IV MTE, implying that the exclusion restriction provides not much extra identifying power when the identification based on functional forms holds. The misspecified linear and quadratic MTE estimators perform satisfactorily as well, mainly because the true MTE curve is close to linear. When the exclusion restriction is violated as in Design 2, the upper-right panel of Figure G.1 shows that the IV-based MTE estimator deteriorates drastically due to misspecification. In Design 3, Assumption NL is violated and the identification based on functional forms fails to hold. The bottom-left panel of Figure G.1 confirms that in this case, the normal, linear, quadratic, and semiparametric non-IV MTE estimators are underidentified, while the IV-based MTE provides much improvement over its non-IV counterpart. Somewhat surprisingly, however, the parametric MTE still performs perfectly, suggesting a potential success of the classical parametric identification based on functional forms, even when the semiparametric identification fails. The bottom-right panel of Figure G.1 shows that estimation based on functional forms is more robust than that based on IV to the violation of identifying conditions.
Besides the curves, we also summarize the simulation results of the MTE estimators by the bias $\left.\sum_{b=1}^{1000}\left[\hat{\Delta}_{b}^{\text{MTE}}\left(\bar{X},v\right)-\Delta^{\text{MTE}}\left(\bar{X},v\right)\right]\right/1000$, the root mean squared error (RMSE) $\sqrt{\left.\sum_{b=1}^{1000}\left[\hat{\Delta}_{b}^{\text{MTE}}\left(\bar{X},v\right)-\Delta^{\text{MTE}}\left(\bar{X},v\right)\right]^{2}\right/1000}$, and the coverage probability of 95% confidence interval $\left.\sum_{b=1}^{1000}1\left\{\left|\hat{\Delta}_{b}^{\text{MTE}}\left(\bar{X},v\right)-\Delta^{\text{MTE}}\left(\bar{X},v\right)\right|\leq 1.96\cdot SE\left(\hat{\Delta}^{\text{MTE}}\left(\bar{X},v\right)\right)\right\}\right/1000$, where $SE\left(\hat{\Delta}^{\text{MTE}}\left(\bar{X},v\right)\right)$ is approximated by the simulated standard deviation of the estimator. Table G.3 reports the summary statistics of $\hat{\Delta}_{b}^{\text{MTE}}\left(\bar{X},0.5\right)$ for Designs 1-4, along with those of the ATE estimator and of $\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ that is the estimated partial effect of the continuous covariate on treatment effects. In Design 1, the coverage probabilities of all the estimators are close to the true size 95%, partly demonstrating the validity of a simple $t$-test when the identification is attained. In Design 2, the IV-based estimator is heavily biased in both estimation and inference, while in Design 3, the non-IV estimators are biased except the fully parametric one, which is consistent with the finding of Figure G.1. The summary statistics in Design 4 show that estimation based on functional forms is much more robust than that based on IV in terms of inference as well.
In the remaining designs, we relax the exclusion restriction and no longer report the results for the IV-based estimator. Design 5 considers a case that allows for nonzero correlation between the continuous covariate and structural treatment error but Assumption CMI holds, as in Example 3. The upper-left panel of Figure G.2 shows that the parametric MTE curve is still almost completely overlapped with the true MTE curve, while the other four MTE curves are more biased than in the uncorrelated case (Designs 1-2). Although biased, the semiparametric MTE estimator is fairly robust in terms of inference as shown by Table G.4. Design 6 is inspired by Example 6, where nonlinearity comes from the multiplicative heteroscedasticity in the treatment error. In this design, the quadratic treatment functional form of the parametric MTE is wrongly specified. Although the misspecification does not lead to large bias in estimation, it does lead to apparent invalidation of inference revealed by the severely biased coverage probability of the parametric estimator. In comparison, the normal estimator that allows for a nonparametric treatment equation performs satisfactorily in terms of the coverage probability. Designs 7 and 8 specify normal polynomial functional forms on the unobservable heterogeneity of MTE, under which the considered parametric second steps are all misspecified. Therefore, the semiparametric estimator that allows for nonparametric functional form of the unobservable heterogeneity of MTE has the best overall performance in these two designs, as illustrated by the lower panels of Figure G.2 and the right half part of Table G.4.
Figures G.3 and G.4 depict the MTE curves in additional designs. Designs 9 and 10 consider nonnormally distributed treatment error $U_{i}$, under which the parametric MTE estimator is misspecified and expected to be inconsistent. However, the upper-left panel of Figure G.3 shows that the parametric MTE still works in Design 9. This is mainly because the MTE estimator depends only on the propensity score (namely, the conditional expectation of treatment) but not on regression coefficients of the treatment equation. The estimation of the propensity score is actually a prediction problem. Some forms of misspecification may have a less serious impact on the prediction than on the regression. In Design 9, the $t\left(3\right)$ distribution of the treatment error is symmetric about zero and resembles the normal distribution, thus exerting only a small influence on the parametric MTE. In comparison, the $t\left(2\right)$ distributed treatment error in Design 10 is more fat-tailed and therefore leads to more biased parametric MTE. The nonnormally distributed treatment error also results in nonnormal unobservable heterogeneity of MTE. As in Designs 7 and 8, the semiparametric MTE performs the best overall in such case. Design 11 considers a cubic treament function $\mu\left(\cdot\right)$, under which the parametric MTE is notably biased due to severe misspecification. The last panel of Figure G.3 for Design 12 shows that the heteroskedasticity in outcome errors makes little difference to the performance of the MTE estimators by comparing with the benchmark design.
Figure G.4 investigates the impact of nonlinearity in the outcome equation, and the results are also as expected. That is, if we correctly include the nonlinear terms during implementation, the MTE estimator still performs well as shown by the dotted lines that stand for the correctly specified semiparametric estimator. However, if we omit the nonlinear terms in the outcome equation, the MTE estimator may be biased, and the magnitude of bias depends on the specific forms of nonlinearity. When the nonlinear term is an interaction of two discrete covariates as in Design 13, the omission of it has quite limited influence on the MTE estimators. Nevertheless, the omitted variable problem will become obvious if the interaction term pertains to the continuous covariate as in Design 14. The lower panels of Figure G.4 further show that omitting a quadratic term or an exponential term in the outcome equation will induce more substantial bias in the estimation of MTE.
In summary, the fully parametric estimator and the estimators that specify nonparametric first step and parametric second step are the first choice when we are confident that the specification is correct or at least nearly correct. Otherwise, the semiparametric estimator is perferred because its performance is the most robust across different designs.
Table F.1: An overview of the testing and estimation procedures
Step 0: Test Assumptions S and NL
Step 0.1: List the data cells for all possible values of discrete covariates and the corresponding conditional supports of the propensity score, as in Table F.2.
Step 0.2: Find a data cell that satisfies Assumption S and has at least 50 observations.
Note: A method of quick validation of Assumption S is to check whether the present data cell has full support or whether its support is overlapped with all other data cells’ supports.
Step 0.3: Test Assumption NL for the data cell selected by Step 0.2.
If there is only one continuous covariate, then plot the propensity score function for the selected data cell as in Figure 1, and visually inspect whether the function is nonmonotonic.
If there are at least two continuous covariates, then test the nonlinearity of the propensity score function by methods from the literature on single-index models (e.g., Fan and Li, 1996; Stute and Zhu, 2005; Xia, 2009; Maistre and Patilea, 2019).
Step 0.4: If Assumption NL holds, go to Step 1. Otherwise, return to Step 0.2-0.3 until finding one data cell for which Assumptions S and NL both hold.
Step 1: Estimate the treatment equation $D=1\left\{\mu\left(X\right)\geq U\right\}=1\left\{\pi\left(X\right)\geq V\right\}$
Case 1.1: Nonparametric (kernel) estimation. Suppose that both the functional form of $\mu\left(\cdot\right)$ and the distribution of the structural treatment error $U$ are nonparametrically specified. Then, estimate the propensity score $\pi\left(X\right)$ by the kernel estimator (12).
Note: We recommend to use the Gaussian kernel (i.e., standard normal density) as $k_{1}\left(\cdot\right)$ and the rule-of-thumb bandwidth $h_{1}=sd\left(X^{C}\right)\cdot n^{-1\left/\left(4+\dim\left(X^{C}\right)\right)\right.}$, where $sd\left(X^{C}\right)$ is the sample standard deviation of the continuous covariate.
Case 1.2: Semiparametric (single-index) estimation. Suppose that $\mu\left(\cdot\right)$ is parametrically specified as a linear index and the distribution of $U$ is nonparametrically specified. Then, estimate $\pi\left(X\right)$ by methods from the literature on single-index models (e.g., Powell et al., 1989; Ichimura, 1993; Klein and Spady, 1993; Lewbel, 2000).
Case 1.3: Parametric (Probit/Logit) estimation. Suppose that $\mu\left(\cdot\right)$ is parametrically specified and $U$ is assumed to follow a normal/logistic distribution. Then, perform a Probit/Logit regression on the treatment equation and predict $\pi\left(X\right)$.
Step 2: Estimate the outcome equation (8) and MTE (9)
Case 2.1: Semiparametric (partially linear) estimation. Suppose that the functional form of $g_{d}\left(\cdot\right)$ is nonparametrically specified. Then, estimate $\beta_{d}$ by the kernel-weighted pairwise difference estimator (13) and $g_{d}\left(\cdot\right)$ by the local linear estimator.
Note: We recommend to use the Gaussian kernel and rule-of-thumb bandwidths $h_{2}=h_{3}=sd\left(\hat{P}\right)\cdot n^{-1\left/5\right.}$ when estimating $\beta_{d}$ and $g_{d}\left(\cdot\right)$.
Case 2.2: Parametric estimation. Suppose that $g_{d}\left(\cdot\right)$ is parametrically specified as $g_{d}\left(\cdot;\theta_{d}\right)$ for a finite-dimensional parameter $\theta_{d}$. We consider three commonly used specifications.
Case 2.2.1: Polynomial specification. Suppose that $g_{d}\left(\cdot\right)$ is a polynomial function. Then, regress $Y$ on $X$, $\hat{P}$, and the powers of $\hat{P}$ for the treated and untreated groups to estimate $\beta_{d}$ and $g_{d}\left(\cdot\right)$.
Case 2.2.2: Normal specification. Suppose that $\left(U,U_{d}\right)$ follows a bivariate normal distribution, which implies $g_{d}\left(p\right)=\theta_{1d}\left.\phi\left(\Phi^{-1}\left(p\right)\right)\right/\left(p-1+d\right)$. Therefore, regress $Y$ on $X$ and $\left.\phi\left(\Phi^{-1}\left(\hat{P}\right)\right)\right/\left(\hat{P}-1+d\right)$ for each group $d=0,1$ to estimate $\beta_{d}$ and $g_{d}\left(\cdot\right)$.
Case 2.2.3: Normal polynomial specification, which includes the normal specification as a special case when the order of polynomial is one. If the order is set two, then add a term $\Phi^{-1}\left(\hat{P}\right)\left.\phi\left(\Phi^{-1}\left(\hat{P}\right)\right)\right/\left(\hat{P}-1+d\right)$ to the above regression. If the order is set three, further add $\left(\Phi^{-2}\left(\hat{P}\right)+2\right)\left.\phi\left(\Phi^{-1}\left(\hat{P}\right)\right)\right/\left(\hat{P}-1+d\right)$. If the order is set four, further add $\left(\Phi^{-3}\left(\hat{P}\right)+3\Phi^{-1}\left(\hat{P}\right)\right)\left.\phi\left(\Phi^{-1}\left(\hat{P}\right)\right)\right/\left(\hat{P}-1+d\right)$, and so on.
Step 2.3: Estimate the MTE. After obtaining the estimates of $\beta_{d}$ and $g_{d}\left(\cdot\right)$, plug them into the equation (14) to construct the estimator for the MTE.
Table F.2: Support of the nonparametrically estimated propensity score in each data cell
Data cell Age Gender Race Parental education Observations Support of $\hat{\pi}_{0}\left(X^{C}\right)$
1 30 Male Hispanic High school or lower 70 [0, 0.208]
2 30 Male Hispanic College or higher 10 [0, 0.288]
3 30 Male Black High school or lower 93 [0, 0.588]
4 30 Male Black College or higher 21 [0, 0.813]
5 30 Male White High school or lower 147 [0, 0.177]
6 30 Male White College or higher 78 [0, 0.208]
7 30 Female Hispanic High school or lower 65 [0, 0.434]
8 30 Female Hispanic College or higher 11 0
9 30 Female Black High school or lower 100 [0, 0.671]
10 30 Female Black College or higher 20 [0, 0.641]
11 30 Female White High school or lower 125 [0, 0.759]
12 30 Female White College or higher 73 [0, 0.179]
13 31 Male Hispanic High school or lower 87 [0, 0.374]
14 31 Male Hispanic College or higher 19 [0, 0.442]
15 31 Male Black High school or lower 112 [0.030, 0.548]
16 31 Male Black College or higher 24 [0, 0.531]
17 31 Male White High school or lower 148 [0, 0.242]
18 31 Male White College or higher 92 [0, 0.125]
19 31 Female Hispanic High school or lower 97 [0, 0.737]
20 31 Female Hispanic College or higher 14 [0, 0.788]
21 31 Female Black High school or lower 121 [0.249, 0.817]
22 31 Female Black College or higher 24 [0, 0.639]
23 31 Female White High school or lower 149 [0, 0.178]
24 31 Female White College or higher 75 [0, 0.221]
25 32 Male Hispanic High school or lower 65 [0, 0.771]
26 32 Male Hispanic College or higher 13 0
27 32 Male Black High school or lower 133 [0, 0.623]
28 32 Male Black College or higher 23 [0, 0.949]
29 32 Male White High school or lower 175 [0, 0.194]
30 32 Male White College or higher 90 [0, 0.061]
Table F.2 (cont.): Support of the nonparametrically estimated propensity score in each data cell
Data cell Age Gender Race Parental education Observations Support of $\hat{\pi}_{0}\left(X^{C}\right)$
31 32 Female Hispanic High school or lower 94 [0, 0.642]
32 32 Female Hispanic College or higher 17 [0, 0.351]
33 32 Female Black High school or lower 130 [0.021, 0.999]
34 32 Female Black College or higher 34 [0, 1]
35 32 Female White High school or lower 155 [0, 0.526]
36 32 Female White College or higher 81 [0, 0.118]
37 33 Male Hispanic High school or lower 72 [0, 0.567]
38 33 Male Hispanic College or higher 15 [0, 0.743]
39 33 Male Black High school or lower 123 [0, 0.818]
40 33 Male Black College or higher 25 [0, 0.596]
41 33 Male White High school or lower 141 [0, 0.154]
42 33 Male White College or higher 95 [0, 0.133]
43 33 Female Hispanic High school or lower 83 [0, 0.481]
44 33 Female Hispanic College or higher 10 [0, 0.447]
45 33 Female Black High school or lower 125 [0.258, 0.628]
46 33 Female Black College or higher 27 [0, 0.925]
47 33 Female White High school or lower 152 [0, 0.140]
48 33 Female White College or higher 93 [0, 0.350]
49 34 Male Hispanic High school or lower 85 [0, 0.864]
50 34 Male Hispanic College or higher 16 [0, 0.673]
51 34 Male Black High school or lower 101 [0.176, 1]
52 34 Male Black College or higher 27 [0, 0.790]
53 34 Male White High school or lower 162 [0, 0.204]
54 34 Male White College or higher 85 [0, 0.057]
55 34 Female Hispanic High school or lower 73 [0, 0.443]
56 34 Female Hispanic College or higher 11 0
57 34 Female Black High school or lower 109 [0.271, 0.586]
58 34 Female Black College or higher 24 [0, 0.856]
59 34 Female White High school or lower 143 [0, 0.5]
60 34 Female White College or higher 76 [0, 0.055]
Figure F.1: Test of Assumption NL1 for all data cells with size larger than 50
Figure F.1 (cont.): Test of Assumption NL1 for all data cells with size larger than 50
Table G.1: Simulation designs
Design Exclusion restriction Other difference from Design 1 Consistent estimators Inconsistent estimators
1 Yes: $\beta_{1}^{C}=\beta_{0}^{C}=0$ - P, N, S, SIV L, Q
2 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ - P, N, S L, Q, SIV
3 Yes: $\beta_{1}^{C}=\beta_{0}^{C}=0$ Linear treatment function: $\mu\left(X_{i}\right)=X_{i}^{C}+\sum_{j=1}^{5}\frac{1}{j}\left(X_{ji}^{D}-0.5\right)$. SIV P, N, L, Q, S
4 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ Linear treatment function: $\mu\left(X_{i}\right)=X_{i}^{C}+\sum_{j=1}^{5}\frac{1}{j}\left(X_{ji}^{D}-0.5\right)$. - P, N, L, Q, S, SIV
5 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $X_{i}^{C}=\sqrt{1-0.75^{2}}U_{i}+0.75\varrho_{i}$ s.t. $Cov\left(U_{i},X_{i}^{C}\right)\neq 0$. P, N, S L, Q
6 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ (i) Linear treatment function: $\mu\left(X_{i}\right)=X_{i}^{C}+\sum_{j=1}^{5}\frac{1}{j}\left(X_{ji}^{D}-0.5\right)$, and (ii) Heteroskedastic $U_{i}$: $U_{i}=\exp\left(X_{i}^{C}+\frac{1}{2}\left[\left(X_{i}^{C}\right)^{2}-1\right]\right)\epsilon_{i}$, where $\epsilon_{i}\sim N\left(0,1\right)$ and is independent of $X_{i}^{C}$. N, S P, L, Q
7 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ Normal quadratic MTE: $U_{1i}=-\left(\frac{3}{8}U_{i}^{2}+\frac{3}{4}U_{i}-\frac{3}{8}\right)+\sqrt{1-\frac{27}{32}}\varrho_{i}$ s.t. $E\left[\left.U_{1i}\right|V_{i}=v\right]=\frac{3}{8}-\frac{3}{4}\Phi^{-1}\left(v\right)-\frac{3}{8}\Phi^{-2}\left(v\right)$, where $V_{i}=\Phi\left(U_{i}\right)$. S P, N, L, Q
8 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ Normal cubic MTE: $U_{1i}=-\frac{1}{4}U_{i}^{3}+\frac{1}{4}\varrho_{i}$ s.t. $E\left[\left.U_{1i}\right|V_{i}=v\right]=-\frac{1}{4}\Phi^{-3}\left(v\right)$. S P, N, L, Q
9 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $U_{i}\sim t\left(3\right)$ and is standardized to have unit variance. S P, N, L, Q
10 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $U_{i}\sim t\left(2\right)$. S P, N, L, Q
11 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $\mu\left(X_{i}\right)=X_{i}^{C}+\frac{1}{2}\left[\left(X_{i}^{C}\right)^{2}-1\right]+\frac{1}{3}\left(X_{i}^{C}\right)^{3}+\sum_{j=1}^{5}\frac{1}{j}\left(X_{ji}^{D}-0.5\right)$. N, S P, L, Q
12 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ Heteroskedastic $U_{di}$: $U_{1i}=-0.75U_{i}+\sqrt{1-0.75^{2}}\varrho_{i}$ with $\varrho_{i}\sim N\left(0,\exp\left(2X_{i}^{C}-2\right)\right)$, and $U_{0i}\sim N\left(0,\left(X_{i}^{C}\right)^{2}\right)$. P, N, S L, Q
13 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $Y_{1i}=2X_{i}^{C}+\sum_{j=1}^{5}\left(X_{ji}^{D}-0.5\right)+\left(X_{1i}^{D}-0.5\right)\left(X_{2i}^{D}-0.5\right)+U_{1i}$. Correctly specified S P, N, L, Q, S
14 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $Y_{1i}=2X_{i}^{C}+\sum_{j=1}^{5}\left(X_{ji}^{D}-0.5\right)+X_{i}^{C}\left(X_{1i}^{D}-0.5\right)+U_{1i}$. Correctly specified S P, N, L, Q, S
15 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $Y_{1i}=2X_{i}^{C}+\sum_{j=1}^{5}\left(X_{ji}^{D}-0.5\right)+\left[\left(X_{i}^{C}\right)^{2}-1\right]+U_{1i}$. Correctly specified S P, N, L, Q, S
16 No: $\beta_{1}^{C}=2,\beta_{0}^{C}=1$ $Y_{1i}=2X_{i}^{C}+\sum_{j=1}^{5}\left(X_{ji}^{D}-0.5\right)+0.5\left[\exp\left(X_{i}^{C}\right)-\exp\left(0.5\right)\right]+U_{1i}$. Correctly specified S P, N, L, Q, S
Notes: P, N, L, Q, S, and SIV stand for parametric, normal, linear, quadratic, semiparametric, and semiparametric IV estimators for the MTE, respectively. Please refer to Table G.2 for details of the estimators. In Design 1, $U_{i}$ follows a standard normal distribution that is independent of $X_{i}$, and $\mu\left(\cdot\right)$ is quadratic as in (G.1). Design 2 differs from Design 1 only in relaxing the exclusion restriction.
Table G.2: MTE estimators considered in the simulation
Estimator Step 1 Step 2
Parametric MTE Case 1.3: Probit regression of $D_{i}$ on $X_{i}^{C}$, $\left(X_{i}^{C}\right)^{2}$, and $X_{ji}^{D}$. Case 2.2.2: Linear regression of $Y_{i}$ on $X_{i}$ and $\left.\phi\left(\Phi^{-1}\left(\hat{P}_{i}\right)\right)\right/\left(\hat{P}_{i}-1+d\right)$ for each group $d=0,1$.
Normal MTE Case 1.1: Nonparametric regression of $D_{i}$ on $X_{i}^{C}$ and $X_{ji}^{D}$. Case 2.2.2: Linear regression of $Y_{i}$ on $X_{i}$ and $\left.\phi\left(\Phi^{-1}\left(\hat{P}_{i}\right)\right)\right/\left(\hat{P}_{i}-1+d\right)$ for each group $d=0,1$.
Linear MTE Case 1.1: Nonparametric regression of $D_{i}$ on $X_{i}^{C}$ and $X_{ji}^{D}$. Case 2.2.1: Linear regression of $Y_{i}$ on $X_{i}$ and $\hat{P}_{i}$ for each group.
Quadratic MTE Case 1.1: Nonparametric regression of $D_{i}$ on $X_{i}^{C}$ and $X_{ji}^{D}$. Case 2.2.1: Linear regression of $Y_{i}$ on $X_{i}$, $\hat{P}_{i}$, and $\hat{P}_{i}^{2}$ for each group.
Semiparametric MTE Case 1.1: Nonparametric regression of $D_{i}$ on $X_{i}^{C}$ and $X_{ji}^{D}$. Case 2.1: Partially linear regression of $Y_{i}$ on $X_{i}$ and $g_{d}\left(\hat{P}_{i}\right)$ by (13).
Semiparametric IV-MTE Case 1.1: Nonparametric regression of $D_{i}$ on $X_{i}^{C}$ and $X_{ji}^{D}$. Case 2.1: Partially linear regression of $Y_{i}$ on $X_{ji}^{D}$ and $g_{d}\left(\hat{P}_{i}\right)$ by (13) with $X_{i}^{C}$ excluded.
Correctly specified semiparametric MTE Case 1.1: Nonparametric regression of $D_{i}$ on $X_{i}^{C}$ and $X_{ji}^{D}$. Case 2.1: Partially linear regression of $Y_{i}$ on $X_{i}$, $g_{d}\left(\hat{P}_{i}\right)$, and the true nonlinear term in Designs 13-16.
Notes: The italics refer to the cases presented in Table F.1. The semiparametric IV-MTE estimator is considered only in Designs 1-4, while the correctly specified semiparametric MTE estimator is considered only in Designs 13-16.
Figure G.1: Simulated MTE curves for Designs 1-4
Notes: The parametric MTE estimator specifies both parametric first step and parametric second step. The normal, linear, quadratic MTE estimators specify nonparametric first step and parametric second step. The semiparametric MTE estimator specifies nonparametric first step and partially linear (semiparametric) second step. The above five MTE estimators impose no exclusion restriction and base the identification on different functional forms between the treatment and outcome equations, while the semiparametric IV-MTE estimator imposes the exclusion restriction $\beta_{1}^{C}=\beta_{0}^{C}=0$ but otherwise is the same as the semiparametric MTE. Please refer to Table G.2 for more details about the estimators
Table G.3: Simulation results for Designs 1-4
Design 1 Design 2 Design 3 Design 4
Bias RMSE Coverage Bias RMSE Coverage Bias RMSE Coverage Bias RMSE Coverage
Parametric estimator
[…]
ATE 0.051 0.107 0.927 0.033 0.101 0.942 0.045 0.206 0.948 0.045 0.207 0.944
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ -0.001 0.054 0.948 -0.001 0.054 0.948 -0.003 0.125 0.949 -0.003 0.125 0.949
Normal estimator
[…]
ATE 0.041 0.119 0.931 0.032 0.117 0.939 0.278 0.323 0.601 0.278 0.323 0.601
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ -0.035 0.067 0.909 -0.035 0.067 0.909 -0.168 0.193 0.574 -0.168 0.193 0.574
Linear estimator
[…]
ATE -0.015 0.118 0.951 -0.024 0.121 0.954 0.300 0.355 0.652 0.300 0.355 0.651
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ -0.024 0.064 0.924 -0.024 0.064 0.924 -0.188 0.215 0.576 -0.188 0.215 0.576
Quadratic estimator
[…]
ATE 0.043 0.142 0.934 0.034 0.141 0.937 0.538 0.584 0.350 0.537 0.584 0.349
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ -0.028 0.067 0.923 -0.028 0.067 0.923 -0.261 0.284 0.363 -0.261 0.284 0.363
Semiparametric estimator
[…]
ATE 0.041 0.168 0.940 0.044 0.194 0.947 0.621 0.684 0.441 0.629 0.693 0.440
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ -0.029 0.068 0.917 -0.029 0.068 0.922 -0.282 0.307 0.358 -0.287 0.311 0.345
Semiparametric IV estimator
[…]
ATE -0.016 0.119 0.947 6.324 6.330 0 0.005 0.108 0.955 6.868 6.872 0
Notes: RMSE stands for the root mean squared error, and Coverage stands for the coverage probability of asymptotic 95% confidence interval. For the semiparametric IV estimator, $\beta_{1}^{C}$ and $\beta_{0}^{C}$ are assumed to be zero and thus not estimated.
Figure G.2: Simulated MTE curves for Designs 5-8
Notes: The parametric MTE estimator specifies both parametric first step and parametric second step. The normal, linear, quadratic MTE estimators specify nonparametric first step and parametric second step. And the semiparametric MTE estimator specifies nonparametric first step and partially linear (semiparametric) second step.
Table G.4: Simulation results for Designs 5-8
Design 5 Design 6 Design 7 Design 8
Bias RMSE Coverage Bias RMSE Coverage Bias RMSE Coverage Bias RMSE Coverage
Parametric estimator
[…]
ATE 0.042 0.087 0.911 -0.011 0.183 0.347 0.267 0.280 0.124 -0.039 0.110 0.932
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ -0.002 0.038 0.949 0.064 0.098 0.315 -0.062 0.078 0.724 0.038 0.070 0.899
Normal estimator
[…]
ATE -0.083 0.121 0.846 0.106 0.152 0.844 0.285 0.301 0.176 -0.066 0.138 0.916
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ 0.005 0.038 0.954 -0.013 0.054 0.949 -0.083 0.097 0.628 0.018 0.065 0.943
Linear estimator
[…]
ATE -0.128 0.159 0.735 0.037 0.118 0.943 0.156 0.187 0.670 0.002 0.127 0.957
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ 0.007 0.039 0.950 -0.013 0.054 0.956 -0.033 0.061 0.890 -0.024 0.066 0.932
Quadratic estimator
[…]
ATE -0.044 0.121 0.936 0.168 0.222 0.795 0.000 0.124 0.959 0.181 0.234 0.792
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ 0.003 0.039 0.950 -0.005 0.055 0.956 -0.020 0.055 0.933 -0.041 0.074 0.892
Semiparametric estimator
[…]
ATE -0.076 0.163 0.917 0.130 0.209 0.887 0.035 0.155 0.943 0.042 0.179 0.951
$\hat{\beta}_{1}^{C}-\hat{\beta}_{0}^{C}$ 0.006 0.039 0.946 -0.013 0.059 0.941 -0.027 0.058 0.914 -0.018 0.066 0.947
Notes: RMSE stands for the root mean squared error, and Coverage stands for the coverage probability of asymptotic 95% confidence interval.
Figure G.3: Simulated MTE curves for Designs 9-12
Notes: The parametric MTE estimator specifies both parametric first step and parametric second step. The normal, linear, quadratic MTE estimators specify nonparametric first step and parametric second step. And the semiparametric MTE estimator specifies nonparametric first step and partially linear (semiparametric) second step.
Figure G.4: Simulated MTE curves for Designs 13-16
Notes: The parametric, normal, linear, quadratic, and semiparametric MTE estimators are defined as before without including any nonlinear term in the outcome equation, while the correctly specified semiparametric MTE estimator includes the true nonlinear term in the outcome equation.